### v1

# Ke Xua,b, Fen-Fen Yanga,b, Chenggui Yuanc

Abstract
In this work, we establish the existence and uniqueness of solutions to McKean-Vlasov stochastic differential equations (SDEs) driven by Lévy processes with common noise on an infinite time horizon, by means of a contraction mapping principle in the space of probability measures. In addition, we analyse the propagation of chaos for Lévy-driven McKean-Vlasov SDEs in the presence of common noise.
Keywords: McKean-Vlasov equations; Lévy processes; Common noise; Moment estimates; Propagation of chaos
1 Introduction
Let $(\Omega,\mathcal{F},\mathbb{P})$ be a complete probability space. Consider the $d$-dimensional stochastic process $X=(X_{t})_{t\geq 0}$ satisfying the following conditional McKean-Vlasov stochastic differential equation (SDE) with jumps given by:
$\mathrm{d}X(t)$ $=b\bigl{(}X(t),\mathcal{L}^{1}(X(t))\bigr{)}\mathrm{d}t+\sigma\bigl{(}X(t),\mathcal{L}^{1}(X(t))\bigr{)}\mathrm{d}W_{t}$ (1.1)
$\quad+c\bigl{(}X(t),\mathcal{L}^{1}(X(t))\bigr{)}\mathrm{d}Z_{t}+\sigma^{0}\bigl{(}X(t),\mathcal{L}^{1}(X(t))\bigr{)}\mathrm{d}W^{0}_{t},$
for $t\geq 0$, with initial value $X(0)=X_{0}$ and its distribution $\mu=\mathcal{L}(X_{0})$. Here $\mathcal{L}^{1}(X_{t})$ denotes the marginal law of $X$ at time $t$, $W=(W_{t})_{t\geq 0}$ is a $d$-dimensional standard Brownian motion. Denote $\mathbb{R}_{0}^{d}:=\mathbb{R}^{d}\setminus\{0\}$ and let $\mathcal{B}(\mathbb{R}_{+}\times\mathbb{R}_{0}^{d})$ be the Borel $\sigma$-algebra on $\mathbb{R}_{+}\times\mathbb{R}_{0}^{d}$.
$Z_{t}=\int_{0}^{t}\int_{\mathbb{R}_{0}^{d}}z\widetilde{N}(\mathrm{d}s,\mathrm{d}z),\quad t\geq 0,$
where $\widetilde{N}$ is the compensated Poisson random measure associated with the Poisson random measure $N$ on $(\mathbb{R}_{+}\times\mathbb{R}_{0}^{d},\mathcal{B}(\mathbb{R}_{+}\times\mathbb{R}_{0}^{d}))$ with intensity measure $\nu(\mathrm{d}z)\mathrm{d}s$. The compensated Poisson random measure is given by
$\widetilde{N}(\mathrm{d}t,\mathrm{d}z):=N(\mathrm{d}t,\mathrm{d}z)-\nu(\mathrm{d}z)\mathrm{d}t.$
Explicitly, for $t>0$ and $A\in\mathcal{B}(\mathbb{R}_{0}^{d})$,
$N([0,t]\times A):=\sum_{s\in(0,t]}\mathbf{1}_{\{\Delta Z_{s}\in A\}},$
with the jump size of $Z$ at time $s$ defined by
$\Delta Z_{s}:=\begin{cases}Z_{s}-Z_{s-},&s>0,\\ 0,&s=0,\end{cases}\quad\text{where}\quad Z_{s-}:=\lim_{u\uparrow s}Z_{u}.$
Moreover $Z=(Z_{t})_{t\geq 0}$ is a $d$-dimensional centered pure jump Lévy process independent of $W$, whose Lévy measure $\nu$ satisfies
$\int_{\mathbb{R}^{d}}(1\wedge|z|^{2})\nu(\mathrm{d}z)<+\infty.$
Let $|\cdot|$ denote the Euclidean norm on $\mathbb{R}^{d}$. Let $(\mathcal{F}^{0}_{t})_{t\geq 0}$, $(\mathcal{F}^{1}_{t})_{t\geq 0}$ be two complete filtered probability spaces satisfying the usual condition. Let $(W_{t}^{0})_{t\geq 0}$ and $(W_{t})_{t\geq 0}$ be independent $d$-dimensional Brownian motions defined on $(\Omega^{0},\mathcal{F}^{0},\mathbb{P}^{0})$ and $(\Omega^{1},\mathcal{F}^{1},\mathbb{P}^{1})$, respectively. We construct the product probability space $(\Omega,\mathcal{F},\mathbb{P})$ by setting
$\Omega:=\Omega^{0}\times\Omega^{1},\quad(\mathcal{F},\mathbb{P}):=\text{completion of }(\mathcal{F}^{0}\otimes\mathcal{F}^{1},\,\mathbb{P}^{0}\otimes\mathbb{P}^{1}),$
and equipping it with the filtration $(\mathcal{F}_{t})_{t\geq 0}$ obtained as the complete, right-continuous augmentation of $(\mathcal{F}^{0}_{t}\otimes\mathcal{F}^{1}_{t})_{t\geq 0}$. An element $\omega\in\Omega$ means $\omega=(\omega^{0},\omega^{1})$ with $\omega^{0}\in\Omega^{0}$ and $\omega^{1}\in\Omega^{1}$. Expectations with respect to $\mathbb{P}^{0}$ and $\mathbb{P}^{1}$ are denoted by $\mathbb{E}^{0}$ and $\mathbb{E}^{1}$, respectively. For $i\in\{1,2\}$, let $(\Omega_{i}^{1},\mathcal{F}_{i}^{1},\mathbb{P}_{i}^{1})$ be independent copies of $(\Omega^{1},\mathcal{F}^{1},\mathbb{P}^{1})$, with $\mathbb{E}^{1}_{i}$ denoting expectation under $\mathbb{P}_{i}^{1}$. We write $\Omega_{i}:=\Omega^{0}\times\Omega_{i}^{1}$ for the corresponding copies of $\Omega$. Denote by $\mathcal{P}(\mathbb{R}^{d})$ the set of all Borel probability measures on $\mathbb{R}^{d}$ endowed with the topology of weak convergence. For $k>0$, define
$\mathcal{P}_{k}(\mathbb{R}^{d}):=\left\{\mu\in\mathcal{P}(\mathbb{R}^{d}):\int_{\mathbb{R}^{d}}|x|^{k}\,\mu(\mathrm{d}x)<\infty\right\}.$
For $\mu,\nu\in\mathcal{P}_{k}(\mathbb{R}^{d})$, the $k$-Wasserstein distance is given by
$\mathcal{W}_{k}(\mu,\nu):=\inf_{\pi\in\Pi(\mu,\nu)}\left(\int_{\mathbb{R}^{d}\times\mathbb{R}^{d}}|x-y|^{k}\,\pi(\mathrm{d}x,\mathrm{d}y)\right)^{\frac{1}{k\vee 1}},$
where $\Pi(\mu,\nu)$ denotes the set of all couplings of $\mu$ and $\nu$. It is known that $(\mathcal{P}_{k}(\mathbb{R}^{d}),\mathcal{W}_{k})$ forms a Polish metric space. We denote $\delta_{0}$ the Dirac measure concentrated at the origin in $\mathbb{R}^{d}$. Let $\xi$ be a random variable on $\Omega$ measurable with respect to $\mathcal{F}^{0}\otimes\mathcal{F}^{1}$. For any fixed $\omega^{0}\in\Omega^{0}$, the mapping
$\Omega^{1}\ni\omega^{1}\longmapsto\xi(\omega^{0},\omega^{1})$
defines a random variable on $(\Omega^{1},\mathcal{F}^{1},\mathbb{P}^{1})$, whose distribution is denoted by $\mathcal{L}(\xi(\omega^{0},\cdot))$. Given a random variable $\xi$ on $(\Omega,\mathcal{F},\mathbb{P})$, we define
$\mathcal{L}^{1}(\xi):=\mathcal{L}(\xi(\omega^{0},\cdot)),$
which can be interpreted as a mapping from $\Omega^{0}$ into $\mathcal{P}(\mathbb{R}^{d})$. This object represents the conditional distribution of $\xi$ given $\mathcal{F}^{0}$. In particular, if the initial condition $X(0)$ in (1.1) is assumed to be defined on $(\Omega^{1},\mathcal{F}^{1},\mathbb{P}^{1})$, then $\mathcal{L}^{1}(X(t))$ provides a version of the conditional law of $X(t)$ given $W^{0}$. Henceforth, we shall restrict our attention to initial conditions living on $(\Omega^{1},\mathcal{F}^{1},\mathbb{P}^{1})$.
The existence and uniqueness theory for strong solutions of McKean-Vlasov SDEs under Lipschitz and linear growth conditions (with respect to both the state variable and the measure) is well-established (see, e.g., [34], [38]). Wang [42] establishes well-posedness via Picard iteration method resolving pathwise intractability in the Brownian-driven case. In the context of delay McKean-Vlasov SDEs with super-linear drift, Bao et al. [5] propose a Milstein-based antithetic MLMC framework attaining optimal complexity without Lévy area simulation under Brownian noise. Liu et al. [24] establish strong and weak well-posedness of McKean-Vlasov SDEs under locally monotone coefficients in state and distribution variables, and derived a large deviation principle with applications to granular media, plasma, kinetic, porous media, and Navier-Stokes equations.
Compared with Brownian-driven McKean-Vlasov SDEs, the McKean-Vlasov SDEs driven by Lévy processes have been much less studied. Bao and Wang [7] establish existence, uniqueness, and multiplicity of stationary distributions for Lévy-driven McKean-Vlasov SDEs. Song et al. [30] develop large and moderate deviation principles, along with explicit rate functions, for Lévy-driven McKean-Vlasov SDEs using a weak-convergence approach. In [26], Jourdain et al. investigate the existence, uniqueness, and particle approximations for McKean-Vlasov SDEs driven by Lévy processes. More recently, Mehri et al. [35] and Neelima et al. [37] address the well-posedness and propagation of chaos results for McKean-Vlasov SDEs with delay driven by Lévy noise.
The Brownian-driven McKean-Vlasov SDEs with common noise have been applied to various fields including finance, physics, biology, mean-field games, and machine learning. The fundamental theory of the McKean-Vlasov SDEs with common noise is well established. One can refer to Carmona and Delarue [13], Kumar et al. [27] for the well-posedness and conditional propagation of chaos, Hammersley et al. [25] for the weak well-posedness, Maillet [36], Bao and Wang [8] for the long-time behavior. Yuan et al. [16] design delay feedback control strategies for stabilizing McKean-Vlasov SDEs with common noise, proving the global solution existence and uniqueness. Shao and Wei [39] study the well-posedness and conditional propagation of chaos for McKean-Vlasov SDEs with common noise, establishing explicit convergence rates.
An epidemic outbreak occurs when a contagious disease spreads rapidly through a population due to increased transmission rates. In such circumstances, it is reasonable to incorporate a jump process to capture sudden surges in infection numbers triggered by super-spreader events. Based on this perspective, the evolution of the infected population can be modelled by a jump diffusion. Moreover, when global mobility is high, localized outbreaks can quickly propagate across regions, resulting in a pandemic (i.e., a worldwide health crisis in which many areas experience simultaneous infection waves). This observation shows that epidemic dynamics are subject to system-wide randomness (e.g., environmental factors, global travel) that impact all communities. In deep learning, training dynamics can be modeled by Lévy-driven McKean-Vlasov SDEs, where parameter updates follow distribution-dependent stochastic evolution with heavy-tailed noise. Mini-batch gradients and restarts induce jump-like perturbations, while the mean-field term captures interactions among parameters in wide-network limits. In federated or distributed settings, common randomness (e.g., global schedules, network delays) produces correlated dynamics across learners. This framework helps explain heavy-tailed gradient noise, escape from saddles via jumps, and inspires neural SDE architectures that combine diffusion with jumps for expressive modeling. The above considerations motivate the study of Lévy-driven McKean-Vlasov SDEs with common noise [[9],[10]]. In recent years, significant advances have been achieved for Lévy-driven McKean-Vlasov SDEs under common noise, covering topics such as well-posedness, deep learning approaches, optimal stopping problems, optimal and impulse control, as well as stochastic maximum principles; see, for instance, [[1], [2], [3], [9], [10], [22], [23]] for further references.
This paper investigates the infinite time horizon approximation for Lévy-driven McKean-Vlasov SDEs with common noise. The primary contribution of this study lies in addressing a research gap by extending previous work on McKean-Vlasov SDE with Lévy-driven process, where the drift and diffusion coefficients are not globally Lipschitz continuous. Tran et al. [41] propose a tamed-adaptive Euler-Maruyama scheme for Lévy-driven McKean-Vlasov SDEs with superlinear and non-globally Lipschitz coefficients, proving strong convergence on finite and infinite horizons. In particular, this study builds upon the results of Tran et al. [41], introducing the effect of common noise-an important factor that has not been examined in the context of such SDEs. The incorporation of common noise enhances the model, providing a more realistic framework for situations where noise simultaneously impacts all components, a consideration that has often been neglected in prior research.
The main contributions of this article are reflected in the following three aspects.
  1. 1.
    Establishing existence, uniqueness, and moment estimates for solutions under non-globally Lipschitz coefficients (Assumptions (A1)-(A7)), generalizing results in [27, 37, 40].
  2. 2.
    We extend the results of Tran et al. [41] by establishing the boundedness of high-order moments of the solutions and demonstrating the propagation of chaos in the presence of common noise.
  3. 3.
    We apply the theoretical proofs presented in this paper to practical examples and provide corresponding numerical analysis and results.
The remainder of the paper is organized as follows. In Section 2, we establish the well-posedness of solutions to McKean-Vlasov SDE driven by Lévy process with common noise, including existence, uniqueness, and second-order moment estimates under non-globally Lipschitz coefficients. In Section 3, we investigate the propagation of chaos for the associated particle systems. In Section 4, we present the corresponding numerical analysis of the McKean-Vlasov SDE driven by Lévy process with common noise. In Section 5, we provide specific examples to conduct numerical simulations.
2 Well-posedness of Solutions
To establish the well-posedness of (1.1), we assume that the drift, diffusion and jump coefficients $b$, $\sigma$, $\sigma^{0}$, $c$ and the Lévy measure $\nu$ satisfy the following conditions.
(A1) There exists a positive constant $L$ such that
$2\langle x,b(x,\mu)\rangle+|\sigma(x,\mu)|^{2}+|\sigma^{0}(x,\mu)|^{2}+|c(x,\mu)|^{2}\int_{\mathbb{R}^{d}_{0}}|z|^{2}\,\nu({{\rm{d}}}z)\leq L(1+|x|^{2}+\mathcal{W}_{2}^{2}(\mu,\delta_{0})),$
for any $x\in\mathbb{R}^{d}$ and $\mu\in\mathcal{P}_{2}(\mathbb{R}^{d})$.
(A2) There exist constants $L_{1}\in\mathbb{R}$ and $L_{2}\geq 0$ such that
$2\langle x-\overline{x},b(x,\mu)-b(\overline{x},\overline{\mu})\rangle+|\sigma(x,\mu)-\sigma(\overline{x},\overline{\mu})|^{2}+|\sigma^{0}(x,\mu)-\sigma^{0}(\overline{x},\overline{\mu})|^{2}$
$+|c(x,\mu)-c(\overline{x},\overline{\mu})|^{2}\int_{\mathbb{R}^{d}_{0}}|z|^{2}\,\nu({{\rm{d}}}z)\leq L_{1}|x-\overline{x}|^{2}+L_{2}\mathcal{W}_{2}^{2}(\mu,\overline{\mu}),$
for any $x,\overline{x}\in\mathbb{R}^{d}$ and $\mu,\overline{\mu}\in\mathcal{P}_{2}(\mathbb{R}^{d})$.
(A3) There exist constants $L>0$ and $\ell\geq 0$ such that
$|b(x,\mu)-b(\overline{x},\overline{\mu})|\leq L(1+|x|^{\ell}+|\overline{x}|^{\ell})\left(|x-\overline{x}|+\mathcal{W}_{2}(\mu,\overline{\mu})\right),$
for any $x,\overline{x}\in\mathbb{R}^{d}$ and $\mu,\overline{\mu}\in\mathcal{P}_{2}(\mathbb{R}^{d})$.
(A4) There exists an even integer $p_{0}\in[2,+\infty)$ such that
$\int_{|z|>1}|z|^{2p_{0}}\,\nu({{\rm{d}}}z)<\infty\quad\text{and}\quad\int_{0<|z|\leq 1}|z|\,\nu({{\rm{d}}}z)<\infty.$
(A5) There exists a positive constant $L_{3}$ such that
$|c(x,\mu)|\leq L_{3}\left(1+|x|+\mathcal{W}_{2}(\mu,\delta_{0})\right),$
for any $x\in\mathbb{R}^{d}$ and $\mu\in\mathcal{P}_{2}(\mathbb{R}^{d})$.
(A6) For the even integer $p_{0}\in[2,+\infty)$ given in (A4) and the positive constant $L_{3}$ given in (A5), there exist constants $\gamma_{1}\in\mathbb{R}$, $\gamma_{2}\geq 0$, and $\eta\geq 0$ such that
$\langle x,b(x,\mu)\rangle+\frac{p_{0}-1}{2}(|\sigma(x,\mu)|^{2}+|\sigma^{0}(x,\mu)|^{2})+\frac{1}{2L_{3}}|c(x,\mu)|^{2}\int_{\mathbb{R}^{d}_{0}}|z|\left(\left(1+L_{3}|z|\right)^{p_{0}-1}-1\right)\nu({{\rm{d}}}z)$
$\leq\gamma_{1}|x|^{2}+\gamma_{2}\mathcal{W}_{2}^{2}(\mu,\delta_{0})+\eta,$
for any $x\in\mathbb{R}^{d}$ and $\mu\in\mathcal{P}_{2}(\mathbb{R}^{d})$, where $\gamma=\gamma_{1}+\gamma_{2}$ represents the dissipation rate.
(A7) There exist constants $\tilde{L}_{1}\in\mathbb{R}$ and $\tilde{L}_{2}\geq 0$ such that
$|\sigma(x,\mu)-\sigma(\overline{x},\overline{\mu})|^{2}+|\sigma^{0}(x,\mu)-\sigma^{0}(\overline{x},\overline{\mu})|^{2}+|c(x,\mu)-c(\overline{x},\overline{\mu})|^{2}\int_{\mathbb{R}^{d}_{0}}|z|^{2}\,\nu({{\rm{d}}}z)$
$\leq\tilde{L}_{1}|x-\overline{x}|^{2}+\tilde{L}_{2}\mathcal{W}_{2}^{2}(\mu,\overline{\mu}),$
for any $x,\overline{x}\in\mathbb{R}^{d}$ and $\mu,\overline{\mu}\in\mathcal{P}_{2}(\mathbb{R}^{d})$.
Theorem 2.1.
[…]
$\sup_{t\in[0,T]}\mathbb{E}[|X_{t}|^{2}]\leq K,$
where $K:=K(\mathbb{E}|X^{\mu}_{0}|^{2},d,L,L_{1},\gamma_{1},\gamma_{2},T)$ is a positive constant.
Proof.
To make the study easy to follow, we divide the proof into the following steps.
  1. $\bf(1)$}
    Constructing Metric Spaces
[…]
$\rho_{\infty}(\xi,\tilde{\xi}):=\sup_{t\in[0,T]}\left\{e^{-\lambda t}|\xi_{t}-\tilde{\xi}_{t}|\right\},\quad\xi,\tilde{\xi}\in\mathcal{D}([0,T];\mathbb{R}^{d}),\quad\lambda>0.$
Then, the space $\mathcal{D}([0,T];\mathbb{R}^{d})$ endowed with the distance $\rho_{\infty}$ is a complete space.
For any given $T>0$ and $p\geq 1$, let $\mathscr{Y}_{T}(p)$ be a space consisting of all processes $\eta:[0,T]\times\Omega\to\mathbb{R}^{d}$ satisfying
  • (i)
    For almost all $\omega\in\Omega$, $\eta_{.}(\omega):[0,T]\to\mathbb{R}^{d}$ is right-continuous with left limits;
  • (ii)
    The process $t\mapsto\eta_{t}$ is $\mathcal{F}_{t}$-adapted and satisfies $\sup_{t\in[0,T]}\mathbb{E}[|\eta_{t}|^{p}]<\infty$.
Based on the distance $\rho_{\infty}$ on $\mathcal{D}([0,T];\mathbb{R}^{d})$, for each $p\geq 1,$ with $\lambda$ as above, we define a metric on $\mathscr{Y}_{T}(p)$ by
$\bm{\rho}_{p,\infty}(\eta,\tilde{\eta}):=\bigg{[}\mathbb{E}\rho^{p}_{\infty}(\eta,\tilde{\eta})\bigg{]}^{1/p},\quad\eta,\tilde{\eta}\in\mathscr{Y}_{T}(p).$
The following lemma focuses on the completeness of $(\mathscr{Y}_{T}(p),\bm{\rho}_{p,\infty})$, which is essentially due to Shao et al. [40, Lemma 2.3 ]. We include a complete proof for readers’ convenience.
Lemma 2.1.
The metric space $(\mathscr{Y}_{T}(p),\bm{\rho}_{p,\infty})$ is complete for every $p\geq 1$.
Proof.
Let $(X^{(n)})_{n\geq 1}$ be a Cauchy sequence in $(\mathscr{Y}_{T}(p),\bm{\rho}_{p,\infty})$, i.e.,
$\bm{\rho}_{p,\infty}(X^{(n)},X^{(m)})^{p}\to 0\quad\text{as }n,m\to\infty.$
Take a subsequence $(X^{(n_{k})})$ such that
$\bm{\rho}_{p,\infty}(X^{(n_{k})},X^{(n_{k+1})})<\frac{1}{2^{k}},\quad k\geq 1,$
which implies
$\mathbb{P}\left(\rho_{\infty}(X^{(n_{k})},X^{(n_{k+1})})\geq\frac{1}{k^{2}}\right)\leq\frac{k^{2p}}{2^{kp}},\quad k\geq 1,$
Since $\sum_{k\geq 1}\frac{k^{2p}}{2^{kp}}<\infty$, the Borel-Cantelli lemma yields
$\mathbb{P}\left(\rho_{\infty}(X^{(n_{k})},X^{(n_{k+1})})\geq\frac{1}{k^{2}},\text{ i.o.}\right)=0.$
Thus, there exists a measurable set $\Xi$ with $\mathbb{P}(\Xi)=1$ such that for every $\omega\in\Xi$,
$\sum_{k=1}^{\infty}\rho_{\infty}(X^{(n_{k})}(\omega),X^{(n_{k+1})}(\omega))<\infty.$
By the completeness of $\mathcal{D}([0,T];\mathbb{R}^{d})$ with respect to $\rho_{\infty}$, for every $\omega\in\Xi$, there exists $\tilde{X}(\omega)\in\mathcal{D}([0,T];\mathbb{R}^{d})$ such that
$\lim_{k\to\infty}\rho_{\infty}\bigl{(}X^{(n_{k})}(\omega),\,\tilde{X}(\omega)\bigr{)}=0,\quad\omega\in\Xi.$
We now aim to prove that $\tilde{X}\in\mathscr{Y}_{T}(p)$ and that
$\bm{\rho}_{p,\infty}\big{(}X^{(n_{k})},\tilde{X}\big{)}\xrightarrow{\;k\to\infty\;}0.$
We first show that the Cauchy sequence $(X^{(n)})_{n\geq 1}$ is uniformly $p$-integrable. Since $(X^{(n)})$ is Cauchy in $(\mathscr{Y}_{T}(p),\bm{\rho}_{p,\infty})$, there exists some fixed index $n_{0}$ and a constant $C<\infty$ such that
$\sup_{n\geq 1}\bm{\rho}_{p,\infty}\big{(}X^{(n)},X^{(n_{0})}\big{)}^{p}=\sup_{n\geq 1}\mathbb{E}\Big{[}\sup_{t\in[0,T]}e^{-\lambda t}|X^{(n)}_{t}-X^{(n_{0})}_{t}|^{p}\Big{]}\leq C.$
Using the elementary inequality $|a+b|^{p}\leq 2^{p-1}(|a|^{p}+|b|^{p})$, we obtain for every $n\geq 1$
$\mathbb{E}\Big{[}\sup_{t\in[0,T]}e^{-\lambda t}|X^{(n)}_{t}|^{p}\Big{]}$ $\leq 2^{p-1}\mathbb{E}\Big{[}\sup_{t\in[0,T]}e^{-\lambda t}|X^{(n)}_{t}-X^{(n_{0})}_{t}|^{p}\Big{]}+2^{p-1}\mathbb{E}\Big{[}\sup_{t\in[0,T]}e^{-\lambda t}|X^{(n_{0})}_{t}|^{p}\Big{]}$
$\leq 2^{p-1}C+2^{p-1}\mathbb{E}\Big{[}\sup_{t\in[0,T]}e^{-\lambda t}|X^{(n_{0})}_{t}|^{p}\Big{]}=:C_{1}<\infty.$
Hence $\sup_{n\geq 1}\mathbb{E}\big{[}\sup_{t\in[0,T]}e^{-\lambda t}|X^{(n)}_{t}|^{p}\big{]}\leq C_{1}<\infty$. Now consider the limit process $\tilde{X}$. By Fatou’s lemma and the Cauchy property of $X^{(n_{k})}\to\tilde{X}$,
$\mathbb{E}\Big{[}\sup_{t\in[0,T]}e^{-\lambda t}|\tilde{X}_{t}|^{p}\Big{]}\leq\liminf_{k\to\infty}\mathbb{E}\Big{[}\sup_{t\in[0,T]}e^{-\lambda t}|X^{(n_{k})}_{t}|^{p}\Big{]}\leq C_{1}<\infty.$
[…]
$\sup_{t\in[0,T]}\mathbb{E}[|\tilde{X}_{t}|^{p}]\leq e^{\lambda T}\,\mathbb{E}\Big{[}\sup_{s\in[0,T]}e^{-\lambda s}|\tilde{X}_{s}|^{p}\Big{]}.$
Since the right-hand side has already been shown to be finite (bounded by $C_{1}$), we conclude
$\sup_{t\in[0,T]}\mathbb{E}[|\tilde{X}_{t}|^{p}]\leq e^{\lambda T}C_{1}<\infty.$
Thus we have $\sup_{t\in[0,T]}\mathbb{E}[|\tilde{X}_{t}|^{p}]<\infty$.
Note that for every $\omega\in\Xi$, the path $\tilde{X}(\omega)$ belongs to $\mathcal{D}([0,T];\mathbb{R}^{d})$. Thus, it remains only to verify that $\tilde{X}_{t}$ is $\mathcal{F}_{t}$-measurable for each $t\in[0,T]$. Let $\tilde{\Gamma}$ denote the law of $\tilde{X}$ as a random element of $\mathcal{D}([0,T];\mathbb{R}^{d})$. We introduce the set
$\mathcal{F}_{\tilde{\Gamma}}:=\{\,t\in[0,T]\mid\tilde{\Gamma}(J_{t})=0\,\},$
where
$J_{t}:=\{\,X\in\mathcal{D}([0,T];\mathbb{R}^{d}):X_{t}\neq X_{t-}\,\}$
[…]
$\lim_{k\to\infty}|X_{t}^{(n_{k})}-\tilde{X}_{t}|=0,\quad\mathbb{P}\text{-a.e.},\forall t\in\mathcal{F}_{\tilde{\Gamma}}$.
Since $X_{t}^{(n_{k})}$ is $\mathcal{F}_{t}$-adapted, $\tilde{X}_{t}$ is $\mathcal{F}_{t}$-adapted for $t\in\mathcal{F}_{\tilde{\Gamma}}$. Due to right‐continuity of the paths, for any $t\in(0,T)$, we choose a sequence $t_{n}\downarrow t$ such that $\tilde{X}_{t_{n}}\;\xrightarrow{\;n\to\infty\;}\;\tilde{X}_{t}\quad\mathbb{P}\text{-a.s.}.$ Since $\mathcal{F}_{t+}=\mathcal{F}_{t}$, it follows that $\tilde{X}_{t}\;\in\;\mathcal{F}_{t}.$ Hence, $\tilde{X}\in\mathscr{Y}_{T}(p)$. By Fatou’s lemma and the Cauchy property,
[…]
$=$ $\lim_{n\to\infty}\bigg{[}\mathbb{E}\sup_{t\in[0,T]}e^{-\lambda t}|X^{(n)}_{t}-\lim_{l\to\infty}X^{(n_{l})}_{t}|^{p}\bigg{]}^{1/p}$
$\leq$ $\lim_{n\to\infty}\mathop{\underline{\lim}}\limits_{l\to\infty}\bigg{[}\mathbb{E}\sup_{t\in[0,T]}e^{-\lambda t}|X^{(n)}_{t}-X^{(n_{l})}_{t}|^{p}\bigg{]}^{1/p}=0.$
Thus, $(\mathscr{Y}_{T}(p),\bm{\rho}_{p,\infty})$ is complete. ∎
  1. $\bf(2)$}
    Freezing the distribution
For a given $\eta\in\mathscr{Y}_{T}(2)$, denote $\mu_{t}=\mathcal{L}(\eta_{t}|\mathcal{F}_{t}^{0})$, we construct the following classic SDE
$Y_{t}^{\mu}=Y_{0}^{\mu}+\int_{0}^{t}b(Y_{s}^{\mu},\mu_{s}){{\rm{d}}}s+\int_{0}^{t}\sigma(Y_{s}^{\mu},\mu_{s}){{\rm{d}}}W_{s}+\int_{0}^{t}c(Y_{s}^{\mu},\mu_{s}){{\rm{d}}}Z_{s}+\int_{0}^{t}\sigma^{0}(Y_{s}^{\mu},\mu_{s}){{\rm{d}}}W_{s}^{0},$ (2.1)
almost surely for any $t\in[0,T],$ with $Y_{0}^{\mu}=X_{0}$. According to [37, Theorem 2.1], the classical SDE (2.1) admits a unique solution with $\sup_{t\in[0,T]}\mathbb{E}\bigl{[}\lvert Y_{t}^{\mu}\rvert^{2}\bigr{]}<\infty.$ Moreover, we have $(Y_{t}^{\mu})_{t\in[0,T]}\in\mathscr{Y}_{T}(2)$.
  1. $\bf(3)$}
    Constructing a contractive mapping
Define the mapping
$\Phi:\mathscr{Y}_{T}(2)\longrightarrow\mathscr{Y}_{T}(2),\quad\eta\mapsto Y^{\mu}.$
We show that $\Phi$ is a strict contraction in $(\mathscr{Y}_{T}(2),\bm{\rho}_{2,\infty})$. For another $\tilde{\eta}\in\mathscr{Y}_{T}(2)$, let $(Y_{t}^{\nu})$ solve SDE (2.1) by replacing $\mu_{t}=\mathcal{L}(\eta_{t}|\mathcal{F}_{t}^{0})$ there with $\nu_{t}=\mathcal{L}(\tilde{\eta}_{t}|\mathcal{F}_{t}^{0})$ for $t\in[0,T]$. Applying Itô’s formula to $|Y_{t}^{\mu}-Y_{t}^{\nu}|^{2}$, taking the supremum over $t\in[0,T]$, taking expectation and using the Assumption (A2), we get
$\mathbb{E}\Big{[}\sup_{0\leq s\leq t}\big{|}Y_{s}^{\mu}-Y_{s}^{\nu}\big{|}^{2}\Big{]}$ (2.2)
$\leq\mathbb{E}\sup_{0\leq r\leq t}\int_{0}^{r}\Big{[}2\langle Y_{s}^{\mu}-Y_{s}^{\nu},b(Y_{s}^{\mu},\mu_{s})-b(Y_{s}^{\nu},\nu_{s})\rangle$
$\quad+|\sigma(Y_{s}^{\mu},\mu_{s})-\sigma(Y_{s}^{\nu},\nu_{s})|^{2}+|\sigma^{0}(Y_{s}^{\mu},\mu_{s})-\sigma^{0}(Y_{s}^{\nu},\nu_{s})|^{2}\Big{]}{\rm{d}}s$
$\quad+\mathbb{E}[\sup_{0\leq r\leq t}\int_{0}^{r}\int_{\mathbb{R}_{0}^{d}}|(c(Y_{s}^{\mu},\mu_{s})-c(Y_{s}^{\nu},\nu_{s}))z|^{2}\nu({\rm{d}}z){\rm{d}}s]$
$\quad+2\,\mathbb{E}\left[\sup_{0\leq r\leq t}\int_{0}^{r}\left\langle Y_{s}^{\mu}-Y_{s}^{\nu},(\sigma(Y_{s}^{\mu},\mu_{s})-\sigma(Y_{s}^{\nu},\nu_{s}))\mathrm{d}W_{s}\right\rangle\right]$
$\quad+2\,\mathbb{E}\left[\sup_{0\leq r\leq t}\int_{0}^{r}\left\langle Y_{s}^{\mu}-Y_{s}^{\nu},(\sigma^{0}(Y_{s}^{\mu},\mu_{s})-\sigma^{0}(Y_{s}^{\nu},\nu_{s}))\mathrm{d}W_{s}^{0}\right\rangle\right]$
$\quad+2\,\mathbb{E}\left[\sup_{0\leq r\leq t}\int_{0}^{r}\int_{\mathbb{R}_{0}^{d}}\Big{(}|Y_{s^{-}}^{\mu}-Y_{s^{-}}^{\nu}+\delta_{s}(z)|^{2}-|Y_{s^{-}}^{\mu}-Y_{s^{-}}^{\nu}|^{2}\Big{)}\widetilde{N}(\mathrm{d}s,\mathrm{d}z)\right]$
$\leq\mathbb{E}\Big{[}\!\sup_{0\leq r\leq t}\int_{0}^{r}\big{(}L_{1}|Y_{s}^{\mu}-Y_{s}^{\nu}|^{2}+L_{2}\mathcal{W}_{2}^{2}(\mu_{s},\nu_{s})\big{)}\,\mathrm{d}s\Big{]}$
$\quad+2\,\mathbb{E}\left[\sup_{0\leq r\leq t}\int_{0}^{r}\left\langle Y_{s}^{\mu}-Y_{s}^{\nu},(\sigma(Y_{s}^{\mu},\mu_{s})-\sigma(Y_{s}^{\nu},\nu_{s}))\mathrm{d}W_{s}\right\rangle\right]$
$\quad+2\,\mathbb{E}\left[\sup_{0\leq r\leq t}\int_{0}^{r}\left\langle Y_{s}^{\mu}-Y_{s}^{\nu},(\sigma^{0}(Y_{s}^{\mu},\mu_{s})-\sigma^{0}(Y_{s}^{\nu},\nu_{s}))\mathrm{d}W_{s}^{0}\right\rangle\right]$
$\quad+2\,\mathbb{E}\left[\sup_{0\leq r\leq t}\int_{0}^{r}\int_{\mathbb{R}_{0}^{d}}\Big{(}|Y_{s^{-}}^{\mu}-Y_{s^{-}}^{\nu}+\delta_{s}(z)|^{2}-|Y_{s^{-}}^{\mu}-Y_{s^{-}}^{\nu}|^{2}\Big{)}\widetilde{N}(\mathrm{d}s,\mathrm{d}z)\right]$
Applying the Burkholder-Davis-Gundy (BDG) inequality, Jensen’s inequality and using the Assumption (A7) to each stochastic integral, we obtain
$\mathbb{E}\left[\sup_{0\leq r\leq t}\left|\int_{0}^{r}\langle Y_{s}^{\mu}-Y_{s}^{\nu},(\sigma(Y_{s}^{\mu},\mu_{s})-\sigma(Y_{s}^{\nu},\nu_{s}))\mathrm{d}W_{s}\rangle\right|\right]$
$\quad\leq C_{1}\,\mathbb{E}\left[\left(\int_{0}^{t}|Y_{s}^{\mu}-Y_{s}^{\nu}|^{2}\cdot|\sigma(Y_{s}^{\mu},\mu_{s})-\sigma(Y_{s}^{\nu},\nu_{s})|^{2}\mathrm{d}s\right)^{1/2}\right]$
$\quad\leq C_{1}\,\mathbb{E}\left[\left(\int_{0}^{t}|Y_{s}^{\mu}-Y_{s}^{\nu}|^{2}\cdot\left(\tilde{L}_{1}|Y_{s}^{\mu}-Y_{s}^{\nu}|^{2}+\tilde{L}_{2}\mathcal{W}_{2}^{2}(\mu_{s},\nu_{s})\right)\mathrm{d}s\right)^{1/2}\right]$
$\quad\leq K_{1}\int_{0}^{t}\left(\mathbb{E}[\sup_{0\leq r\leq s}e^{-\lambda r}|Y_{r}^{\mu}-Y_{r}^{\nu}|^{2}+\sup_{0\leq r\leq s}e^{-\lambda r}\mathcal{W}_{2}^{2}(\mu_{r},\nu_{r})]\right){\rm{d}}s,$
where $K_{1}:=K_{1}(C_{1},\tilde{L}_{1},\tilde{L}_{2})$ and similarly for the $W^{0}$ term and the jump term with $K_{2},K_{3}$ respectively. Therefore,
$\mathbb{E}\Big{[}\sup_{0\leq s\leq t}\big{|}Y_{s}^{\mu}-Y_{s}^{\nu}\big{|}^{2}\Big{]}$ (2.3)
$\leq\mathbb{E}\Big{[}\!\sup_{0\leq r\leq t}\int_{0}^{r}\big{(}L_{1}|Y_{s}^{\mu}-Y_{s}^{\nu}|^{2}\big{)}\,\mathrm{d}s\Big{]}+\int_{0}^{t}\mathbb{E}[\sup_{0\leq r\leq s}(K_{1}+K_{2}+K_{3})|Y_{r}^{\mu}-Y_{r}^{\nu}|^{2}$
$\quad+\sup_{0\leq r\leq s}(L_{2}+K_{1}+K_{2}+K_{3})\mathcal{W}_{2}^{2}(\mu_{r},\nu_{r})]{\rm{d}}s$
$\leq\int_{0}^{t}\mathbb{E}[\sup_{0\leq r\leq s}(L_{1}+K_{1}+K_{2}+K_{3})|Y_{r}^{\mu}-Y_{r}^{\nu}|^{2}+\sup_{0\leq r\leq s}(L_{2}+K_{1}+K_{2}+K_{3})\mathcal{W}_{2}^{2}(\mu_{r},\nu_{r})]{\rm{d}}s$
$\leq\hat{C}\int_{0}^{t}\left(\mathbb{E}[\sup_{0\leq r\leq s}|Y_{r}^{\mu}-Y_{r}^{\nu}|^{2}+\sup_{0\leq r\leq s}\mathcal{W}_{2}^{2}(\mu_{r},\nu_{r})]\right){\rm{d}}s,$
where $\hat{C}:=\hat{C}(L_{1},L_{2},K_{1},K_{2},K_{3})$.
Note that for any $s\geq 0$,
$\sup_{0\leq r\leq s}|Y_{r}^{\mu}-Y_{r}^{\nu}|^{2}=\sup_{0\leq r\leq s}\left(e^{\lambda r}\cdot e^{-\lambda r}|Y_{r}^{\mu}-Y_{r}^{\nu}|^{2}\right)\leq e^{\lambda s}\sup_{0\leq r\leq s}e^{-\lambda r}|Y_{r}^{\mu}-Y_{r}^{\nu}|^{2}.$ (2.4)
For convenience, we denote
$\phi(t):=\mathbb{E}\left[\sup_{0\leq s\leq t}e^{-\lambda s}\left|Y_{s}^{\mu}-Y_{s}^{\nu}\right|^{2}\right].$
Taking expectations to (2.4), we obtain
$\mathbb{E}\left[\sup_{0\leq r\leq s}|Y_{r}^{\mu}-Y_{r}^{\nu}|^{2}\right]\leq e^{\lambda s}\,\phi(s).$
Similarly,
$\sup_{0\leq r\leq s}\mathcal{W}_{2}^{2}(\mu_{r},\nu_{r})\leq e^{\lambda s}\sup_{0\leq r\leq s}e^{-\lambda r}\mathcal{W}_{2}^{2}(\mu_{r},\nu_{r}).$
Hence, according to (2.3) we have
$\phi(t)$ $\leq\mathbb{E}\Big{[}\sup_{0\leq s\leq t}\big{|}Y_{s}^{\mu}-Y_{s}^{\nu}\big{|}^{2}\Big{]}$
$\leq\hat{C}\int_{0}^{t}\left(\mathbb{E}[\sup_{0\leq r\leq s}|Y_{r}^{\mu}-Y_{r}^{\nu}|^{2}+\sup_{0\leq r\leq s}\mathcal{W}_{2}^{2}(\mu_{r},\nu_{r})]\right){\rm{d}}s$
$\leq\hat{C}\int_{0}^{t}\left(e^{\lambda s}\,\phi(s)+e^{\lambda s}\mathbb{E}[\sup_{0\leq r\leq s}e^{-\lambda r}\mathcal{W}_{2}^{2}(\mu_{r},\nu_{r})]\right){\rm{d}}s$
$\leq\hat{C}e^{\lambda T}\int_{0}^{t}\left(\,\phi(s)+\mathbb{E}[\sup_{0\leq r\leq s}e^{-\lambda r}\mathcal{W}_{2}^{2}(\mu_{r},\nu_{r})]\right){\rm{d}}s.$
By using Gronwall’s inequality, we have
$\phi(t)$ $\leq\hat{C}e^{\lambda T}\int_{0}^{t}\mathbb{E}[\sup_{0\leq r\leq s}e^{-\lambda r}\mathcal{W}_{2}^{2}(\mu_{r},\nu_{r})]e^{\hat{C}e^{\lambda T}(t-s)}{\rm{d}}s$
$\leq\hat{C}e^{\lambda T}\cdot\mathbb{E}[\sup_{0\leq r\leq t}e^{-\lambda r}\mathcal{W}_{2}^{2}(\mu_{r},\nu_{r})]\int_{0}^{t}e^{\hat{C}e^{\lambda T}(t-s)}{\rm{d}}s$
$=(e^{\hat{C}e^{\lambda T}t}-1)\mathbb{E}[\sup_{0\leq r\leq t}e^{-\lambda r}\mathcal{W}_{2}^{2}(\mu_{r},\nu_{r})].$
Therefore, we have
$\phi(t)$ $=\mathbb{E}[\sup_{0\leq s\leq t}e^{-\lambda s}\left|Y_{s}^{\mu}-Y_{s}^{\nu}\right|^{2}]$
$\leq(e^{\hat{C}e^{\lambda T}t}-1)\cdot\mathbb{E}[\sup_{0\leq r\leq t}e^{-\lambda r}\mathcal{W}_{2}^{2}(\mu_{r},\nu_{r})]$
$\leq(e^{\hat{C}e^{\lambda T}t}-1)\cdot\mathbb{E}[\sup_{t\in[0,T]}e^{-\lambda t}|X_{t}-\tilde{X}_{t}|^{2}].$
Combining the definition of the metric on space $(\mathscr{Y}_{T}(2),\bm{\rho}_{2,\infty})$, it holds that
[…]
$\leq(e^{\hat{C}e^{\lambda T}t}-1)^{1/2}\bigg{[}\mathbb{E}\sup_{t\in[0,T]}e^{-\lambda t}|X_{t}-\tilde{X}_{t}|^{2}\bigg{]}^{1/2}$
[…]
Taking $\lambda=ln{(\frac{ln2}{2T\hat{C}})}^{\frac{1}{T}},$ then we have $0<(e^{\hat{C}e^{\lambda T}t}-1)^{1/2}<1$.
By the contraction mapping principle, the mapping
$\Phi:\mathscr{Y}_{T}(2)\to\mathscr{Y}_{T}(2)$
is a strict contraction on the complete metric space $\bigl{(}\mathscr{Y}_{T}(2),\bm{\rho}_{2,\infty}\bigr{)}$. Hence there exists a unique fixed point $X^{*}\in\mathscr{Y}_{T}(2)$ such that
$\Phi(X^{*})\;=\;X^{*}.$
By definition of $\Phi$, $X^{*}$ satisfies
$X^{*}_{t}$ $=X^{\mu}_{0}+\int_{0}^{t}b\bigl{(}X^{*}_{s},\mathcal{L}(X^{*}_{s}\mid\mathcal{F}^{0}_{s})\bigr{)}\,{{\rm{d}}}s+\int_{0}^{t}\sigma\bigl{(}X^{*}_{s},\mathcal{L}(X^{*}_{s}\mid\mathcal{F}^{0}_{s})\bigr{)}\,{{\rm{d}}}W_{s}+\int_{0}^{t}c\bigl{(}X^{*}_{s},\mathcal{L}(X^{*}_{s}\mid\mathcal{F}^{0}_{s})\bigr{)}\,{{\rm{d}}}Z_{s}$
$+\int_{0}^{t}\sigma^{0}\bigl{(}X^{*}_{s},\mathcal{L}(X^{*}_{s}\mid\mathcal{F}^{0}_{s})\bigr{)}\,{\rm{d}}W^{0}_{s},$
which is precisely equation (1.1). Uniqueness follows immediately: if $Y\in\mathscr{Y}_{T}(2)$ is any other solution of (1.1), then $Y=\Phi(Y)$, and by uniqueness of the fixed point $Y=X^{*}$. Therefore, SDE (1.1) with initial distribution $X^{\mu}_{0}$ has a unique solution.
  1. $\bf(4)$}
    Proving the second-order moment boundedness
Apply Itô’s formula to $|X_{t}|^{2}$. Using Assumption (A1)
$2\langle x,b(x,\mu)\rangle+|\sigma(x,\mu)|^{2}+|\sigma^{0}(x,\mu)|^{2}+\int_{Z}|c(x,\mu)z|^{2}\nu(dz)\leq L(1+|x|^{2}+\mathcal{W}_{2}^{2}(\mu,\delta_{0})).$
Taking expectation and noting $\mathbb{E}[\mathcal{W}_{2}^{2}(\mathcal{L}^{1}(X_{s}),\delta_{0})]\leq\mathbb{E}|X_{s}|^{2}$, we derive
$\mathbb{E}|X_{t}|^{2}\leq|x_{0}|^{2}+L\int_{0}^{t}(1+2\mathbb{E}|X_{s}|^{2}){\rm{d}}s.$
By Gronwall’s inequality, we obtain that
$\sup_{t\in[0,T]}\mathbb{E}|X_{t}|^{2}\leq\left(|x_{0}|^{2}+LT\right)e^{2LT}:=K.$
This establishes existence, uniqueness, and the second-order moment boundedness. ∎
Remark 2.2.
Similar with Tran et al.[41]. It follows from Assumption (A6) that,
$\langle x,b(x,\mu)\rangle+\frac{p_{0}-1}{2}(|\sigma(x,\mu)|^{2}+|\sigma^{0}(x,\mu)|^{2})+\frac{1}{2L_{3}}|c(x,\mu)|^{2}\int_{\mathbb{R}^{d}_{0}}|z|\left(\left(1+L_{3}|z|\right)^{p_{0}-1}-1\right)\nu({{\rm{d}}}z)$
$\leq\gamma_{1}|x|^{2}+\gamma_{2}\mathcal{W}_{2}^{2}(\mu,\delta_{0})+\eta,$
for any $p\in[2,p_{0}]$, $x\in\mathbb{R}^{d}$ and $\mu\in\mathcal{P}_{2}(\mathbb{R}^{d})$.
Proposition 2.3.
[…]
$\mathbb{E}[|X_{t}|^{p}]\leq\begin{cases}C_{p}(1+e^{\gamma pt})&\text{if }\gamma\neq 0,\\ C_{p}(1+t)^{p/2}&\text{if }\gamma=0,p=2\text{ or }\gamma=0,\gamma_{2}>0,p\in(2,p_{0}],\\ C_{p}(1+t)^{p}&\text{if }\gamma=0,\gamma_{2}=0,p\in(2,p_{0}],\end{cases}$ (2.5)
where $\gamma=\gamma_{1}+\gamma_{2}$.
$\text{If }\gamma<0,\text{ then }\sup_{t\geq 0}\mathbb{E}\bigl{[}|X_{t}|^{p}\bigr{]}\leq 2C_{p}.$
Proof.
We hereby follow the methodology introduced in [41, Proposition 2.3 ].
Step 1: It follows from Theorem 2.1 that $\sup_{t\in[0,T]}\mathbb{E}[|X_{t}|^{2}]\leq K,K:=K(\mathbb{E}|X^{\mu}_{0}|^{2},d,L,L_{1},\gamma_{1},\gamma_{2},T).$
Step 2: We first demonstrate that for any even natural number $p\in[2,p_{0}]$ and for any $T>0$, the following holds,
$\sup_{t\in[0,T]}\mathbb{E}[|X_{t}|^{p}]\leq C(T,p).$ (2.6)
Observe that (2.6) is valid for the case $p=2$ as established in Step 1. Assume now, as an induction hypothesis, that (2.6) holds for every even integer $q\in[2,\,p-2]$, namely
$\sup_{t\in[0,T]}\mathbb{E}\big{[}\,|X_{t}|^{q}\,\big{]}\leq C(T,q).$ (2.7)
Let $\lambda\in\mathbb{R}$ and consider any even integer $p\in[2,\,p_{0}]$. By applying Itô’s formula to the process $e^{-\lambda t}|X_{t}|^{p}$, we obtain
$e^{-\lambda t}|X_{t}|^{p}=|x_{0}|^{p}+\int_{0}^{t}e^{-\lambda s}\Big{[}-\lambda|X_{s}|^{p}+p|X_{s}|^{p-2}\langle X_{s},b(X_{s},\mathcal{L}^{1}_{X_{s}})\rangle+\frac{p}{2}|X_{s}|^{p-2}|\sigma(X_{s},\mathcal{L}^{1}_{X_{s}})|^{2}$ (2.8)
$+\frac{p}{2}|X_{s}|^{p-2}|\sigma^{0}(X_{s},\mathcal{L}^{1}_{X_{s}})|^{2}+\frac{p(p-2)}{2}|X_{s}|^{p-4}\big{|}X_{s}^{\top}\sigma(X_{s},\mathcal{L}^{1}_{X_{s}})\big{|}^{2}$
$+\frac{p(p-2)}{2}|X_{s}|^{p-4}\big{|}X_{s}^{\top}\sigma^{0}(X_{s},\mathcal{L}^{1}_{X_{s}})\big{|}^{2}\Big{]}{{\rm{d}}}s+p\int_{0}^{t}e^{-\lambda s}|X_{s}|^{p-2}\Big{[}\langle X_{s},\sigma(X_{s},\mathcal{L}^{1}_{X_{s}}){{\rm{d}}}W_{s}\rangle$
$+\langle X_{s},\sigma^{0}(X_{s},\mathcal{L}^{1}_{X_{s}}){{\rm{d}}}W^{0}_{s}\rangle\Big{]}+\int_{0}^{t}\int_{\mathbb{R}_{0}^{d}}e^{-\lambda s}\Big{(}|X_{s^{-}}+c(X_{s^{-}},\mathcal{L}^{1}_{X_{s^{-}}})z|^{p}-|X_{s^{-}}|^{p}\Big{)}\widetilde{N}({{\rm{d}}}s,{{\rm{d}}}z)$
$+\int_{0}^{t}\int_{\mathbb{R}_{0}^{d}}e^{-\lambda s}\Big{(}|X_{s}+c(X_{s},\mathcal{L}^{1}_{X_{s}})z|^{p}-|X_{s}|^{p}-p|X_{s}|^{p-2}\langle X_{s},c(X_{s},\mathcal{L}^{1}_{X_{s}})z\rangle\Big{)}\nu({{\rm{d}}}z){{\rm{d}}}s.$
To handle the last integral in (2.8), it suffices to apply the binomial theorem, which yields, for any $t\geq 0$,
$|X_{s}+c(X_{s},\mathcal{L}^{1}_{X_{s}})z|^{p}=\left(|X_{s}|^{2}+|c(X_{s},\mathcal{L}^{1}_{X_{s}})z|^{2}+2\langle X_{s},c(X_{s},\mathcal{L}^{1}_{X_{s}})z\rangle\right)^{p/2}$ (2.9)
$=|X_{s}|^{p}+\frac{p}{2}|X_{s}|^{p-2}\left(|c(X_{s},\mathcal{L}^{1}_{X_{s}})z|^{2}+2\langle X_{s},c(X_{s},\mathcal{L}^{1}_{X_{s}})z\rangle\right)$
$\quad+\sum_{i=2}^{p/2}\binom{p/2}{i}|X_{s}|^{p-2i}\left(|c(X_{s},\mathcal{L}^{1}_{X_{s}})z|^{2}+2\langle X_{s},c(X_{s},\mathcal{L}^{1}_{X_{s}})z\rangle\right)^{i}$
$=|X_{s}|^{p}+p|X_{s}|^{p-2}\langle X_{s},c(X_{s},\mathcal{L}^{1}_{X_{s}})z\rangle+\frac{p}{2}|X_{s}|^{p-2}|c(X_{s},\mathcal{L}^{1}_{X_{s}})z|^{2}$
$\quad+\sum_{i=2}^{p/2}\binom{p/2}{i}|X_{s}|^{p-2i}\left(|c(X_{s},\mathcal{L}^{1}_{X_{s}})z|^{2}+2\langle X_{s},c(X_{s},\mathcal{L}^{1}_{X_{s}})z\rangle\right)^{i}.$
Lemma 2.4.
For any $s\geq 0$, we have
$\mathcal{W}^{2}_{2}(\mathcal{L}^{1}_{X_{s}},\delta_{0})=\mathbb{E}[|X_{s}|^{2}\mid\mathcal{F}^{0}],$
where $\mathcal{L}^{1}_{X_{s}}$ denotes the conditional law of $X_{s}$ given $\mathcal{F}^{0}$.
Proof.
Recall that for two probability measures $\mu$ and $\nu$ on $\mathbb{R}^{d}$, the 2-Wasserstein distance is defined as
$\mathcal{W}_{2}^{2}(\mu,\nu)=\inf_{\pi\in\Pi(\mu,\nu)}\iint_{\mathbb{R}^{d}\times\mathbb{R}^{d}}|x-y|^{2}{\rm{d}}\pi(x,y),$
where $\Pi(\mu,\nu)$ denotes the set of all couplings of $\mu$ and $\nu$.
In our case, $\nu=\delta_{0}$ is the Dirac measure at the origin. Since $\delta_{0}$ is a deterministic measure, there is only one possible coupling $\pi$ between $\mathcal{L}^{1}_{X_{s}}$ and $\delta_{0}$, namely the product measure
$\pi=\mathcal{L}^{1}_{X_{s}}\otimes\delta_{0}.$
Now compute the cost with this specific coupling
$\mathcal{W}_{2}^{2}(\mathcal{L}^{1}_{X_{s}},\delta_{0})$ $=\iint_{\mathbb{R}^{d}\times\mathbb{R}^{d}}|x-y|^{2}{\rm{d}}\pi(x,y)$
$=\iint_{\mathbb{R}^{d}\times\mathbb{R}^{d}}|x-y|^{2}{\rm{d}}(\mathcal{L}^{1}_{X_{s}}\otimes\delta_{0})(x,y)$
$=\int_{\mathbb{R}^{d}}\left(\int_{\mathbb{R}^{d}}|x-y|^{2}\delta_{0}({\rm{d}}y)\right)\mathcal{L}^{1}_{X_{s}}({\rm{d}}x).$
Since $\delta_{0}$ is concentrated at $y=0$, the inner integral evaluates to
$\int_{\mathbb{R}^{d}}|x-y|^{2}\delta_{0}({\rm{d}}y)=|x-0|^{2}=|x|^{2}.$
Therefore,
$\mathcal{W}_{2}^{2}(\mathcal{L}^{1}_{X_{s}},\delta_{0})=\int_{\mathbb{R}^{d}}|x|^{2}\mathcal{L}^{1}_{X_{s}}({\rm{d}}x)=\mathbb{E}[|X_{s}|^{2}\mid\mathcal{F}^{0}].$
This completes the proof. ∎
Next, by repeatedly applying the binomial theorem, using Assumption (A5), the identity
$\mathcal{W}^{2}_{2}(\mathcal{L}^{1}_{X_{s}},\delta_{0})=\mathbb{E}[|X_{s}|^{2}\mid\mathcal{F}^{0}],$
and Young’s inequality
$y|x|^{p-3}\leq\frac{1}{2}(|x|^{p-2}+y^{2}|x|^{p-4}),$
which holds for any $x\in\mathbb{R}^{d}$, $y>0$, as well as the binomial identities
$\sum_{j=0}^{i}\binom{i}{j}a^{j}=(1+a)^{i},\quad\sum_{j=0}^{i}\binom{i}{j}ja^{j}=ia(1+a)^{i-1}$
valid for any $a\in\mathbb{R}$, we obtain,
$|X_{s}|^{p-2i}\Big{(}|c(X_{s},\mathcal{L}^{1}_{X_{s}})z|^{2}+2\langle X_{s},c(X_{s},\mathcal{L}^{1}_{X_{s}})z\rangle\Big{)}^{i}$
$=|X_{s}|^{p-2i}\sum_{j=0}^{i}\binom{i}{j}|c(X_{s},\mathcal{L}^{1}_{X_{s}})z|^{2i-2j}2^{j}\langle X_{s},c(X_{s},\mathcal{L}^{1}_{X_{s}})z\rangle^{j}$
$\leq\sum_{j=0}^{i}\binom{i}{j}2^{j}|X_{s}|^{p-2i+j}|c(X_{s},\mathcal{L}^{1}_{X_{s}})|^{2i-j}|z|^{2i-2j}$
$=|c(X_{s},\mathcal{L}^{1}_{X_{s}})|^{2}\sum_{j=0}^{i}\binom{i}{j}2^{j}|X_{s}|^{p-2i+j}|c(X_{s},\mathcal{L}^{1}_{X_{s}})|^{2i-j-2}|z|^{2i-j}$
$\leq|c(X_{s},\mathcal{L}^{1}_{X_{s}})|^{2}\sum_{j=0}^{i}\binom{i}{j}2^{j}|X_{s}|^{p-2i+j}|z|^{2i-j}L_{3}^{2i-j-2}(1+|X_{s}|+\mathcal{W}_{2}(\mathcal{L}^{1}_{X_{s}},\delta_{0}))^{2i-j-2}$
$=|c(X_{s},\mathcal{L}^{1}_{X_{s}})|^{2}\sum_{j=0}^{i}\binom{i}{j}2^{j}|X_{s}|^{p-2i+j}|z|^{2i-j}L_{3}^{2i-j-2}\Big{(}1+|X_{s}|+\sqrt{\mathbb{E}[|X_{s}|^{2}|\mathcal{F}^{0}]}\Big{)}^{2i-j-2}$
$=|c(X_{s},\mathcal{L}^{1}_{X_{s}})|^{2}\sum_{j=0}^{i}\binom{i}{j}2^{j}|X_{s}|^{p-2i+j}|z|^{2i-j}L_{3}^{2i-j-2}\Bigg{(}|X_{s}|^{2i-j-2}$
$\quad+(2i-j-2)\Big{(}1+\sqrt{\mathbb{E}[|X_{s}|^{2}|\mathcal{F}^{0}]}\Big{)}|X_{s}|^{2i-j-3}+\sum_{k=2}^{2i-j-2}\binom{2i-j-2}{k}$
$\quad\times\Big{(}1+\sqrt{\mathbb{E}[|X_{s}|^{2}|\mathcal{F}^{0}]}\Big{)}^{k}|X_{s}|^{2i-j-2-k}\Bigg{)}$
$=|c(X_{s},\mathcal{L}^{1}_{X_{s}})|^{2}\sum_{j=0}^{i}\binom{i}{j}2^{j}|z|^{2i-j}L_{3}^{2i-j-2}\Bigg{(}|X_{s}|^{p-2}+(2i-j-2)\Big{(}1+\sqrt{\mathbb{E}[|X_{s}|^{2}|\mathcal{F}^{0}]}\Big{)}|X_{s}|^{p-3}$
$\quad+\sum_{k=2}^{2i-j-2}\binom{2i-j-2}{k}\Big{(}1+\sqrt{\mathbb{E}[|X_{s}|^{2}|\mathcal{F}^{0}]}\Big{)}^{k}|X_{s}|^{p-2-k}\Bigg{)}$
$\leq|c(X_{s},\mathcal{L}^{1}_{X_{s}})|^{2}\sum_{j=0}^{i}\binom{i}{j}2^{j}|z|^{2i-j}L_{3}^{2i-j-2}\Bigg{(}|X_{s}|^{p-2}+\frac{(2i-j-2)}{2}\Big{(}|X_{s}|^{p-2}$
$\quad+\Big{(}1+\sqrt{\mathbb{E}[|X_{s}|^{2}|\mathcal{F}^{0}]}\Big{)}^{2}|X_{s}|^{p-4}\Bigg{)}+\sum_{k=2}^{2i-j-2}\binom{2i-j-2}{k}\Big{(}1+\sqrt{\mathbb{E}[|X_{s}|^{2}|\mathcal{F}^{0}]}\Big{)}^{k}|X_{s}|^{p-2-k}\Bigg{)}$
$=|c(X_{s},\mathcal{L}^{1}_{X_{s}})|^{2}|X_{s}|^{p-2}\sum_{j=0}^{i}\binom{i}{j}2^{j}|z|^{2i-j}L_{3}^{2i-j-2}(i-\frac{j}{2})$
$\quad+|c(X_{s},\mathcal{L}^{1}_{X_{s}})|^{2}\sum_{j=0}^{i}\binom{i}{j}2^{j}|z|^{2i-j}L_{3}^{2i-j-2}\Bigg{(}\Big{(}i-\frac{j}{2}-1\Big{)}\Big{(}1+\sqrt{\mathbb{E}[|X_{s}|^{2}|\mathcal{F}^{0}]}\Big{)}^{2}|X_{s}|^{p-4}$
$\quad+\sum_{k=2}^{2i-j-2}\binom{2i-j-2}{k}\Big{(}1+\sqrt{\mathbb{E}[|X_{s}|^{2}|\mathcal{F}^{0}]}\Big{)}^{k}|X_{s}|^{p-2-k}\Bigg{)}$
$=|c(X_{s},\mathcal{L}^{1}_{X_{s}})|^{2}|X_{s}|^{p-2}\big{(}\frac{i}{L_{3}^{2}}(L_{3}^{2}|z|^{2}+2L_{3}|z|)^{i}-\frac{|z|}{L_{3}}i(L_{3}^{2}|z|^{2}+2L_{3}|z|)^{i}\big{)}$
$\quad+|c(X_{s},\mathcal{L}^{1}_{X_{s}})|^{2}\sum_{j=0}^{i}\binom{i}{j}2^{j}|z|^{2i-j}L_{3}^{2i-j-2}\Bigg{(}\Big{(}i-\frac{j}{2}-1\Big{)}\Big{(}1+\sqrt{\mathbb{E}[|X_{s}|^{2}|\mathcal{F}^{0}]}\Big{)}^{2}|X_{s}|^{p-4}$
$\quad+\sum_{k=2}^{2i-j-2}\binom{2i-j-2}{k}\Big{(}1+\sqrt{\mathbb{E}[|X_{s}|^{2}|\mathcal{F}^{0}]}\Big{)}^{k}|X_{s}|^{p-2-k}\Bigg{)}$
$\leq|c(X_{s},\mathcal{L}^{1}_{X_{s}})|^{2}|X_{s}|^{p-2}\left(\frac{i}{L_{3}^{2}}(L_{3}^{2}|z|^{2}+2L_{3}|z|)^{i}-\frac{|z|i}{L_{3}}(L_{3}^{2}|z|^{2}+2L_{3}|z|)^{i-1}\right)$
$\quad+{L_{3}^{2}}\Big{(}1+|X_{s}|+\sqrt{\mathbb{E}[|X_{s}|^{2}|\mathcal{F}^{0}]}\Big{)}^{2}\sum_{j=0}^{i}\binom{i}{j}2^{j}|z|^{2i-j}L_{3}^{2i-j-2}\Bigg{(}(i-\frac{j}{2}-1)(1+\sqrt{\mathbb{E}[|X_{s}|^{2}|\mathcal{F}^{0}]})^{2}|X_{s}|^{p-4}$
$\quad+\sum_{k=2}^{2i-j-2}\binom{2i-j-2}{k}(1+\sqrt{\mathbb{E}[|X_{s}|^{2}|\mathcal{F}^{0}]})^{k}|X_{s}|^{p-2-k}\Bigg{)}$
$=|c(X_{s},\mathcal{L}^{1}_{X_{s}})|^{2}|X_{s}|^{p-2}\frac{i|z|}{L_{3}}(1+L_{3}|z|)(L_{3}^{2}|z|^{2}+2L_{3}|z|)^{i-1}$
$\quad+\sum_{j=0}^{i}|z|^{2i-j}Q_{p}\big{(}p-2,2i-j,|X_{s}|,1+\sqrt{\mathbb{E}[|X_{s}|^{2}|\mathcal{F}^{0}]}\big{)}$
$where\quad 2i-j\geq 2\quad\text{and}\quad$
$Q_{q}(m,n,x,y)=\sum_{\ell_{1}\leq m,\,\ell_{2}\leq n,\,\ell_{1}+\ell_{2}=q}c_{\ell_{1}\ell_{2}}x^{\ell_{1}}y^{\ell_{2}}.$
Combined with the identity
$\sum_{i=2}^{p/2}\binom{p/2}{i}ia^{i-1}=\frac{p}{2}\left((1+a)^{p/2-1}-1\right),$
which holds for any $a\in\mathbb{R}$, this yields,
$\sum_{i=2}^{p/2}\binom{p/2}{i}|X_{s}|^{p-2i}\left(|c(X_{s},\mathcal{L}^{1}_{X_{s}})z|^{2}+2\langle X_{s},c(X_{s},\mathcal{L}^{1}_{X_{s}})z\rangle\right)^{i}$ (2.10)
$\leq\frac{|z|}{L_{3}}\left(1+L_{3}|z|\right)|c(X_{s},\mathcal{L}^{1}_{X_{s}})|^{2}|X_{s}|^{p-2}\sum_{i=2}^{p/2}\binom{p/2}{i}i\left(L_{3}^{2}|z|^{2}+2L_{3}|z|\right)^{i-1}$
$\quad+\sum_{i=2}^{p/2}\sum_{j=0}^{i}\binom{p/2}{i}|z|^{2i-j}Q_{p}(p-2,2i-j,|X_{s}|,1+\sqrt{\mathbb{E}[|X_{s}|^{2}|\mathcal{F}^{0}]})$
$=\frac{p}{2}|X_{s}|^{p-2}|c(X_{s},\mathcal{L}^{1}_{X_{s}})|^{2}\frac{|z|}{L_{3}}\left((1+L_{3}|z|)^{p-1}-L_{3}|z|-1\right)$
$\quad+\sum_{i=2}^{p/2}\sum_{j=0}^{i}\binom{p/2}{i}|z|^{2i-j}Q_{p}(p-2,2i-j,|X_{s}|,1+\sqrt{\mathbb{E}[|X_{s}|^{2}|\mathcal{F}^{0}]}).$
As a consequence of (2.9) and (2.10), we conclude that for any $s\geq 0$,
$|X_{s}+c(X_{s},\mathcal{L}^{1}_{X_{s}})z|^{p}-|X_{s}|^{p}-p|X_{s}|^{p-2}\langle X_{s},c(X_{s},\mathcal{L}^{1}_{X_{s}})z\rangle$ (2.11)
$\leq\frac{p}{2}|X_{s}|^{p-2}|c(X_{s},\mathcal{L}^{1}_{X_{s}})|^{2}\Bigg{(}|z|^{2}+\frac{|z|}{L_{3}}\left((1+L_{3}|z|)^{p-1}-L_{3}|z|-1\right)\Bigg{)}$
$\quad+\sum_{i=2}^{p/2}\sum_{j=0}^{i}\binom{p/2}{i}|z|^{2i-j}Q_{p}(p-2,2i-j,|X_{s}|,1+\sqrt{\mathbb{E}[|X_{s}|^{2}|\mathcal{F}^{0}]})$
$=\frac{p}{2L_{3}}|X_{s}|^{p-2}|c(X_{s},\mathcal{L}^{1}_{X_{s}})|^{2}|z|\left((1+L_{3}|z|)^{p-1}-1\right)$
$\quad+\sum_{i=2}^{p/2}\sum_{j=0}^{i}\binom{p/2}{i}|z|^{2i-j}Q_{p}(p-2,2i-j,|X_{s}|,1+\sqrt{\mathbb{E}[|X_{s}|^{2}|\mathcal{F}^{0}]}).$
Under Assumption (A4), the Lévy measure satisfies
$\int_{|z|>1}|z|^{2p_{0}}\nu({\rm{d}}z)<\infty,$
which ensures the convergence of the higher-order jump integrals in (2.9). Therefore, by substituting (2.11) into (2.8), we obtain that for any $t\geq 0$,
$e^{-\lambda t}|X_{t}|^{p}\leq|x_{0}|^{p}+p\int_{0}^{t}e^{-\lambda s}|X_{s}|^{p-2}\Bigg{[}-\frac{\lambda}{p}|X_{s}|^{2}+\langle X_{s},b(X_{s},\mathcal{L}^{1}_{X_{s}})\rangle+\frac{p-1}{2}|\sigma(X_{s},\mathcal{L}^{1}_{X_{s}})|^{2}$
$+\frac{p-1}{2}|\sigma^{0}(X_{s},\mathcal{L}^{1}_{X_{s}})|^{2}+\frac{1}{2L_{3}}|c(X_{s},\mathcal{L}^{1}_{X_{s}})|^{2}\int_{\mathbb{R}^{d}_{0}}|z|\bigg{(}(1+L_{3}|z|)^{p-1}-1\bigg{)}\nu({{\rm{d}}}z)\Bigg{]}{{\rm{d}}}s$
$+\int_{0}^{t}\int_{\mathbb{R}^{d}_{0}}e^{-\lambda s}\sum_{i=2}^{p/2}\sum_{j=0}^{i}\binom{p/2}{i}|z|^{2i-j}Q_{p}\bigg{(}p-2,2i-j,|X_{s}|,1+\sqrt{\mathbb{E}[|X_{s}|^{2}|\mathcal{F}^{0}]}\bigg{)}\nu({{\rm{d}}}z){{\rm{d}}}s$
$+p\int_{0}^{t}e^{-\lambda s}|X_{s}|^{p-2}\bigg{[}\langle X_{s},\sigma(X_{s},\mathcal{L}^{1}_{X_{s}}){{\rm{d}}}W_{s}\rangle+\langle X_{s},\sigma^{0}(X_{s},\mathcal{L}^{1}_{X_{s}}){{\rm{d}}}W^{0}_{s}\rangle\Bigg{]}$
$+\int_{0}^{t}\int_{\mathbb{R}^{d}_{0}}e^{-\lambda s}\Big{(}|X_{s-}+c(X_{s-},\mathcal{L}^{1}_{X_{s}-})z|^{p}-|X_{s-}|^{p}\Big{)}\tilde{N}({{\rm{d}}}s,{{\rm{d}}}z).$
Now, for each $N>0$, define the stopping time
$\tau_{N}:=\inf\{t\geq 0:|X_{t}|\geq N\}.$
By choosing $\lambda=\gamma_{1}p$ and applying Assumption (A6), Remark 2.2, together with the identity
$\mathcal{W}_{2}^{2}(\mathcal{L}^{1}_{X_{s}},\delta_{0})=\mathbb{E}[|X_{s}|^{2}\mid\mathcal{F}^{0}],$
we obtain,
$\mathbb{E}\left[e^{-\gamma_{1}p(t\wedge\tau_{N})}|X_{t\wedge\tau_{N}}|^{p}\right]\leq|x_{0}|^{p}+p\int_{0}^{t}e^{-\gamma_{1}ps}\mathbb{E}\Big{[}|X_{s}|^{p-2}\Big{]}\Big{(}\gamma_{2}\mathcal{W}_{2}^{2}(\mathcal{L}^{1}_{X_{s}},\delta_{0})+\eta\Big{)}{{\rm{d}}}s$ (2.12)
$\quad+\int_{0}^{t}\int_{\mathbb{R}^{d}_{0}}e^{-\gamma_{1}ps}\sum_{i=2}^{p/2}\sum_{j=0}^{i}\binom{p/2}{i}|z|^{2i-j}\mathbb{E}\left[Q_{p}\left(p-2,2i-j,|X_{s}|,1+\sqrt{\mathbb{E}[|X_{s}|^{2}|\mathcal{F}^{0}]}\right)\right]\nu({{\rm{d}}}z){{\rm{d}}}s$
$\leq|x_{0}|^{p}+p\int_{0}^{t}e^{-\gamma_{1}ps}\mathbb{E}\left[|X_{s}|^{p-2}\right]\big{(}\gamma_{2}\mathbb{E}\left[|X_{s}|^{2}|\mathcal{F}^{0}\right]+\eta\big{)}{{\rm{d}}}s$
$\quad+\int_{0}^{t}\int_{\mathbb{R}^{d}_{0}}e^{-\gamma_{1}ps}\sum_{i=2}^{p/2}\sum_{j=0}^{i}\binom{p/2}{i}|z|^{2i-j}\mathbb{E}\left[Q_{p}\left(p-2,2i-j,|X_{s}|,1+\sqrt{\mathbb{E}[|X_{s}|^{2}|\mathcal{F}^{0}]}\right)\right]\nu({{\rm{d}}}z){{\rm{d}}}s.$
Next, using the inequality
$e^{-\gamma_{1}p(t\wedge\tau_{N})}\geq e^{-\gamma_{1}pt},$
and the induction hypothesis (2.7), there exists a positive constant $C(T,p)$, independent of $N$, such that
$\sup_{t\in[0,T]}\mathbb{E}\left[|X_{t\wedge\tau_{N}}|^{p}\right]\leq C(T,p).$ (2.13)
This yields
$\sup_{t\in[0,T]}\mathbb{P}(\tau_{N}<t)\leq\frac{C(T,p)}{N^{p}},$
which implies that $\tau_{N}\to\infty$ almost surely as $N\to\infty$. Now, letting $N\uparrow\infty$ and applying Fatou’s lemma to the left-hand side of (2.13), we obtain
$\sup_{t\in[0,T]}\mathbb{E}\left[|X_{t}|^{p}\right]\leq C(T,p).$
Thus, by the principle of induction, we have established (2.6).
Step 3: We next aim to prove (2.5) for every even natural number $p\in[2,p_{0}]$.
First, by applying (2.8) with $\lambda=2\gamma$ and $p=2$, and using Assumption (A6), we obtain,
$e^{-2\gamma t}|X_{t}|^{2}=|x_{0}|^{2}+\int_{0}^{t}e^{-2\gamma s}\bigg{[}-2\gamma|X_{s}|^{2}+2\langle X_{s},b(X_{s},\mathcal{L}^{1}_{X_{s}})\rangle+|\sigma(X_{s},\mathcal{L}^{1}_{X_{s}})|^{2}+|\sigma^{0}(X_{s},\mathcal{L}^{1}_{X_{s}})|^{2}\bigg{]}{{\rm{d}}}s$
$\quad+2\int_{0}^{t}e^{-2\gamma s}\bigg{[}\langle X_{s},\sigma(X_{s},\mathcal{L}^{1}_{X_{s}}){{\rm{d}}}W_{s}\rangle+\langle X_{s},\sigma^{0}(X_{s},\mathcal{L}^{1}_{X_{s}}){{\rm{d}}}W^{0}_{s}\rangle\Bigg{]}$
$\quad+\int_{0}^{t}\int_{\mathbb{R}^{d}_{0}}e^{-2\lambda s}\Big{(}|X_{s-}+c(X_{s-},\mathcal{L}^{1}_{X_{s}-})z|^{2}-|X_{s-}|^{2}\Big{)}\tilde{N}({{\rm{d}}}s,{{\rm{d}}}z)$
$\quad+\int_{0}^{t}\int_{\mathbb{R}^{d}_{0}}e^{-2\gamma s}|c(X_{s},\mathcal{L}^{1}_{X_{s}})z|^{2}\nu({{\rm{d}}}z){{\rm{d}}}s$
$\leq|x_{0}|^{2}+2\int_{0}^{t}e^{-2\gamma s}\bigg{[}-\gamma|X_{s}|^{2}+\langle X_{s},b(X_{s},\mathcal{L}^{1}_{X_{s}})\rangle+\frac{1}{2}|\sigma(X_{s},\mathcal{L}^{1}_{X_{s}})|^{2}+\frac{1}{2}|\sigma^{0}(X_{s},\mathcal{L}^{1}_{X_{s}})|^{2}$
$\quad+\frac{1}{2}|c(X_{s},\mathcal{L}^{1}_{X_{s}})|^{2}\int_{\mathbb{R}^{d}_{0}}|z|^{2}\nu({{\rm{d}}}z)\bigg{]}{{\rm{d}}}s$
$\quad+2\int_{0}^{t}e^{-2\gamma s}\bigg{[}\langle X_{s},\sigma(X_{s},\mathcal{L}^{1}_{X_{s}}){{\rm{d}}}W_{s}\rangle+\langle X_{s},\sigma^{0}(X_{s},\mathcal{L}^{1}_{X_{s}}){{\rm{d}}}W_{s}^{0}\rangle\bigg{]}$
$\quad+\int_{0}^{t}\int_{\mathbb{R}^{d}_{0}}e^{-2\gamma s}\Big{(}|X_{s-}+c(X_{s-},\mathcal{L}^{1}_{X_{s}-})z|^{2}-|X_{s-}|^{2}\Big{)}\tilde{N}({{\rm{d}}}s,{{\rm{d}}}z)$
$\leq|x_{0}|^{2}+2\int_{0}^{t}e^{-2\gamma s}\left(-\gamma_{2}|X_{s}|^{2}+\gamma_{2}\mathcal{W}_{2}^{2}(\mathcal{L}^{1}_{X_{s}},\delta_{0})+\eta\right){{\rm{d}}}s$
$\quad+2\int_{0}^{t}e^{-2\gamma s}\bigg{[}\langle X_{s},\sigma(X_{s},\mathcal{L}^{1}_{X_{s}}){{\rm{d}}}W_{s}+\sigma^{0}(X_{s},\mathcal{L}^{1}_{X_{s}}){{\rm{d}}}W_{s}^{0}\rangle\bigg{]}$
$\quad+\int_{0}^{t}\int_{\mathbb{R}^{d}_{0}}e^{-2\gamma s}\Big{(}|X_{s-}+c(X_{s-},\mathcal{L}^{1}_{X_{s}-})z|^{2}-|X_{s-}|^{2}\Big{)}\tilde{N}({{\rm{d}}}s,{{\rm{d}}}z).$
From the identity
$\mathcal{W}_{2}^{2}\big{(}\mathcal{L}^{1}_{X_{s}},\delta_{0}\big{)}=\mathbb{E}\big{[}\,|X_{s}|^{2}\,\big{|}\,\mathcal{F}^{0}\big{]},$
and by applying the estimate (2.6), we deduce
$\mathbb{E}\!\left[e^{-2\gamma t}|X_{t}|^{2}\right]\leq|x_{0}|^{2}+2\int_{0}^{t}e^{-2\gamma s}\mathbb{E}\!\left[-\gamma_{2}|X_{s}|^{2}+\gamma_{2}\mathbb{E}[|X_{s}|^{2}\mid\mathcal{F}^{0}]+\eta\right]\mathrm{d}s\leq|x_{0}|^{2}+2\eta\int_{0}^{t}e^{-2\gamma s}\,\mathrm{d}s.$
This yields
$\mathbb{E}[|X_{t}|^{2}]\leq\begin{cases}\big{(}|x_{0}|^{2}+\tfrac{\eta}{\gamma}\big{)}e^{2\gamma t}-\tfrac{\eta}{\gamma},&\gamma\neq 0,\\[3.99994pt] |x_{0}|^{2}+2\eta t,&\gamma=0,\end{cases}$
so that (2.5) holds in the case $p=2$. Assume now, as an induction hypothesis, that (2.5) is valid for every even integer $q\in[2,\,p-2]$, that is,
$\mathbb{E}[|X_{t}|^{q}]\leq\begin{cases}C_{q}\big{(}1+e^{\gamma qt}\big{)},&\gamma\neq 0,\\[3.99994pt] C_{q}(1+t)^{q/2},&\gamma=0,\ \big{(}q=2\ \text{or}\ (\gamma_{2}>0,\ q\in(2,p-2])\big{)},\\[3.99994pt] C_{q}(1+t)^{q},&\gamma=0,\ \gamma_{2}=0,\ q\in(2,p-2].\end{cases}$ (2.14)
We prove that (2.5) also holds for the even integer $p$. The argument relies on (2.12), the inductive assumption (2.14), and Assumption (A4).
Case $\gamma\neq 0$.
We have
$\mathbb{E}\!\left[e^{-\gamma_{1}p(t\wedge\tau_{N})}|X_{t\wedge\tau_{N}}|^{p}\right]\leq|x_{0}|^{p}+C\int_{0}^{t}e^{-\gamma_{1}ps}\big{(}1+e^{\gamma ps}\big{)}\mathrm{d}s=|x_{0}|^{p}+C\int_{0}^{t}\big{(}e^{-\gamma_{1}ps}+e^{\gamma_{2}ps}\big{)}\mathrm{d}s.$
Since $\tau_{N}\to\infty$ almost surely as $N\to\infty$, Fatou’s lemma gives
$\mathbb{E}\!\left[e^{-\gamma_{1}pt}|X_{t}|^{p}\right]\leq|x_{0}|^{p}+C\int_{0}^{t}\big{(}e^{-\gamma_{1}ps}+e^{\gamma_{2}ps}\big{)}\mathrm{d}s.$
If $\gamma_{1}=0$, then
$\mathbb{E}[|X_{t}|^{p}]\leq|x_{0}|^{p}+C\big{(}1+t+e^{\gamma_{2}pt}\big{)}\leq C\big{(}1+e^{\gamma pt}\big{)}.$
If $\gamma_{1}\neq 0$, we obtain
$\mathbb{E}[|X_{t}|^{p}]\leq|x_{0}|^{p}e^{\gamma_{1}pt}+\frac{C}{p\gamma_{1}}\big{(}e^{\gamma_{1}pt}-1\big{)}+\frac{C}{p\gamma_{2}}e^{\gamma pt}\leq C\big{(}1+e^{\gamma pt}\big{)}.$
Case $\gamma=0$.
If $\gamma_{2}>0$, then
$\mathbb{E}\!\left[e^{-\gamma_{1}p(t\wedge\tau_{N})}|X_{t\wedge\tau_{N}}|^{p}\right]\leq|x_{0}|^{p}+C\int_{0}^{t}e^{-\gamma_{1}ps}(1+s)^{p/2}\mathrm{d}s\leq|x_{0}|^{p}+C(1+t)^{p/2}\!\!\int_{0}^{t}e^{-\gamma_{1}ps}\mathrm{d}s.$
Letting $N\to\infty$ and using $\gamma_{1}=\gamma_{2}$, we find
$\mathbb{E}\!\left[e^{\gamma_{2}pt}|X_{t}|^{p}\right]\leq|x_{0}|^{p}+\frac{C}{\gamma_{2}p}(1+t)^{p/2}e^{\gamma_{2}pt},$
hence $\mathbb{E}[|X_{t}|^{p}]\leq C(1+t)^{p/2}$.
If $\gamma_{2}=\gamma_{1}=0$, we have
$\mathbb{E}[|X_{t\wedge\tau_{N}}|^{p}]\leq|x_{0}|^{p}+C\int_{0}^{t}(1+s)^{p-1}\mathrm{d}s\leq C(1+t)^{p}.$
Letting $N\to\infty$ yields $\mathbb{E}[|X_{t}|^{p}]\leq C(1+t)^{p}$.
Therefore, (2.5) holds for the given even $p$. By induction, it follows that (2.5) is valid for all even integers $p\in[2,\,p_{0}]$, and by Hölder’s inequality, for all real $p\in[2,\,p_{0}]$. This completes the proof.
3 Propagation of chaos
For each $N\in\mathbb{N}$, let $(W^{i},Z^{i})$, $i=1,\ldots,N$, be independent copies of the pair $(W,Z)$. Denote by $N^{i}(dt,dz)$ the Poisson random measure corresponding to the jumps of the Lévy process $Z^{i}$ with intensity measure $\nu(dz)dt$, and define the compensated Poisson random measure by
$\widetilde{N}^{i}({\rm{d}}t,{\rm{d}}z):=N^{i}({\rm{d}}t,{\rm{d}}z)-\nu({\rm{d}}z){\rm{d}}t.$
Then, the Lévy-Itô decomposition of $Z^{i}$ is given by
$Z_{t}^{i}=\int_{0}^{t}\int_{\mathbb{R}_{0}^{d}}z\,\widetilde{N}^{i}({\rm{d}}s,{\rm{d}}z),\quad t\geq 0.$
We now consider the system of non-interacting particles associated with the Lévy-driven McKean-Vlasov SDE with common noise (1.1), where the state $X^{i}=(X_{t}^{i})_{t\geq 0}$ of the $i$-th particle is defined by
$\begin{split}X_{t}^{i}&=x_{0}+\int_{0}^{t}b(X_{s}^{i},\mathcal{L}^{1}_{X_{s}^{i}}){{\rm{d}}}s+\int_{0}^{t}\sigma(X_{s}^{i},\mathcal{L}^{1}_{X_{s}^{i}}){{\rm{d}}}W_{s}^{i}+\int_{0}^{t}\sigma^{0}(X_{s}^{i},\mathcal{L}^{1}_{X_{s}^{i}}){{\rm{d}}}W^{0}_{s}\\ &+\int_{0}^{t}\int_{\mathbb{R}_{0}^{d}}c\left(X_{s-}^{i},\mathcal{L}^{1}_{X_{s-}^{i}}\right)z\widetilde{N}^{i}({{\rm{d}}}s,{{\rm{d}}}z),\end{split}$ (3.1)
for any $t\geq 0$ and $i\in\{1,\ldots,N\}$.
For $\bm{x}^{N}:=(x_{1},x_{2},\ldots,x_{N}),\bm{y}^{N}:=(y_{1},y_{2},\ldots,y_{N})\in\mathbb{R}^{dN}$, we have
$\mathcal{W}_{2}^{2}(\mu^{\bm{x}^{N}},\delta_{0})=\frac{1}{N}\sum_{i=1}^{N}|x_{i}|^{2}.$
Here, the empirical measure is defined by
$\mu^{\bm{x}^{N}}({\rm{d}}x):=\frac{1}{N}\sum_{i=1}^{N}\delta_{x_{i}}({\rm{d}}x),$
where $\delta_{x}$ denotes the Dirac measure concentrated at $x$. Moreover, a standard estimate for the Wasserstein distance between two empirical measures $\mu^{\bm{x}^{N}}$ and $\mu^{\bm{y}^{N}}$ is given by
$\mathcal{W}_{2}^{2}(\mu^{\bm{x}^{N}},\mu^{\bm{y}^{N}})\leq\frac{1}{N}\sum_{i=1}^{N}|x_{i}-y_{i}|^{2}=\frac{1}{N}\left|\bm{x}^{N}-\bm{y}^{N}\right|^{2},$
(see inequality (1.24) in [18]). The true measure $\mathcal{L}_{X_{t}}$ at each time $t$ is approximated by the empirical measure
$\mu_{t}^{\bm{X}^{N}}({\rm{d}}x):=\frac{1}{N}\sum_{i=1}^{N}\delta_{X_{t}^{i,N}}({\rm{d}}x),$
where $\bm{X}^{N}=(\bm{X}_{t}^{N})_{t\geq 0}=(X_{t}^{1,N},\ldots,X_{t}^{N,N})_{t\geq 0}^{\top}$ denotes the system of interacting particles, which is the solution to an $\mathbb{R}^{dN}$-valued SDE driven by Lévy processes with common noise. The components $X^{i,N}=(X_{t}^{i,N})_{t\geq 0}$ correspond to the state of the $i$-th particle. Then, we have
$\begin{split}X_{t}^{i,N}&=x_{0}+\int_{0}^{t}b(X_{s}^{i,N},\mu_{s}^{\bm{X}^{N}}){{\rm{d}}}s+\int_{0}^{t}\sigma(X_{s}^{i,N},\mu_{s}^{\bm{X}^{N}}){{\rm{d}}}W_{s}^{i}+\int_{0}^{t}\sigma^{0}(X_{s}^{i,N},\mu_{s}^{\bm{X}^{N}}){{\rm{d}}}W^{0}_{s}\\ &\quad+\int_{0}^{t}\int_{\mathbb{R}_{0}^{d}}c\left(X_{s-}^{i,N},\mu_{s-}^{\bm{X}^{N}}\right)z\widetilde{N}^{i}({\rm{d}}s,{\rm{d}}z),\end{split}$ (3.2)
[…]
$\max_{i\in\{1,\dots,N\}}\sup_{t\in[0,T]}\mathbb{E}\big{[}|X^{i,N}_{t}|^{2}\big{]}\leq K,$
for any $N\in\mathbb{N}$, where the constant $K>0$ is independent of $N$.
Proposition 3.1.
(see [41]) For each $p\in[2,p_{0}]$, there exists a constant $C_{p}>0$ such that, for all $t\geq 0$,
$\max_{1\leq i\leq N}\mathbb{E}\!\left[|X_{t}^{i,N}|^{p}\right]\leq\begin{cases}C_{p}\big{(}1+e^{\gamma pt}\big{)},&\gamma\neq 0,\\[3.00003pt] C_{p}(1+t)^{p/2},&\gamma=0,\ p=2\ \text{or}\ (\gamma=0,\ \gamma_{2}>0,\ p\in(2,p_{0}]),\\[3.00003pt] C_{p}(1+t)^{p},&\gamma=0,\ \gamma_{2}=0,\ p\in(2,p_{0}],\end{cases}$
where $\gamma:=\gamma_{1}+\gamma_{2}$. In the special case $\gamma<0$, one further has
$\max_{1\leq i\leq N}\ \sup_{t\geq 0}\ \mathbb{E}\!\left[\,|X_{t}^{i,N}|^{p}\,\right]\leq 2C_{p}.$
Proof.
The argument follows the same reasoning as in the proof of Proposition 2.3 and is therefore not repeated here. ∎
Next, we state a result concerning the propagation of chaos, which plays a central role in the convergence analysis as $N\to\infty$. For clarity, we first introduce the following definition,
$\varphi(N)=\begin{cases}N^{-1/2}&\text{if }d<4,\\ N^{-1/2}\ln N&\text{if }d=4,\\ N^{-2/d}&\text{if }d>4,\end{cases}$ (3.3)
(see Theorem 1 of [21]).
Proposition 3.2.
Assume that all conditions in Proposition 3.1 hold and that Assumption (A2) holds for $L_{1}\in\mathbb{R}$, $L_{2}\geq 0$. Then, we have
$\max_{i\in\{1,\ldots,N\}}\sup_{t\in[0,T]}\mathbb{E}\left[|X_{t}^{i}-X_{t}^{i,N}|^{2}\right]\leq C_{T}\varphi(N),$ (3.4)
for any $N\in\mathbb{N}$, where the positive constant $C_{T}$ does not depend on $N$.
Assume further that $L_{1}+L_{2}<0$ and $\gamma=\gamma_{1}+\gamma_{2}<0$. Then, we have
$\max_{i\in\{1,\ldots,N\}}\sup_{t\geq 0}\mathbb{E}\left[|X_{t}^{i}-X_{t}^{i,N}|^{2}\right]\leq C\varphi(N),$ (3.5)
for any $N\in\mathbb{N}$, where the positive constant $C$ does not depend on $N$ and $T$.
Proof.
The proof of the theorem follows the classical approach presented in [41]. It is important to emphasize that the main novelty of our result is the extension of the estimate to the setting with common noise. Observe that for any $t\geq 0$,
$X_{t}^{i}-X_{t}^{i,N}$ $=\int_{0}^{t}\left(b(X_{s}^{i},\mathcal{L}^{1}_{X_{s}^{i}})-b(X_{s}^{i,N},\mu_{s}^{\bm{X}^{N}})\right){{\rm{d}}}s+\int_{0}^{t}\left(\sigma(X_{s}^{i},\mathcal{L}^{1}_{X_{s}^{i}})-\sigma(X_{s}^{i,N},\mu_{s}^{\bm{X}^{N}})\right){\rm{d}}W_{s}^{i}$
$\quad+\int_{0}^{t}\left(\sigma^{0}(X_{s}^{i},\mathcal{L}^{1}_{X_{s}^{i}})-\sigma^{0}(X_{s}^{i,N},\mu_{s}^{\bm{X}^{N}})\right){\rm{d}}W_{s}^{0}$
$\quad+\int_{0}^{t}\int_{\mathbb{R}_{0}^{d}}\left(c(X_{s-}^{i},\mathcal{L}^{1}_{X_{s-}^{i}})-c(X_{s-}^{i,N},\mu_{s-}^{\bm{X}^{N}})\right)z\widetilde{N}^{i}({\rm{d}}s,{\rm{d}}z).$
Next, for $\lambda\in\mathbb{R}$, by applying Itô’s formula and invoking Assumption (A2), $L_{1}\in\mathbb{R}$, and $L_{2}\geq 0$, we deduce that for all $t\geq 0$,
$e^{-\lambda t}|X_{t}^{i}-X_{t}^{i,N}|^{2}=\int_{0}^{t}e^{-\lambda s}\Big{[}-\lambda|X_{s}^{i}-X_{s}^{i,N}|^{2}+2\langle X_{s}^{i}-X_{s}^{i,N},b(X_{s}^{i},\mathcal{L}^{1}_{X_{s}^{i}})-b(X_{s}^{i,N},\mu_{s}^{\bm{X}^{N}})\rangle$
$\quad+|\sigma(X_{s}^{i},\mathcal{L}^{1}_{X_{s}^{i}})-\sigma(X_{s}^{i,N},\mu_{s}^{\bm{X}^{N}})|^{2}+|\sigma^{0}(X_{s}^{i},\mathcal{L}^{1}_{X_{s}^{i}})-\sigma^{0}(X_{s}^{i,N},\mu_{s}^{\bm{X}^{N}})|^{2}\Big{]}{\rm{d}}s$
$\quad+2\int_{0}^{t}e^{-\lambda s}\bigg{[}\langle X_{s}^{i}-X_{s}^{i,N},(\sigma(X_{s}^{i},\mathcal{L}^{1}_{X_{s}^{i}})-\sigma(X_{s}^{i,N},\mu_{s}^{\bm{X}^{N}})){\rm{d}}W_{s}^{i}\rangle$
$\quad+\langle X_{s}^{i}-X_{s}^{i,N},(\sigma^{0}(X_{s}^{i},\mathcal{L}^{1}_{X_{s}^{i}})-\sigma^{0}(X_{s}^{i,N},\mu_{s}^{\bm{X}^{N}})){\rm{d}}W_{s}^{0}\rangle\bigg{]}$
$\quad+\int_{0}^{t}\int_{\mathbb{R}_{0}^{d}}e^{-\lambda s}\left(|X_{s-}^{i}-X_{s-}^{i,N}+(c(X_{s-}^{i},\mathcal{L}^{1}_{X_{s-}^{i}})-c(X_{s-}^{i,N},\mu_{s-}^{\bm{X}^{N}}))z|^{2}-|X_{s-}^{i}-X_{s-}^{i,N}|^{2}\right)\widetilde{N}^{i}({\rm{d}}s,{\rm{d}}z)$
$\quad+\int_{0}^{t}\int_{\mathbb{R}_{0}^{d}}e^{-\lambda s}|(c(X_{s-}^{i},\mathcal{L}^{1}_{X_{s-}^{i}})-c(X_{s-}^{i,N},\mu_{s-}^{\bm{X}^{N}}))z|^{2}\nu({\rm{d}}z){\rm{d}}s$
$\leq\int_{0}^{t}e^{-\lambda s}\left(-\lambda|X_{s}^{i}-X_{s}^{i,N}|^{2}+2\left\langle X_{s}^{i}-X_{s}^{i,N},b(X_{s}^{i},\mathcal{L}^{1}_{X_{s}^{i}})-b(X_{s}^{i,N},\mu_{s}^{X^{N}})\right\rangle\right){\rm{d}}s$
$\quad+\int_{0}^{t}e^{-\lambda s}\left(|\sigma(X_{s}^{i},\mathcal{L}^{1}_{X_{s}^{i}})-\sigma(X_{s}^{i,N},\mu_{s}^{X^{N}})|^{2}+|\sigma^{0}(X_{s}^{i},\mathcal{L}^{1}_{X_{s}^{i}})-\sigma^{0}(X_{s}^{i,N},\mu_{s}^{X^{N}})|^{2}\right){\rm{d}}s$
$\quad+\int_{0}^{t}e^{-\lambda s}|c(X_{s-}^{i},\mathcal{L}^{1}_{X_{s-}^{i}})-c(X_{s-}^{i,N},\mu_{s-}^{X^{N}})|^{2}\int_{\mathbb{R}_{0}^{d}}|z|^{2}\nu({\rm{d}}z){\rm{d}}s$
$\quad+2\int_{0}^{t}e^{-\lambda s}\bigg{[}\left\langle X_{s}^{i}-X_{s}^{i,N},\left(\sigma(X_{s}^{i},\mathcal{L}^{1}_{X_{s}^{i}})-\sigma(X_{s}^{i,N},\mu_{s}^{X^{N}})\right){\rm{d}}W_{s}^{i}\right\rangle$
$\quad+\left\langle X_{s}^{i}-X_{s}^{i,N},\left(\sigma^{0}(X_{s}^{i},\mathcal{L}^{1}_{X_{s}^{i}})-\sigma^{0}(X_{s}^{i,N},\mu_{s}^{X^{N}})\right){\rm{d}}W_{s}^{0}\right\rangle\bigg{]}$
$\quad+\int_{0}^{t}\int_{\mathbb{R}_{0}^{d}}e^{-\lambda s}\left(|X_{s-}^{i}-X_{s-}^{i,N}+\left(c(X_{s-}^{i},\mathcal{L}^{1}_{X_{s-}^{i}})-c(X_{s-}^{i,N},\mu_{s-}^{X^{N}})\right)z|^{2}-|X_{s-}^{i}-X_{s-}^{i,N}|^{2}\right)\widetilde{N}^{i}({\rm{d}}s,{\rm{d}}z)$
$\leq\int_{0}^{t}e^{-\lambda s}\left(-\lambda|X_{s}^{i}-X_{s}^{i,N}|^{2}+L_{1}|X_{s}^{i}-X_{s}^{i,N}|^{2}+L_{2}\mathcal{W}_{2}^{2}(\mathcal{L}^{1}_{X_{s}^{i}},\mu_{s}^{X^{N}})\right){\rm{d}}s$
$\quad+2\int_{0}^{t}e^{-\lambda s}\bigg{[}\left\langle X_{s}^{i}-X_{s}^{i,N},\left(\sigma(X_{s}^{i},\mathcal{L}^{1}_{X_{s}^{i}})-\sigma(X_{s}^{i,N},\mu_{s}^{X^{N}})\right){\rm{d}}W_{s}^{i}\right\rangle$
$\quad+\left\langle X_{s}^{i}-X_{s}^{i,N},\left(\sigma^{0}(X_{s}^{i},\mathcal{L}^{1}_{X_{s}^{i}})-\sigma^{0}(X_{s}^{i,N},\mu_{s}^{X^{N}})\right){\rm{d}}W_{s}^{0}\right\rangle\bigg{]}$
$\quad+\int_{0}^{t}\int_{\mathbb{R}_{0}^{d}}e^{-\lambda s}\left(|X_{s-}^{i}-X_{s-}^{i,N}+\left(c(X_{s-}^{i},\mathcal{L}^{1}_{X_{s-}^{i}})-c(X_{s-}^{i,N},\mu_{s-}^{X^{N}})\right)z|^{2}-|X_{s-}^{i}-X_{s-}^{i,N}|^{2}\right)\widetilde{N}^{i}({\rm{d}}s,{\rm{d}}z).$
By taking the expectation and utilizing the bound
$\mathbb{E}\left[\mathcal{W}_{2}^{2}\big{(}\mathcal{L}^{1}_{X_{s}^{i}},\mathcal{L}^{1}_{X_{s}^{i,N}}\big{)}\right]\leq\mathbb{E}\left[|X_{s}^{i}-X_{s}^{i,N}|^{2}\right],$
it follows that, for any $\varepsilon>0$,
$\begin{split}&e^{-\lambda t}\mathbb{E}\left[|X_{t}^{i}-X_{t}^{i,N}|^{2}\right]\\ &\leq\int_{0}^{t}e^{-\lambda s}\bigg{(}(-\lambda+L_{1})\mathbb{E}\left[|X_{s}^{i}-X_{s}^{i,N}|^{2}\right]+L_{2}\mathbb{E}\left[\mathcal{W}_{2}^{2}(\mathcal{L}^{1}_{X_{s}^{i}},\mu_{s}^{\bm{X}^{N}})\right]\bigg{)}{\rm{d}}s\\ &\leq\int_{0}^{t}e^{-\lambda s}\left((-\lambda+L_{1}+L_{2}+\varepsilon)\mathbb{E}\left[|X_{s}^{i}-X_{s}^{i,N}|^{2}\right]+L_{2}\left(1+\frac{L_{2}}{\varepsilon}\right)\mathbb{E}\left[\mathcal{W}_{2}^{2}(\mathcal{L}^{1}_{X_{s}^{i,N}},\mu_{s}^{\bm{X}^{N}})\right]\right){\rm{d}}s.\end{split}$ (3.6)
Moreover, by Proposition 3.1, for any $p\in(4,p_{0}]$, we have
$\max_{i\in\{1,\dots,N\}}\sup_{t\in[0,T]}\mathbb{E}\left[\left|X^{i,N}_{t}\right|^{p}\right]\leq C_{T},$
for some constant $C_{T}>0$. Following Fournier and Guillin [21] and Sznitman [38], for any $\mu\in\mathcal{P}_{q}(\mathbb{R}^{d})$ with $q>4$, there exists a constant $C=C(d,q,M_{q}(\mu))$ such that the empirical measure $\mu^{N}$ satisfies
$\mathbb{E}\left[\mathcal{W}_{2}^{2}(\mu^{N},\mu)\right]\leq C\varphi(N),$
where $\varphi(N)$ is defined in (3.3). Although the original result in Fournier and Guillin [21] assumes independent particles, this bound still holds for our system with common noise $W^{0}$ due to the exchangeability of the particles conditioned on $\mathcal{F}^{0}$. The rest of the proof proceeds similarly by controlling the term $\mathbb{E}\left[\mathcal{W}_{2}^{2}\big{(}\mathcal{L}^{1}_{X^{i,N}_{s}},\mu^{X^{N}}_{s}\big{)}\right]$ using the triangle inequality and the exchangeability of the particles. If $\mu\in\mathcal{P}_{q}(\mathbb{R}^{d})$ for some $q>4$, that is, if
$M_{q}(\mu)^{q}:=\int_{\mathbb{R}^{d}}|x|^{q}\,\mu({\rm{d}}x)<\infty,$
then for each dimension $d\geq 1$, there exists a constant $C=C(d,q,M_{q}(\mu))$ such that, for all $N\geq 2$,
$\begin{split}\mathbb{E}\left[\mathcal{W}_{2}(\bar{\mu}^{N},\mu)^{2}\right]&\leq C\begin{cases}N^{-1/2}&\text{if }d<4,\\ N^{-1/2}\log N&\text{if }d=4,\\ N^{-2/d}&\text{if }d>4,\end{cases}\\ &=C\varphi(N).\end{split}$
Next, let $(X_{t}^{i,N})_{i=1}^{N}$ be a collection of random variables that are identically distributed but not independent, and define the empirical measure by
$\mu_{t}^{\bm{x}^{N}}:=\frac{1}{N}\sum_{i=1}^{N}\delta_{x_{t}^{i,N}}.$
Meanwhile, let $(X_{t}^{i})_{i=1}^{N}$ be a sequence of independent and identically distributed random variables, and define the corresponding empirical measure by
$\bar{\mu}_{t}^{N}:=\frac{1}{N}\sum_{i=1}^{N}\delta_{x_{t}^{i}}.$
In this case, we have
[…]
Consequently, it suffices to choose $\lambda=L_{1}+L_{2}+\epsilon$ in (LABEL:3.6) to conclude that
$\begin{split}&e^{-\lambda t}\mathbb{E}\left[|X_{t}^{i}-X_{t}^{i,N}|^{2}\right]\\ &\leq\int_{0}^{t}e^{-\lambda s}\left(L_{2}\left(1+\frac{L_{2}}{\varepsilon}\right)\mathbb{E}\left[\mathcal{W}_{2}^{2}(\mathcal{L}^{1}_{X_{s}^{i,N}},\mu_{s}^{\bm{X}^{N}})\right]\right){{\rm{d}}}s\\ &\leq L_{2}\left(1+\frac{L_{2}}{\varepsilon}\right)\int_{0}^{t}e^{-\lambda s}\left(6\mathbb{E}[|X_{s}^{i}-X_{s}^{i,N}|^{2}]+3C\varphi(N)\right){{{\rm{d}}}}s\\ &=6L_{2}\left(1+\frac{L_{2}}{\varepsilon}\right)\int_{0}^{t}e^{-\lambda s}\mathbb{E}[|X_{s}^{i}-X_{s}^{i,N}|^{2}]{{{\rm{d}}}}s+3C\varphi(N)L_{2}\left(1+\frac{L_{2}}{\varepsilon}\right)\int_{0}^{t}e^{-\lambda s}{{\rm{d}}}s.\end{split}$
Then using Gronwall’s inequality, we have
[…]
When $L_{1}+L_{2}<0$ and $\gamma=\gamma_{1}+\gamma_{2}<0$, one can choose $\epsilon>0$ sufficiently small so that $\lambda<0$. Hence, we have
$\begin{split}\mathbb{E}\left[|X_{t}^{i}-X_{t}^{i,N}|^{2}\right]&\leq\frac{3C\varphi(N)L_{2}\left(1+\frac{L_{2}}{\varepsilon}\right)}{\lambda}e^{-6L_{2}\left(1+\frac{L_{2}}{\varepsilon}\right)(1-e^{\lambda t})/\lambda}\\ &\leq\frac{3C\varphi(N)L_{2}\left(1+\frac{L_{2}}{\varepsilon}\right)}{\lambda}e^{-6L_{2}\left(1+\frac{L_{2}}{\varepsilon}\right)/\lambda}\\ &=C\varphi(N),\end{split}$
for any $t\in[0,T]$, where the positive constant $C$ does not depend on time. Finally, when $\gamma<0$, it follows from Proposition 3.1 that for any $p\in(4,p_{0}]$,
$\max_{i\in\{1,\dots,N\}}\sup_{t\geq 0}\mathbb{E}\left[\left|X^{i,N}_{t}\right|^{q}\right]\leq C,$
where the positive constant $C$ does not depend on $T$. The desired proof follows. ∎
4 Tamed-adaptive Euler-Maruyama method
4.1 Tamed coefficients and adaptive time control
For each $\Delta\in(0,1)$ define the tamed coefficients
$\sigma_{\Delta}(x,\mu)$ $:=\frac{\sigma(x,\mu)}{1+\Delta^{1/2}|\sigma(x,\mu)|(1+|x|)},$ (4.1)
$c_{\Delta}(x,\mu)$ $:=\frac{c(x,\mu)}{1+\Delta^{1/2}|c(x,\mu)|(1+|x|+|b(x,\mu)|)},$ (4.2)
$\sigma^{0}_{\Delta}(x,\mu)$ $:=\frac{\sigma^{0}(x,\mu)}{1+\Delta^{1/2}|\sigma^{0}(x,\mu)|(1+|x|)}.$ (4.3)
The adaptive time control function is taken as
$h(x,\mu)=\frac{h_{0}}{(1+|b(x,\mu)|+|\sigma(x,\mu)|+|x|^{l})^{2}+|c(x,\mu)|^{p_{0}}},$ (4.4)
with $h_{0}>0$ (we set $h_{0}=1$ in proofs), and $l,p_{0}$ as in (A3) and (A4). We will use the standard taming bounds
$|\sigma_{\Delta}(x,\mu)|(1+|x|)\leq\Delta^{-1/2},$ (4.5)
$|c_{\Delta}(x,\mu)|(1+|x|)\leq\Delta^{-1/2},$ (4.6)
$|\sigma(x,\mu)-\sigma_{\Delta}(x,\mu)|\leq C\sqrt{\Delta}\,|\sigma(x,\mu)|^{2}(1+|x|),$ (4.7)
$|\sigma^{0}(x,\mu)-\sigma^{0}_{\Delta}(x,\mu)|\leq C\sqrt{\Delta}\,|\sigma^{0}(x,\mu)|^{2}(1+|x|).$ (4.8)
4.2 Continuous-time tamed adaptive scheme with common noise
Fix $N\in\mathbb{N}$. For $i=1,\dots,N$, let $W^{i}$ be independent Brownian motions, $W^{0}$ a common Brownian motion, and $Z^{i}$ independent Lévy processes with common Lévy measure $\nu$. Define the empirical measure $\mu_{\widehat{X}^{N}_{t}}:=\frac{1}{N}\sum_{j=1}^{N}\delta_{\widehat{X}^{j,N}_{t}}$.
The continuous-time tamed adaptive Euler-Maruyama approximation $\widehat{X}^{i,N}_{t}$ is the Itô process
$\widehat{X}^{i,N}_{t}$ $=x_{0}+\int_{0}^{t}b\bigl{(}\widehat{X}^{i,N}_{s},\mu_{\widehat{X}^{N}_{s}}\bigr{)}\,{\rm{d}}s+\int_{0}^{t}\sigma_{\Delta}\!\bigl{(}\widehat{X}^{i,N}_{s},\mu_{\widehat{X}^{N}_{s}}\bigr{)}\,{\rm{d}}W^{i}_{s}$
$\quad+\int_{0}^{t}\sigma^{0}_{\Delta}\!\bigl{(}\widehat{X}^{i,N}_{s},\mu_{\widehat{X}^{N}_{s}}\bigr{)}\,{\rm{d}}W^{0}_{s}+\int_{0}^{t}\int_{\mathbb{R}^{d}_{0}}c_{\Delta}\!\bigl{(}\widehat{X}^{i,N}_{s-},\mu_{\widehat{X}^{N}_{s-}}\bigr{)}z\,\widetilde{N}^{i}({\rm{d}}s,{\rm{d}}z),$ (4.9)
where $\widetilde{N}^{i}$ is the compensated Poisson random measure of $Z^{i}$.
Remark 4.1.
[…]
$\max_{1\leq i\leq N}\mathbb{E}\Big{[}\sup_{0\leq t\leq T}|\widehat{X}^{i,N}_{t}|^{p}\Big{]}\leq C.$
Proof.
The proof of this remark can be established by adapting the methodology presented in [41]. ∎
Theorem 4.2.
Assume (A1)-(A7) and use the definitions (4.1)-(4.4). Let $X^{i,N}$ be the interacting particle system (4.2) and $\widehat{X}^{i,N}$ the continuous tamed-adaptive approximation (4.2). Then for any $T>0$ there exists $C_{T}>0$ such that
$\max_{1\leq i\leq N}\sup_{0\leq t\leq T}\mathbb{E}\bigl{[}|X^{i,N}_{t}-\widehat{X}^{i,N}_{t}|^{2}\bigr{]}\leq C_{T}\Delta.$
If moreover $L_{1}+L_{2}<0$ and $\gamma<0$, the constant can be chosen independent of $T$, i.e.
$\max_{1\leq i\leq N}\sup_{t\geq 0}\mathbb{E}\bigl{[}|X^{i,N}_{t}-\widehat{X}^{i,N}_{t}|^{2}\bigr{]}\leq C\Delta.$
Proof.
We denote by $X^{i,N}$ the exact interacting particle system with common noise
$X^{i,N}_{t}$ $=x_{0}+\int_{0}^{t}b\bigl{(}X^{i,N}_{s},\mu_{X^{N}_{s}}\bigr{)}\,{\rm{d}}s+\int_{0}^{t}\sigma\bigl{(}X^{i,N}_{s},\mu_{X^{N}_{s}}\bigr{)}\,{\rm{d}}W^{i}_{s}$
$\quad+\int_{0}^{t}\sigma^{0}\bigl{(}X^{i,N}_{s},\mu_{X^{N}_{s}}\bigr{)}\,{\rm{d}}W^{0}_{s}+\int_{0}^{t}\int_{\mathbb{R}^{d}_{0}}c\bigl{(}X^{i,N}_{s-},\mu_{X^{N}_{s-}}\bigr{)}z\,\widetilde{N}^{i}({\rm{d}}s,{\rm{d}}z).$ (4.10)
Set the error $e^{i,N}_{t}:=X^{i,N}_{t}-\widehat{X}^{i,N}_{t}$. Applying Itô’s formula to $|e^{i,N}_{t}|^{2}$ yields
$|e^{i,N}_{t}|^{2}$ $=2\int_{0}^{t}\big{\langle}e^{i,N}_{s},\,b(X^{i,N}_{s},\mu_{X^{N}_{s}})-b(\widehat{X}^{i,N}_{s},\mu_{\widehat{X}^{N}_{s}})\big{\rangle}{\rm{d}}s$
$\quad+\int_{0}^{t}\big{\|}\sigma(X^{i,N}_{s},\mu_{X^{N}_{s}})-\sigma_{\Delta}(\widehat{X}^{i,N}_{s},\mu_{\widehat{X}^{N}_{s}})\big{\|}^{2}{\rm{d}}s$
$\quad+\int_{0}^{t}\big{\|}\sigma^{0}(X^{i,N}_{s},\mu_{X^{N}_{s}})-\sigma^{0}_{\Delta}(\widehat{X}^{i,N}_{s},\mu_{\widehat{X}^{N}_{s}})\big{\|}^{2}{\rm{d}}s$
$\quad+\int_{0}^{t}\int_{\mathbb{R}^{d}_{0}}\big{|}c(X^{i,N}_{s-},\mu_{X^{N}_{s-}})-c_{\Delta}(\widehat{X}^{i,N}_{s-},\mu_{\widehat{X}^{N}_{s-}})\big{|}^{2}|z|^{2}\nu(dz)\,{\rm{d}}s$
$\quad+\mathcal{M}^{i,N}_{t},$ (4.11)
where $\mathcal{M}^{i,N}_{t}$ is a local martingale collecting the stochastic integral remainders.
Decompose each coefficient difference into a consistency part and a taming part, e.g.
$).\sigma(X,\mu)-\sigma_{\Delta}(\widehat{X},\widehat{\mu})=\big{(}\sigma(X,\mu)-\sigma(\widehat{X},\widehat{\mu})\big{)}+\big{(}\sigma(\widehat{X},\widehat{\mu})-\sigma_{\Delta}(\widehat{X},\widehat{\mu}$$\big{)}$.
Using Assumptions (A2)-(A3) and the taming control (4.8), for any small $\varepsilon>0$ we can bound
$\big{\|}\sigma(X,\mu)-\sigma_{\Delta}(\widehat{X},\widehat{\mu})\big{\|}^{2}$ $\leq(1+\varepsilon)\big{\|}\sigma(X,\mu)-\sigma(\widehat{X},\widehat{\mu})\big{\|}^{2}+C(\varepsilon)\Delta(1+|\widehat{X}|^{q_{1}}),$ (4.12)
$\big{\|}\sigma^{0}(X,\mu)-\sigma^{0}_{\Delta}(\widehat{X},\widehat{\mu})\big{\|}^{2}$ $\leq(1+\varepsilon)\big{\|}\sigma^{0}(X,\mu)-\sigma^{0}(\widehat{X},\widehat{\mu})\big{\|}^{2}+C(\varepsilon)\Delta(1+|\widehat{X}|^{q_{2}}),$ (4.13)
with analogous estimates for $c$; here $q_{1},q_{2}$ reflect coefficient growth and are handled by moment bounds.
Plugging these decompositions into (4.2), using (A2), Young’s and BDG’s inequalities, yields for suitable constants $C_{1},C_{2}$ and a polynomial moment term $R_{s}$, which is bounded in expectation on $[0,T]$
$|e^{i,N}_{t}|^{2}\leq\int_{0}^{t}\big{(}C_{1}|e^{i,N}_{s}|^{2}+C_{2}\mathcal{W}_{2}^{2}(\mu_{X^{N}_{s}},\mu_{\widehat{X}^{N}_{s}})\big{)}\,{\rm{d}}s+C\Delta\int_{0}^{t}(1+R_{s})\,{\rm{d}}s+\mathcal{M}^{i,N}_{t}.$ (4.14)
Take expectations in (4.14) and use the standard bound
$\mathcal{W}_{2}^{2}(\mu_{X^{N}_{s}},\mu_{\widehat{X}^{N}_{s}})\leq\frac{1}{N}\sum_{j=1}^{N}|e^{j,N}_{s}|^{2}.$
Let $\mathcal{E}_{t}:=\max_{1\leq i\leq N}\mathbb{E}[|e^{i,N}_{t}|^{2}]$. Then from (4.14) we obtain
$\mathcal{E}_{t}\leq C\int_{0}^{t}\mathcal{E}_{s}\,{\rm{d}}s+C\Delta\int_{0}^{t}(1+\mathbb{E}[R_{s}])\,{\rm{d}}s.$
By uniform moment bounds, $\mathbb{E}[R_{s}]\leq C_{T}$ on $[0,T]$. Hence
$\mathcal{E}_{t}\leq C\int_{0}^{t}\mathcal{E}_{s}\,{\rm{d}}s+C_{T}\Delta t,$
and Gronwall’s inequality gives
$\mathcal{E}_{t}\leq C_{T}\Delta,\qquad t\in[0,T].$
Therefore, we have
$\max_{1\leq i\leq N}\sup_{0\leq t\leq T}\mathbb{E}\big{[}|X^{i,N}_{t}-\widehat{X}^{i,N}_{t}|^{2}\big{]}\leq C_{T}\Delta.$
If in addition the contractivity conditions $L_{1}+L_{2}<0$ and $\gamma=\gamma_{1}+\gamma_{2}<0$ hold, then the same reasoning yields a time-uniform estimate
$\max_{1\leq i\leq N}\sup_{t\geq 0}\mathbb{E}\big{[}|X^{i,N}_{t}-\widehat{X}^{i,N}_{t}|^{2}\big{]}\leq C\Delta,$
with $C$ independent of $T$. ∎
5 Numerical Example
In this section, I will provide corresponding numerical examples based on the aforementioned theoretical background. Each of these conditional McKean-Vlasov SDEs involves both idiosyncratic and common sources of randomness, as well as jump components, and the coefficients satisfy the conditions required in (A1)-(A7).
Model 1 (LQ conditional mean-field control I).
${\rm{d}}X_{t}=u(t)\,{\rm{d}}t+\theta\,{\rm{d}}B_{t}^{0}+\sigma\,\mathbb{E}[X_{t}\mid\mathcal{F}_{t}^{B^{0}}]\,{\rm{d}}W_{t}+\int_{\mathbb{R}}\gamma_{0}(z)\,\mathbb{E}[X_{t}\mid\mathcal{F}_{t}^{B^{0}}]\,\tilde{N}({\rm{d}}t,{\rm{d}}z),\quad X_{0}=x_{0}\in\mathbb{R},$
where $B^{0}$ is a common Brownian motion, $W$ is an idiosyncratic Brownian motion, and $\tilde{N}$ is a compensated Poisson random measure. The control process $u$ is adapted to the common filtration, and the jump amplitude $\gamma_{0}(z)$ is assumed to be bounded. Under these conditions, the coefficients satisfy (A1)-(A5) and (A7). In the numerical experiments, the jump distribution is taken to be Gaussian, so that (A4) also holds.

### v2

@@
 Abstract
-In this work, we establish the existence and uniqueness of solutions to McKean-Vlasov stochastic differential equations (SDEs) driven by Lévy processes with common noise on an infinite time horizon, by means of a contraction mapping principle in the space of probability measures.
-In addition, we analyse the propagation of chaos for Lévy-driven McKean-Vlasov SDEs in the presence of common noise.
-Keywords: McKean-Vlasov equations; Lévy processes; Common noise; Moment estimates; Propagation of chaos
+In this work, we establish the existence and uniqueness of solutions to McKean-Vlasov stochastic differential equations (SDEs) driven by Lévy processes with common noise, by means of a contraction mapping principle in the space of probability measures.
+In addition, on the infinite time horizon we analyse the propagation of chaos for McKean-Vlasov SDEs driven by Lévy processes in the presence of common noise, and provide numerical simulations to validate the corresponding model.
+Keywords: McKean-Vlasov equations; Lévy processes; Common noise; Propagation of chaos; Numerical simulation
 1 Introduction
@@
 $\quad+c\bigl(X(t),\mathcal{L}^{1}(X(t))\bigr)\mathrm{d}Z_{t}+\sigma^{0}\bigl(X(t),\mathcal{L}^{1}(X(t))\bigr)\mathrm{d}W^{0}_{t},$
-for $t\geq 0$, with initial value $X(0)=X_{0}$ and its distribution $\mu=\mathcal{L}(X_{0})$.
-Here $\mathcal{L}^{1}(X_{t})$ denotes the marginal law of $X$ at time $t$, $W=(W_{t})_{t\geq 0}$ is a $d$-dimensional standard Brownian motion.
-Denote $\mathbb{R}_{0}^{d}:=\mathbb{R}^{d}\setminus\{0\}$ and let $\mathcal{B}(\mathbb{R}_{+}\times\mathbb{R}_{0}^{d})$ be the Borel $\sigma$-algebra on $\mathbb{R}_{+}\times\mathbb{R}_{0}^{d}$.
+for $t\geq 0$, with initial value $X_{0}=x_{0}$.
+Here $\mathcal{L}^{1}(X(t))$ provides a version of the conditional law of $X(t)$ given $W^{0}$, $W=(W_{t})_{t\geq 0}$ is a $d$-dimensional standard Brownian motion.
+Denote $\mathbb{R}_{0}^{d}:=\mathbb{R}^{d}\setminus\{0\}$ and let $\mathcal{B}(\mathbb{R}_{+}\times\mathbb{R}_{0}^{d})$ be the Borel $\sigma$-algebra on $\mathbb{R}_{+}\times\mathbb{R}_{0}^{d}$, and
 $Z_{t}=\int_{0}^{t}\int_{\mathbb{R}_{0}^{d}}z\widetilde{N}(\mathrm{d}s,\mathrm{d}z),\quad t\geq 0,$
-where $\widetilde{N}$ is the compensated Poisson random measure associated with the Poisson random measure $N$ on $(\mathbb{R}_{+}\times\mathbb{R}_{0}^{d},\mathcal{B}(\mathbb{R}_{+}\times\mathbb{R}_{0}^{d}))$ with intensity measure $\nu(\mathrm{d}z)\mathrm{d}s$.
-The compensated Poisson random measure is given by
-$\widetilde{N}(\mathrm{d}t,\mathrm{d}z):=N(\mathrm{d}t,\mathrm{d}z)-\nu(\mathrm{d}z)\mathrm{d}t.$
+where $\widetilde{N}(\mathrm{d}t,\mathrm{d}z):=N(\mathrm{d}t,\mathrm{d}z)-\nu(\mathrm{d}z)\mathrm{d}t$ is the compensated Poisson random measure associated with the Poisson random measure $N$ on $(\mathbb{R}_{+}\times\mathbb{R}_{0}^{d},\mathcal{B}(\mathbb{R}_{+}\times\mathbb{R}_{0}^{d}))$ with intensity measure $\nu(\mathrm{d}z)\mathrm{d}s$.
 Explicitly, for $t>0$ and $A\in\mathcal{B}(\mathbb{R}_{0}^{d})$,
@@
 $\int_{\mathbb{R}^{d}}(1\wedge|z|^{2})\nu(\mathrm{d}z)<+\infty.$
-Let $|\cdot|$ denote the Euclidean norm on $\mathbb{R}^{d}$.
-Let $(\mathcal{F}^{0}_{t})_{t\geq 0}$, $(\mathcal{F}^{1}_{t})_{t\geq 0}$ be two complete filtered probability spaces satisfying the usual condition.
+Let $|\cdot|$ denote the Euclidean norm on $\mathbb{R}^{d}$, and let $(\mathcal{F}^{0}_{t})_{t\geq 0}$, $(\mathcal{F}^{1}_{t})_{t\geq 0}$ be two complete filtered probability spaces satisfying the usual condition.
 Let $(W_{t}^{0})_{t\geq 0}$ and $(W_{t})_{t\geq 0}$ be independent $d$-dimensional Brownian motions defined on $(\Omega^{0},\mathcal{F}^{0},\mathbb{P}^{0})$ and $(\Omega^{1},\mathcal{F}^{1},\mathbb{P}^{1})$, respectively.
@@
 $\mathcal{L}^{1}(\xi):=\mathcal{L}(\xi(\omega^{0},\cdot)),$
-which can be interpreted as a mapping from $\Omega^{0}$ into $\mathcal{P}(\mathbb{R}^{d})$.
-This object represents the conditional distribution of $\xi$ given $\mathcal{F}^{0}$.
+which represents the conditional distribution of $\xi$ given $\mathcal{F}^{0}$.
 In particular, if the initial condition $X(0)$ in (1.1) is assumed to be defined on $(\Omega^{1},\mathcal{F}^{1},\mathbb{P}^{1})$, then $\mathcal{L}^{1}(X(t))$ provides a version of the conditional law of $X(t)$ given $W^{0}$.
@@
 The existence and uniqueness theory for strong solutions of McKean-Vlasov SDEs under Lipschitz and linear growth conditions (with respect to both the state variable and the measure) is well-established (see, e.g., [25], [29]).
-Wang [42] establishes well-posedness via Picard iteration method resolving pathwise intractability in the Brownian-driven case.
+Wang [33] establishes well-posedness via the Picard iteration method, resolving the pathwise intractability in the case of SDEs driven by Brownian motion.
 In the context of delay McKean-Vlasov SDEs with super-linear drift, Bao et al.
@@
 Liu et al.
-[24] establish strong and weak well-posedness of McKean-Vlasov SDEs under locally monotone coefficients in state and distribution variables, and derived a large deviation principle with applications to granular media, plasma, kinetic, porous media, and Navier-Stokes equations.
-Compared with Brownian-driven McKean-Vlasov SDEs, the McKean-Vlasov SDEs driven by Lévy processes have been much less studied.
-Bao and Wang [7] establish existence, uniqueness, and multiplicity of stationary distributions for Lévy-driven McKean-Vlasov SDEs.
+[19] establish concentration inequalities, exponential convergence in the Wasserstein metric $\mathcal{W}_{1}$, and uniform-in-time propagation of chaos for McKean-Vlasov systems via reflection coupling and new cost functions, with applications to multi-well confinement and bounded interaction potentials.
+Compared with McKean-Vlasov SDEs driven by Brownian motion, the McKean-Vlasov SDEs driven by Lévy processes have been much less studied.
+Bao and Wang [5] establish existence, uniqueness, and multiplicity of stationary distributions for McKean-Vlasov SDEs driven by Lévy processes.
 Song et al.
-[30] develop large and moderate deviation principles, along with explicit rate functions, for Lévy-driven McKean-Vlasov SDEs using a weak-convergence approach.
+[23] develop large and moderate deviation principles, along with explicit rate functions, for McKean-Vlasov SDEs driven by Lévy processes using a weak-convergence approach.
 In [21], Jourdain et al.
@@
 [28] address the well-posedness and propagation of chaos results for McKean-Vlasov SDEs with delay driven by Lévy noise.
-The Brownian-driven McKean-Vlasov SDEs with common noise have been applied to various fields including finance, physics, biology, mean-field games, and machine learning.
+Bao et al.
+[7] establish strong well-posedness of McKean-Vlasov SDEs driven by Lévy processes under weak monotonicity and coercivity, prove both weak and strong propagation of chaos, and extend the results to systems with common noise.
+The McKean-Vlasov SDEs driven by Brownian motion with common noise have been applied to various fields including finance, physics, biology, mean-field games, and machine learning.
 The fundamental theory of the McKean-Vlasov SDEs with common noise is well established.
@@
 Shao and Wei [30] study the well-posedness and conditional propagation of chaos for McKean-Vlasov SDEs with common noise, establishing explicit convergence rates.
-An epidemic outbreak occurs when a contagious disease spreads rapidly through a population due to increased transmission rates.
-In such circumstances, it is reasonable to incorporate a jump process to capture sudden surges in infection numbers triggered by super-spreader events.
-Based on this perspective, the evolution of the infected population can be modelled by a jump diffusion.
-Moreover, when global mobility is high, localized outbreaks can quickly propagate across regions, resulting in a pandemic (i.e., a worldwide health crisis in which many areas experience simultaneous infection waves).
-This observation shows that epidemic dynamics are subject to system-wide randomness (e.g., environmental factors, global travel) that impact all communities.
-In deep learning, training dynamics can be modeled by Lévy-driven McKean-Vlasov SDEs, where parameter updates follow distribution-dependent stochastic evolution with heavy-tailed noise.
-Mini-batch gradients and restarts induce jump-like perturbations, while the mean-field term captures interactions among parameters in wide-network limits.
-In federated or distributed settings, common randomness (e.g., global schedules, network delays) produces correlated dynamics across learners.
-This framework helps explain heavy-tailed gradient noise, escape from saddles via jumps, and inspires neural SDE architectures that combine diffusion with jumps for expressive modeling.
-The above considerations motivate the study of Lévy-driven McKean-Vlasov SDEs with common noise [[9],[10]].
-In recent years, significant advances have been achieved for Lévy-driven McKean-Vlasov SDEs under common noise, covering topics such as well-posedness, deep learning approaches, optimal stopping problems, optimal and impulse control, as well as stochastic maximum principles; see, for instance, [[1], [2], [3], [9], [10], [22], [23]] for further references.
-This paper investigates the infinite time horizon approximation for Lévy-driven McKean-Vlasov SDEs with common noise.
-The primary contribution of this study lies in addressing a research gap by extending previous work on McKean-Vlasov SDE with Lévy-driven process, where the drift and diffusion coefficients are not globally Lipschitz continuous.
+In applications, random jump shocks naturally arise in a variety of systems.
+For instance, an epidemic outbreak occurs when a contagious disease spreads rapidly through a population due to increased transmission rates.
+Super-spreader events may cause sudden surges in infection numbers, which can be modelled by a jump component in a diffusion-type dynamics for the infected population.
+When global mobility is high, localized outbreaks can quickly propagate across regions, resulting in a pandemic, and the epidemic dynamics are then subject to system-wide sources of randomness (e.g., environmental factors, global travel) that impact all communities.
+Similar large-scale effects appear in economics and finance.
+In systemic risk modelling, the balance sheet of each bank is influenced not only by its idiosyncratic shocks but also by macroeconomic factors and market-wide events.
+This can be described by McKean-Vlasov SDEs with common noise and jumps, where the mean-field term captures peer effects and the jump component models sudden losses or liquidity dry-ups; see the interbank lending model studied later.
+Bistable dynamics with random environment, as in a Ginzburg-Landau type equation, provide a prototype for metastable transitions under heavy-tailed shocks, and will serve as our first test model.
+In conditional mean-field control problems, agents interact through conditional expectations of their states while being exposed to common noise representing shared information or policy shocks.
+Linear-quadratic structures then lead to tractable feedback controls and motivate the LQ test models considered in Section 5.
+These examples motivate the study of McKean-Vlasov SDEs driven by Lévy processes with common noise; see, for instance, [8, 9].
+In recent years, significant advances have been achieved for McKean-Vlasov SDEs driven by Lévy processes under common noise, covering topics such as well-posedness, deep learning approaches, optimal stopping problems, optimal and impulse control, as well as stochastic maximum principles; see, among others, [1, 2, 3, 8, 9, 13, 15, 17, 18, 24] for further references.
+This paper investigates the infinite time horizon approximation for McKean-Vlasov SDEs driven by Lévy processes with common noise.
+The primary contribution of this study lies in addressing a research gap by extending previous work to McKean-Vlasov SDE driven by Lévy processes, where the drift and diffusion coefficients are not globally Lipschitz continuous.
 Tran et al.
-[41] propose a tamed-adaptive Euler-Maruyama scheme for Lévy-driven McKean-Vlasov SDEs with superlinear and non-globally Lipschitz coefficients, proving strong convergence on finite and infinite horizons.
+[32] propose a tamed-adaptive Euler-Maruyama scheme for McKean-Vlasov SDEs driven by Lévy processes with superlinear coefficients, proving strong convergence on finite and infinite horizons.
 In particular, this study builds upon the results of Tran et al.
-[41], introducing the effect of common noise-an important factor that has not been examined in the context of such SDEs.
+[32] by introducing the effect of common noise.
+This factor has not been examined in the context of such SDEs.
 The incorporation of common noise enhances the model, providing a more realistic framework for situations where noise simultaneously impacts all components, a consideration that has often been neglected in prior research.
@@
   2. 2.
-    We extend the results of Tran et al. [41] by establishing the boundedness of high-order moments of the solutions and demonstrating the propagation of chaos in the presence of common noise.
+    We extend the results of Tran et al. [32] to establish the boundedness of high-order moments of the solutions and demonstrate the propagation of chaos in the presence of common noise.
   3. 3.
@@
 for any $x,\overline{x}\in\mathbb{R}^{d}$ and $\mu,\overline{\mu}\in\mathcal{P}_{2}(\mathbb{R}^{d})$.
-(A3) There exist constants $L>0$ and $\ell\geq 0$ such that
-$|b(x,\mu)-b(\overline{x},\overline{\mu})|\leq L(1+|x|^{\ell}+|\overline{x}|^{\ell})\left(|x-\overline{x}|+\mathcal{W}_{2}(\mu,\overline{\mu})\right),$
+(A3) There exist constants $\tilde{L}>0$ and $\ell>0$ such that
+$|b(x,\mu)-b(\overline{x},\overline{\mu})|\leq\tilde{L}(1+|x|^{\ell}+|\overline{x}|^{\ell})\left(|x-\overline{x}|+\mathcal{W}_{2}(\mu,\overline{\mu})\right),$
 for any $x,\overline{x}\in\mathbb{R}^{d}$ and $\mu,\overline{\mu}\in\mathcal{P}_{2}(\mathbb{R}^{d})$.
@@
 $\leq\gamma_{1}|x|^{2}+\gamma_{2}\mathcal{W}_{2}^{2}(\mu,\delta_{0})+\eta,$
-for any $x\in\mathbb{R}^{d}$ and $\mu\in\mathcal{P}_{2}(\mathbb{R}^{d})$, where $\gamma=\gamma_{1}+\gamma_{2}$ represents the dissipation rate.
+for any $x\in\mathbb{R}^{d}$ and $\mu\in\mathcal{P}_{2}(\mathbb{R}^{d})$.
 (A7) There exist constants $\tilde{L}_{1}\in\mathbb{R}$ and $\tilde{L}_{2}\geq 0$ such that
@@
 for any $x,\overline{x}\in\mathbb{R}^{d}$ and $\mu,\overline{\mu}\in\mathcal{P}_{2}(\mathbb{R}^{d})$.
+Remark 2.1.
+Similar with Tran et al.
+[32].
+It follows from Assumption (A6) that, for the even integer $p_{0}\in[2,+\infty)$ given in (A4) and the positive constant $L_{3}$ given in (A5), there exist constants $\gamma_{1}\in\mathbb{R}$, $\gamma_{2}\geq 0$, and $\eta\geq 0$ such that
+$\langle x,b(x,\mu)\rangle+\frac{p-1}{2}(|\sigma(x,\mu)|^{2}+|\sigma^{0}(x,\mu)|^{2})+\frac{1}{2L_{3}}|c(x,\mu)|^{2}\int_{\mathbb{R}^{d}_{0}}|z|\left(\left(1+L_{3}|z|\right)^{p-1}-1\right)\nu({{\rm{d}}}z)$
+$\leq\gamma_{1}|x|^{2}+\gamma_{2}\mathcal{W}_{2}^{2}(\mu,\delta_{0})+\eta,$
+for any $p\in[2,p_{0}]$, $x\in\mathbb{R}^{d}$ and $\mu\in\mathcal{P}_{2}(\mathbb{R}^{d})$.
+Example 1.
+To illustrate that the aforementioned hypotheses are valid, we consider the following example: a variant of the Ginzburg-Landau equation. Specifically, consider
+[…]
+[…]
+where $\mu_{t}=\mathcal{L}^{1}(X_{t})$ denotes the conditional distribution of $X_{t}$, $m(\mu_{t})=\int y\,\mu_{t}({\rm d}y)$ is its first moment, $W$ is a Brownian motion, $W^{0}$ a common Brownian motion, and $Z$ a pure-jump Lévy process with Lévy measure $\nu({\rm d}z)$.
+For the jump part we choose a tempered stable Lévy measure on $\mathbb{R}_{0}$:
+$\nu(\mathrm{d}z)=C_{\alpha,\lambda}\,|z|^{-1-\alpha}e^{-\lambda|z|}\,\mathrm{d}z,\qquad\alpha\in(0,1),\;\lambda>0,$
+where $C_{\alpha,\lambda}$ is determined by the normalization condition ensuring that $\nu$ is a valid Lévy measure, i.e.
+$C_{\alpha,\lambda}=\frac{\lambda^{\alpha}}{2\Gamma(-\alpha)},$
+with $\Gamma(\cdot)$ denoting the Gamma function. Note that $\Gamma(s)=\int_{0}^{\infty}t^{s-1}e^{-t}\,\mathrm{d}t$ is the Gamma function. By setting $s=q-\alpha$, we have
+$\int_{\mathbb{R}_{0}}|z|^{q}\,\nu(\mathrm{d}z)=C_{\alpha,\lambda}\int_{\mathbb{R}_{0}}|z|^{q-1-\alpha}e^{-\lambda|z|}\,\mathrm{d}z=2C_{\alpha,\lambda}\,\lambda^{\alpha-q}\,\Gamma(q-\alpha),\qquad q>\alpha.$
+For $q=2$, we obtain
+$\int_{\mathbb{R}_{0}}|z|^{2}\,\nu(\mathrm{d}z)=2C_{\alpha,\lambda}\,\lambda^{\alpha-2}\,\Gamma(2-\alpha),\qquad 0<\alpha<1.$
+Next we verify that the above equation satisfies assumptions (A1)-(A7).
+(1) Since $b(x,\mu)=0.1(x-x^{3}+m(\mu))$, we have
+$2\langle x,b(x,\mu)\rangle$ $=2x\cdot 0.1\big(x-x^{3}+m(\mu)\big)=0.2\big(x^{2}-x^{4}+xm(\mu)\big).$
+Moreover,
+$|\sigma(x,\mu)|^{2}=(0.1x)^{2}=0.01x^{2},\qquad|\sigma^{0}(x,\mu)|^{2}=(0.05x)^{2}=0.0025x^{2}.$
+Also we have $|c(x,\mu)|^{2}\int_{\mathbb{R}_{0}}|z|^{2}\,\nu(\mathrm{d}z)=\sin^{2}(x)\cdot 2C_{\alpha,\lambda}\,\lambda^{\alpha-2}\,\Gamma(2-\alpha)\leq 2C_{\alpha,\lambda}\,\lambda^{\alpha-2}\,\Gamma(2-\alpha)$ and $m(\mu)^{2}\leq\mathcal{W}_{2}^{2}(\mu,\delta_{0}).$ Thus, we have
+$2\langle x,b(x,\mu)\rangle+|\sigma(x,\mu)|^{2}+|\sigma^{0}(x,\mu)|^{2}+|c(x,\mu)|^{2}\int_{\mathbb{R}_{0}}|z|^{2}\,\nu(\mathrm{d}z)$
+$\quad=0.2(x^{2}-x^{4}+xm(\mu))+0.0125x^{2}+\sin^{2}(x)\cdot 2C_{\alpha,\lambda}\,\lambda^{\alpha-2}\,\Gamma(2-\alpha)$
+$\quad\leq 0.2x^{2}+0.0125x^{2}+0.2|x||m(\mu)|+2C_{\alpha,\lambda}\,\lambda^{\alpha-2}\,\Gamma(2-\alpha)$
+$\quad\leq(0.2125+0.1)x^{2}+0.1\mathcal{W}_{2}^{2}(\mu,\delta_{0})+2C_{\alpha,\lambda}\,\lambda^{\alpha-2}\,\Gamma(2-\alpha)$
+$\quad=0.3125x^{2}+0.1\mathcal{W}_{2}^{2}(\mu,\delta_{0})+2C_{\alpha,\lambda}\,\lambda^{\alpha-2}\,\Gamma(2-\alpha).$
+Letting $L=\max\{0.3125,0.1,2C_{\alpha,\lambda}\,\lambda^{\alpha-2}\,\Gamma(2-\alpha)\}+1$, we show that the Assumption (A1) holds.
+(2) Note that
+$|\sigma(x)-\sigma(\overline{x})|^{2}=(0.1)^{2}|x-\overline{x}|^{2}=0.01|x-\overline{x}|^{2},$
+$|\sigma^{0}(x)-\sigma^{0}(\overline{x})|^{2}=(0.05)^{2}|x-\overline{x}|^{2}=0.0025|x-\overline{x}|^{2}.$
+For the jump coefficient, use $|\sin x-\sin y|\leq|x-y|$, so
+$|c(x)-c(\overline{x})|^{2}\int_{\mathbb{R}_{0}}|z|^{2}\,\nu(\mathrm{d}z)\leq 2C_{\alpha,\lambda}\,\lambda^{\alpha-2}\,\Gamma(2-\alpha)|x-\overline{x}|^{2}.$
+Hence,
+$2\langle x-\overline{x},b(x,\mu)-b(\overline{x},\overline{\mu})\rangle$
+$\quad=2(x-\overline{x})\cdot 0.1\big((x-x^{3})-(\overline{x}-\overline{x}^{3})+m(\mu)-m(\overline{\mu})\big)$
+$\quad=0.2(x-\overline{x})\big((x-x^{3})-(\overline{x}-\overline{x}^{3})\big)+0.2(x-\overline{x})(m(\mu)-m(\overline{\mu})).$
+Using the algebraic identity $(x-x^{3})-(\overline{x}-\overline{x}^{3})=(x-\overline{x})\big(1-(x^{2}+x\overline{x}+\overline{x}^{2})\big),$ we have
+$0.2(x-\overline{x})\big((x-x^{3})-(\overline{x}-\overline{x}^{3})\big)=0.2|x-\overline{x}|^{2}\big(1-(x^{2}+x\overline{x}+\overline{x}^{2})\big)\leq 0.2|x-\overline{x}|^{2}.$
+Moreover,
+$0.2(x-\overline{x})(m(\mu)-m(\overline{\mu}))\leq 0.1|x-\overline{x}|^{2}+0.1|m(\mu)-m(\overline{\mu})|^{2}.$
+Combining the above bounds gives
+$2\langle x-\overline{x},b(x,\mu)-b(\overline{x},\overline{\mu})\rangle+|\sigma(x,\mu)-\sigma(\overline{x},\overline{\mu})|^{2}+|\sigma^{0}(x,\mu)-\sigma^{0}(\overline{x},\overline{\mu})|^{2}$
+$\leq\big(0.2+0.1+0.01+0.0025+2C_{\alpha,\lambda}\,\lambda^{\alpha-2}\,\Gamma(2-\alpha)\big)|x-\overline{x}|^{2}+0.1|m(\mu)-m(\overline{\mu})|^{2}$
+$=\Big(0.3125+2C_{\alpha,\lambda}\,\lambda^{\alpha-2}\,\Gamma(2-\alpha)\Big)|x-\overline{x}|^{2}+0.1|m(\mu)-m(\overline{\mu})|^{2}.$
+Since $|m(\mu)-m(\overline{\mu})|\leq\mathcal{W}_{2}(\mu,\overline{\mu})$, we may choose
+$L_{1}:=0.3125+2C_{\alpha,\lambda}\,\lambda^{\alpha-2}\,\Gamma(2-\alpha),\qquad L_{2}:=0.1.$
+Hence Assumption (A2) holds with these constants.
+(3) Due to
+$|b(x,\mu)-b(\overline{x},\overline{\mu})|$ $=0.1\big|(x-x^{3})-(\overline{x}-\overline{x}^{3})+m(\mu)-m(\overline{\mu})\big|$
+$\leq 0.1\big|(x-x^{3})-(\overline{x}-\overline{x}^{3})\big|+0.1|m(\mu)-m(\overline{\mu})|.$
+Applying the mean-value theorem: there exists $\xi$ between $x$ and $\overline{x}$ such that
+$|(x-x^{3})-(\overline{x}-\overline{x}^{3})|=|x-\overline{x}||1-3\xi^{2}|\leq|x-\overline{x}|(1+3\xi^{2})\leq|x-\overline{x}|(1+3(|x|^{2}+|\overline{x}|^{2})).$
+Thus
+$|b(x,\mu)-b(\overline{x},\overline{\mu})|\leq 0.1\big(1+3(|x|^{2}+|\overline{x}|^{2})\big)|x-\overline{x}|+0.1\mathcal{W}_{2}(\mu,\overline{\mu}).$
+This implies that Assumption (A3) holds with $\ell=2$.
+(4) Note that $\Gamma(s,\lambda)=\int_{\lambda}^{\infty}t^{s-1}e^{-t}\,\mathrm{d}t$ is the upper incomplete Gamma function. When $2p_{0}>\alpha$, we have
+$\int_{|z|>1}|z|^{2p_{0}}\,\nu(\mathrm{d}z)$ $=2C_{\alpha,\lambda}\int_{1}^{\infty}r^{2p_{0}-1-\alpha}e^{-\lambda r}\,\mathrm{d}r$
+$=2C_{\alpha,\lambda}\,\lambda^{\alpha-2p_{0}}\,\Gamma(2p_{0}-\alpha,\lambda).$
+Similarly, due to $\gamma(s,\lambda)=\int_{0}^{\lambda}t^{s-1}e^{-t}\,\mathrm{d}t$ is the lower incomplete Gamma function. When $\alpha<1$, we have
+$\int_{0<|z|\leq 1}|z|\,\nu(\mathrm{d}z)$ $=2C_{\alpha,\lambda}\int_{0}^{1}r^{-\alpha}e^{-\lambda r}\,\mathrm{d}r$
+$=2C_{\alpha,\lambda}\,\lambda^{\alpha-1}\,\gamma(1-\alpha,\lambda).$
+Thus Assumption (A4) holds.
+(5) Since $|\sin x|\leq 1\leq 1+|x|+\mathcal{W}_{2}(\mu,\delta_{0})$ for all $x,\mu$, it’s easy to see that Assumption (A5) holds.
+(6) Since $b(x,\mu)=0.1(x-x^{3}+m(\mu))$,
+$\langle x,b(x,\mu)\rangle=0.1\big(x^{2}-x^{4}+xm(\mu)\big)\leq 0.1x^{2}+0.1|x||m(\mu)|\leq 0.1x^{2}+0.05x^{2}+0.05\mathcal{W}_{2}^{2}(\mu,\delta_{0}).$
+Next,
+$\frac{p_{0}-1}{2}(|\sigma(x,\mu)|^{2}+|\sigma^{0}(x,\mu)|^{2})$ $=\frac{p_{0}-1}{2}\cdot 0.0125x^{2}.$
+Using $|c(x,\mu)|^{2}=\sin^{2}(x)\leq 1$ and $L_{3}=1$, we have
+$\langle x,b(x,\mu)\rangle+\frac{p_{0}-1}{2}(|\sigma(x,\mu)|^{2}+|\sigma^{0}(x,\mu)|^{2})+\frac{1}{2L_{3}}|c(x,\mu)|^{2}\int_{\mathbb{R}_{0}}|z|\Big((1+L_{3}|z|)^{p_{0}-1}-1\Big)\nu(\mathrm{d}z)$
+$\quad\leq\Big(0.15+\frac{p_{0}-1}{2}\cdot 0.0125\Big)x^{2}+0.05\mathcal{W}_{2}^{2}(\mu,\delta_{0})+\tfrac{1}{2}\int_{\mathbb{R}_{0}}|z|\Big((1+|z|)^{p_{0}-1}-1\Big)\cdot C_{\alpha,\lambda}\,|z|^{-1-\alpha}e^{-\lambda|z|}\,\mathrm{d}z.$
+Let
+$I:=\tfrac{1}{2}\int_{\mathbb{R}_{0}}|z|\Big((1+|z|)^{p_{0}-1}-1\Big)\nu(\mathrm{d}z).$
+Setting $t=\lambda|z|$, we obtain
+$I$ $=C_{\alpha,\lambda}\int_{0}^{\infty}\Big(\frac{t}{\lambda}\Big)^{-\alpha}\Big(\bigl(1+\tfrac{t}{\lambda}\bigr)^{p_{0}-1}-1\Big)e^{-t}\,\frac{\mathrm{d}t}{\lambda}$
+$=C_{\alpha,\lambda}\,\lambda^{\alpha-1}\int_{0}^{\infty}t^{-\alpha}\Big(\bigl(1+\tfrac{t}{\lambda}\bigr)^{p_{0}-1}-1\Big)e^{-t}\,\mathrm{d}t.$
+As $t\to 0$, the term $\bigl(1+\tfrac{t}{\lambda}\bigr)^{p_{0}-1}-1\sim(p_{0}-1)\tfrac{t}{\lambda}$, so it is integrable if $\alpha<2$.
+As $t\to\infty$, exponential damping $e^{-t}$ ensures integrability for all $p_{0}$.
+Thus the integral $I$ is finite for all $p_{0}\in\mathbb{R}$ and $\alpha\in(0,1)$.
+Hence we may choose
+$\gamma_{1}:=0.15+\frac{p_{0}-1}{2}\cdot 0.0125,\qquad\gamma_{2}:=0.05,\qquad\eta:=\tfrac{1}{2}\int_{\mathbb{R}_{0}}|z|\Big((1+|z|)^{p_{0}-1}-1\Big)\nu(\mathrm{d}z).$
+This proves Assumption (A6).
+(7) Note that
+$|\sigma(x)-\sigma(\overline{x})|^{2}+|\sigma^{0}(x)-\sigma^{0}(\overline{x})|^{2}+|c(x)-c(\overline{x})|^{2}\int_{\mathbb{R}_{0}}|z|^{2}\,\nu(\mathrm{d}z)$
+$\quad=0.01|x-\overline{x}|^{2}+0.0025|x-\overline{x}|^{2}+2C_{\alpha,\lambda}\,\lambda^{\alpha-2}\,\Gamma(2-\alpha)|\sin x-\sin\overline{x}|^{2}$
+$\quad\leq(0.0125+2C_{\alpha,\lambda}\,\lambda^{\alpha-2}\,\Gamma(2-\alpha))|x-\overline{x}|^{2}.$
+Hence choose $\tilde{L}_{1}:=0.0125+2C_{\alpha,\lambda}\,\lambda^{\alpha-2}\,\Gamma(2-\alpha)$ and $\tilde{L}_{2}:=0$. This verifies Assumption (A7).
+Remark 2.2.
+We point out that assumptions (A1), (A2), (A4), (A6) and (A7) does not hold for an isotropic $\alpha$-stable Lévy measure with index $\alpha\in(0,2)$, and the measure $\nu(\mathrm{d}z)=c_{d,\alpha}\,|z|^{-d-\alpha}\,\mathrm{d}z.$ In fact
+$\int_{|z|>1}|z|^{2}\,\nu(\mathrm{d}z)=c_{d,\alpha}\int_{|z|>1}|z|^{2-d-\alpha}\,\mathrm{d}z.$
+Using spherical coordinates,
+$\int_{|z|>1}|z|^{2-d-\alpha}\,\mathrm{d}z=\int_{1}^{\infty}\int_{S^{d-1}}r^{2-d-\alpha}\,r^{d-1}\,\mathrm{d}\sigma(\theta)\,\mathrm{d}r=S_{d-1}\int_{1}^{\infty}r^{1-\alpha}\,\mathrm{d}r,$
+where $S_{d-1}=2\pi^{d/2}/\Gamma(d/2)$ is the surface area of the (d1)(d-1)-sphere.
+The radial integral is
+$\int_{1}^{\infty}r^{1-\alpha}\,\mathrm{d}r=\frac{1}{2-\alpha}\left(r^{2-\alpha}\right)\Big|_{1}^{\infty}.$
+For $\alpha\in(0,2)$ we have $2-\alpha>0$, hence the limit as $r\to\infty$ diverges. Thus
+$\int_{|z|>1}|z|^{2}\,\nu(\mathrm{d}z)=\infty.$
+Therefore the isotropic $\alpha$-stable Lévy measure does not admit a finite second moment.
 Theorem 2.3.
@@
 Proof.
-To make the study easy to follow, we divide the proof into the following steps.
+To facilitate readability, we divide the proof into the following steps.
   1. $\bf(1)$}
@@
 […]
-$\rho_{\infty}(\xi,\tilde{\xi}):=\sup_{t\in[0,T]}\left\{e^{-\lambda t}|\xi_{t}-\tilde{\xi}_{t}|\right\},\quad\xi,\tilde{\xi}\in\mathcal{D}([0,T];\mathbb{R}^{d}),\quad\lambda>0.$
+$\rho_{\infty}(\xi,\tilde{\xi}):=\sup_{t\in[0,T]}\left\{|\xi_{t}-\tilde{\xi}_{t}|\right\},\quad\xi,\tilde{\xi}\in\mathcal{D}([0,T];\mathbb{R}^{d}).$
 Then, the space $\mathcal{D}([0,T];\mathbb{R}^{d})$ endowed with the distance $\rho_{\infty}$ is a complete space.
@@
 $\bm{\rho}_{p,\infty}(\eta,\tilde{\eta}):=\bigg[\mathbb{E}\rho^{p}_{\infty}(\eta,\tilde{\eta})\bigg]^{1/p},\quad\eta,\tilde{\eta}\in\mathscr{Y}_{T}(p).$
-The following lemma focuses on the completeness of $(\mathscr{Y}_{T}(p),\bm{\rho}_{p,\infty})$, which is essentially due to Shao et al.
-[40, Lemma 2.3 ].
-We include a complete proof for readers’ convenience.
-Lemma 2.1.
-The metric space $(\mathscr{Y}_{T}(p),\bm{\rho}_{p,\infty})$ is complete for every $p\geq 1$.
-Proof.
-Let $(X^{(n)})_{n\geq 1}$ be a Cauchy sequence in $(\mathscr{Y}_{T}(p),\bm{\rho}_{p,\infty})$, i.e.,
-$\bm{\rho}_{p,\infty}(X^{(n)},X^{(m)})^{p}\to 0\quad\text{as }n,m\to\infty.$
-Take a subsequence $(X^{(n_{k})})$ such that
-$\bm{\rho}_{p,\infty}(X^{(n_{k})},X^{(n_{k+1})})<\frac{1}{2^{k}},\quad k\geq 1,$
-which implies
-$\mathbb{P}\left(\rho_{\infty}(X^{(n_{k})},X^{(n_{k+1})})\geq\frac{1}{k^{2}}\right)\leq\frac{k^{2p}}{2^{kp}},\quad k\geq 1,$
-Since $\sum_{k\geq 1}\frac{k^{2p}}{2^{kp}}<\infty$, the Borel-Cantelli lemma yields
-$\mathbb{P}\left(\rho_{\infty}(X^{(n_{k})},X^{(n_{k+1})})\geq\frac{1}{k^{2}},\text{ i.o.}\right)=0.$
-Thus, there exists a measurable set $\Xi$ with $\mathbb{P}(\Xi)=1$ such that for every $\omega\in\Xi$,
-$\sum_{k=1}^{\infty}\rho_{\infty}(X^{(n_{k})}(\omega),X^{(n_{k+1})}(\omega))<\infty.$
-By the completeness of $\mathcal{D}([0,T];\mathbb{R}^{d})$ with respect to $\rho_{\infty}$, for every $\omega\in\Xi$, there exists $\tilde{X}(\omega)\in\mathcal{D}([0,T];\mathbb{R}^{d})$ such that
-$\lim_{k\to\infty}\rho_{\infty}\bigl{(}X^{(n_{k})}(\omega),\,\tilde{X}(\omega)\bigr{)}=0,\quad\omega\in\Xi.$
-We now aim to prove that $\tilde{X}\in\mathscr{Y}_{T}(p)$ and that
-$\bm{\rho}_{p,\infty}\big{(}X^{(n_{k})},\tilde{X}\big{)}\xrightarrow{\;k\to\infty\;}0.$
-We first show that the Cauchy sequence $(X^{(n)})_{n\geq 1}$ is uniformly $p$-integrable.
-Since $(X^{(n)})$ is Cauchy in $(\mathscr{Y}_{T}(p),\bm{\rho}_{p,\infty})$, there exists some fixed index $n_{0}$ and a constant $C<\infty$ such that
-$\sup_{n\geq 1}\bm{\rho}_{p,\infty}\big{(}X^{(n)},X^{(n_{0})}\big{)}^{p}=\sup_{n\geq 1}\mathbb{E}\Big{[}\sup_{t\in[0,T]}e^{-\lambda t}|X^{(n)}_{t}-X^{(n_{0})}_{t}|^{p}\Big{]}\leq C.$
-Using the elementary inequality $|a+b|^{p}\leq 2^{p-1}(|a|^{p}+|b|^{p})$, we obtain for every $n\geq 1$
-$\mathbb{E}\Big{[}\sup_{t\in[0,T]}e^{-\lambda t}|X^{(n)}_{t}|^{p}\Big{]}$ $\leq 2^{p-1}\mathbb{E}\Big{[}\sup_{t\in[0,T]}e^{-\lambda t}|X^{(n)}_{t}-X^{(n_{0})}_{t}|^{p}\Big{]}+2^{p-1}\mathbb{E}\Big{[}\sup_{t\in[0,T]}e^{-\lambda t}|X^{(n_{0})}_{t}|^{p}\Big{]}$
-$\leq 2^{p-1}C+2^{p-1}\mathbb{E}\Big{[}\sup_{t\in[0,T]}e^{-\lambda t}|X^{(n_{0})}_{t}|^{p}\Big{]}=:C_{1}<\infty.$
-Hence $\sup_{n\geq 1}\mathbb{E}\big{[}\sup_{t\in[0,T]}e^{-\lambda t}|X^{(n)}_{t}|^{p}\big{]}\leq C_{1}<\infty$.
-Now consider the limit process $\tilde{X}$.
-By Fatou’s lemma and the Cauchy property of $X^{(n_{k})}\to\tilde{X}$,
-$\mathbb{E}\Big{[}\sup_{t\in[0,T]}e^{-\lambda t}|\tilde{X}_{t}|^{p}\Big{]}\leq\liminf_{k\to\infty}\mathbb{E}\Big{[}\sup_{t\in[0,T]}e^{-\lambda t}|X^{(n_{k})}_{t}|^{p}\Big{]}\leq C_{1}<\infty.$
-[…]
-$\sup_{t\in[0,T]}\mathbb{E}[|\tilde{X}_{t}|^{p}]\leq e^{\lambda T}\,\mathbb{E}\Big{[}\sup_{s\in[0,T]}e^{-\lambda s}|\tilde{X}_{s}|^{p}\Big{]}.$
-Since the right-hand side has already been shown to be finite (bounded by $C_{1}$), we conclude
-$\sup_{t\in[0,T]}\mathbb{E}[|\tilde{X}_{t}|^{p}]\leq e^{\lambda T}C_{1}<\infty.$
-Thus we have $\sup_{t\in[0,T]}\mathbb{E}[|\tilde{X}_{t}|^{p}]<\infty$.
-Note that for every $\omega\in\Xi$, the path $\tilde{X}(\omega)$ belongs to $\mathcal{D}([0,T];\mathbb{R}^{d})$.
-Thus, it remains only to verify that $\tilde{X}_{t}$ is $\mathcal{F}_{t}$-measurable for each $t\in[0,T]$.
-Let $\tilde{\Gamma}$ denote the law of $\tilde{X}$ as a random element of $\mathcal{D}([0,T];\mathbb{R}^{d})$.
-We introduce the set
-$\mathcal{F}_{\tilde{\Gamma}}:=\{\,t\in[0,T]\mid\tilde{\Gamma}(J_{t})=0\,\},$
-where
-$J_{t}:=\{\,X\in\mathcal{D}([0,T];\mathbb{R}^{d}):X_{t}\neq X_{t-}\,\}$
-[…]
-$\lim_{k\to\infty}|X_{t}^{(n_{k})}-\tilde{X}_{t}|=0,\quad\mathbb{P}\text{-a.e.},\forall t\in\mathcal{F}_{\tilde{\Gamma}}$.
-Since $X_{t}^{(n_{k})}$ is $\mathcal{F}_{t}$-adapted, $\tilde{X}_{t}$ is $\mathcal{F}_{t}$-adapted for $t\in\mathcal{F}_{\tilde{\Gamma}}$.
-Due to right‐continuity of the paths, for any $t\in(0,T)$, we choose a sequence $t_{n}\downarrow t$ such that $\tilde{X}_{t_{n}}\;\xrightarrow{\;n\to\infty\;}\;\tilde{X}_{t}\quad\mathbb{P}\text{-a.s.}.$ Since $\mathcal{F}_{t+}=\mathcal{F}_{t}$, it follows that $\tilde{X}_{t}\;\in\;\mathcal{F}_{t}.$ Hence, $\tilde{X}\in\mathscr{Y}_{T}(p)$.
-By Fatou’s lemma and the Cauchy property,
-[…]
-$=$ $\lim_{n\to\infty}\bigg{[}\mathbb{E}\sup_{t\in[0,T]}e^{-\lambda t}|X^{(n)}_{t}-\lim_{l\to\infty}X^{(n_{l})}_{t}|^{p}\bigg{]}^{1/p}$
-$\leq$ $\lim_{n\to\infty}\mathop{\underline{\lim}}\limits_{l\to\infty}\bigg{[}\mathbb{E}\sup_{t\in[0,T]}e^{-\lambda t}|X^{(n)}_{t}-X^{(n_{l})}_{t}|^{p}\bigg{]}^{1/p}=0.$
-Thus, $(\mathscr{Y}_{T}(p),\bm{\rho}_{p,\infty})$ is complete. ∎
+The completeness of $(\mathscr{Y}_{T}(p),\bm{\rho}_{p,\infty})$ follows essentially from Lemma 2.3 in Shao et al. [31].
   1. $\bf(2)$}
@@
 almost surely for any $t\in[0,T],$ with $Y_{0}^{\mu}=X_{0}$.
-According to [37, Theorem 2.1], the classical SDE (2.1) admits a unique solution with $\sup_{t\in[0,T]}\mathbb{E}\bigl{[}\lvert Y_{t}^{\mu}\rvert^{2}\bigr{]}<\infty.$ Moreover, we have $(Y_{t}^{\mu})_{t\in[0,T]}\in\mathscr{Y}_{T}(2)$.
+According to [28, Theorem 2.1], the classical SDE (2.1) admits a unique solution with $\sup_{t\in[0,T]}\mathbb{E}\bigl[\lvert Y_{t}^{\mu}\rvert^{2}\bigr]<\infty.$ Moreover, we have $(Y_{t}^{\mu})_{t\in[0,T]}\in\mathscr{Y}_{T}(2)$.
   1. $\bf(3)$}
@@
 We show that $\Phi$ is a strict contraction in $(\mathscr{Y}_{T}(2),\bm{\rho}_{2,\infty})$.
-For another $\tilde{\eta}\in\mathscr{Y}_{T}(2)$, let $(Y_{t}^{\nu})$ solve SDE (2.1) by replacing $\mu_{t}=\mathcal{L}(\eta_{t}|\mathcal{F}_{t}^{0})$ there with $\nu_{t}=\mathcal{L}(\tilde{\eta}_{t}|\mathcal{F}_{t}^{0})$ for $t\in[0,T]$.
-Applying Itô’s formula to $|Y_{t}^{\mu}-Y_{t}^{\nu}|^{2}$, taking the supremum over $t\in[0,T]$, taking expectation and using the Assumption (A2), we get
+For another $\tilde{\eta}\in\mathscr{Y}_{T}(2)$, let $(Y_{t}^{\nu})_{t\in[0,T]}$ solve SDE (2.1) by replacing $\mu_{t}=\mathcal{L}(\eta_{t}|\mathcal{F}_{t}^{0})$ there with $\nu_{t}=\mathcal{L}(\tilde{\eta}_{t}|\mathcal{F}_{t}^{0})$ for $t\in[0,T]$.
+Applying Itô’s formula to $|Y_{t}^{\mu}-Y_{t}^{\nu}|^{2}$, taking the supremum over $0,t]\subset[0,T]$ and expectation, and using the Assumption (A2), we get
 $\mathbb{E}\Big[\sup_{0\leq s\leq t}\big|Y_{s}^{\mu}-Y_{s}^{\nu}\big|^{2}\Big]$ (2.2)
@@
 $\quad+2\,\mathbb{E}\left[\sup_{0\leq r\leq t}\int_{0}^{r}\left\langle Y_{s}^{\mu}-Y_{s}^{\nu},(\sigma^{0}(Y_{s}^{\mu},\mu_{s})-\sigma^{0}(Y_{s}^{\nu},\nu_{s}))\mathrm{d}W_{s}^{0}\right\rangle\right]$
-$\quad+2\,\mathbb{E}\left[\sup_{0\leq r\leq t}\int_{0}^{r}\int_{\mathbb{R}_{0}^{d}}\Big{(}|Y_{s^{-}}^{\mu}-Y_{s^{-}}^{\nu}+\delta_{s}(z)|^{2}-|Y_{s^{-}}^{\mu}-Y_{s^{-}}^{\nu}|^{2}\Big{)}\widetilde{N}(\mathrm{d}s,\mathrm{d}z)\right]$
+$\quad+2\,\mathbb{E}\left[\sup_{0\leq r\leq t}\int_{0}^{r}\int_{\mathbb{R}_{0}^{d}}\Big(|Y_{s^{-}}^{\mu}-Y_{s^{-}}^{\nu}+\big(c(Y_{s^{-}}^{\mu},\mu_{s})-c(Y_{s^{-}}^{\nu},\nu_{s})\big)\,z|^{2}-|Y_{s^{-}}^{\mu}-Y_{s^{-}}^{\nu}|^{2}\Big)\widetilde{N}(\mathrm{d}s,\mathrm{d}z)\right]$
 $\leq\mathbb{E}\Big[\!\sup_{0\leq r\leq t}\int_{0}^{r}\big(L_{1}|Y_{s}^{\mu}-Y_{s}^{\nu}|^{2}+L_{2}\mathcal{W}_{2}^{2}(\mu_{s},\nu_{s})\big)\,\mathrm{d}s\Big]$
@@
 $\quad+2\,\mathbb{E}\left[\sup_{0\leq r\leq t}\int_{0}^{r}\left\langle Y_{s}^{\mu}-Y_{s}^{\nu},(\sigma^{0}(Y_{s}^{\mu},\mu_{s})-\sigma^{0}(Y_{s}^{\nu},\nu_{s}))\mathrm{d}W_{s}^{0}\right\rangle\right]$
-$\quad+2\,\mathbb{E}\left[\sup_{0\leq r\leq t}\int_{0}^{r}\int_{\mathbb{R}_{0}^{d}}\Big{(}|Y_{s^{-}}^{\mu}-Y_{s^{-}}^{\nu}+\delta_{s}(z)|^{2}-|Y_{s^{-}}^{\mu}-Y_{s^{-}}^{\nu}|^{2}\Big{)}\widetilde{N}(\mathrm{d}s,\mathrm{d}z)\right]$
-Applying the Burkholder-Davis-Gundy (BDG) inequality, Jensen’s inequality and using the Assumption (A7) to each stochastic integral, we obtain
-$\mathbb{E}\left[\sup_{0\leq r\leq t}\left|\int_{0}^{r}\langle Y_{s}^{\mu}-Y_{s}^{\nu},(\sigma(Y_{s}^{\mu},\mu_{s})-\sigma(Y_{s}^{\nu},\nu_{s}))\mathrm{d}W_{s}\rangle\right|\right]$
+$\quad+2\,\mathbb{E}\left[\sup_{0\leq r\leq t}\int_{0}^{r}\int_{\mathbb{R}_{0}^{d}}\Big(|Y_{s^{-}}^{\mu}-Y_{s^{-}}^{\nu}+\big(c(Y_{s^{-}}^{\mu},\mu_{s})-c(Y_{s^{-}}^{\nu},\nu_{s})\big)\,z|^{2}-|Y_{s^{-}}^{\mu}-Y_{s^{-}}^{\nu}|^{2}\Big)\widetilde{N}(\mathrm{d}s,\mathrm{d}z)\right].$
+Moreover, due to
+$\mathbb{E}[|Y_{s}^{\mu}-Y_{s}^{\nu}|^{2}]\leq\mathbb{E}\Big[\sup_{0\leq r\leq s}|Y_{r}^{\mu}-Y_{r}^{\nu}|^{2}\Big],\qquad\mathbb{E}[\mathcal{W}_{2}^{2}(\mu_{s},\nu_{s})]\leq\mathbb{E}\Big[\sup_{0\leq r\leq s}\mathcal{W}_{2}^{2}(\mu_{r},\nu_{r})\Big],$
+we obtain
+$\mathbb{E}\Bigg[\int_{0}^{t}\Big(\tilde{L}_{1}|Y_{s}^{\mu}-Y_{s}^{\nu}|^{2}+\tilde{L}_{2}\mathcal{W}_{2}^{2}(\mu_{s},\nu_{s})\Big)\,\mathrm{d}s\Bigg]$
+$=\int_{0}^{t}\Big(\tilde{L}_{1}\,\mathbb{E}[|Y_{s}^{\mu}-Y_{s}^{\nu}|^{2}]+\tilde{L}_{2}\,\mathbb{E}[\mathcal{W}_{2}^{2}(\mu_{s},\nu_{s})]\Big)\,\mathrm{d}s$
+$\leq\int_{0}^{t}\Big(\tilde{L}_{1}\,\mathbb{E}[\sup_{0\leq r\leq s}|Y_{r}^{\mu}-Y_{r}^{\nu}|^{2}]+\tilde{L}_{2}\,\mathbb{E}[\sup_{0\leq r\leq s}\mathcal{W}_{2}^{2}(\mu_{r},\nu_{r})]\Big)\,\mathrm{d}s.$
+Therefore,
+$C_{1}\,\mathbb{E}\Bigg[\Bigg(\int_{0}^{t}|Y_{s}^{\mu}-Y_{s}^{\nu}|^{2}\Big(\tilde{L}_{1}|Y_{s}^{\mu}-Y_{s}^{\nu}|^{2}+\tilde{L}_{2}\mathcal{W}_{2}^{2}(\mu_{s},\nu_{s})\Big)\,\mathrm{d}s\Bigg)^{1/2}\Bigg]$
+$\quad\leq C_{1}\Big(\mathbb{E}[\sup_{0\leq r\leq t}|Y_{r}^{\mu}-Y_{r}^{\nu}|^{2}]\Big)^{1/2}\Bigg(\int_{0}^{t}\Big(\tilde{L}_{1}\,\mathbb{E}[\sup_{0\leq r\leq s}|Y_{r}^{\mu}-Y_{r}^{\nu}|^{2}]+\tilde{L}_{2}\,\mathbb{E}[\sup_{0\leq r\leq s}\mathcal{W}_{2}^{2}(\mu_{r},\nu_{r})]\Big)\,\mathrm{d}s\Bigg)^{1/2}.$
+Applying Young’s inequality $uv\leq\tfrac{1}{2}\epsilon u^{2}+\tfrac{1}{2\epsilon}v^{2}$, let $\epsilon=\frac{1}{4C_{1}}$ we have
+$C_{1}\,\mathbb{E}\Bigg[\Bigg(\int_{0}^{t}|Y_{s}^{\mu}-Y_{s}^{\nu}|^{2}\Big(\tilde{L}_{1}|Y_{s}^{\mu}-Y_{s}^{\nu}|^{2}+\tilde{L}_{2}\mathcal{W}_{2}^{2}(\mu_{s},\nu_{s})\Big)\,\mathrm{d}s\Bigg)^{1/2}\Bigg]$
+$\quad\leq\frac{C_{1}\epsilon}{2}\,\mathbb{E}\Big[\sup_{0\leq r\leq t}|Y_{r}^{\mu}-Y_{r}^{\nu}|^{2}\Big]+\frac{C_{1}}{2\epsilon}\int_{0}^{t}\Big(\tilde{L}_{1}\,\mathbb{E}[\sup_{0\leq r\leq s}|Y_{r}^{\mu}-Y_{r}^{\nu}|^{2}]+\tilde{L}_{2}\,\mathbb{E}[\sup_{0\leq r\leq s}\mathcal{W}_{2}^{2}(\mu_{r},\nu_{r})]\Big)\,\mathrm{d}s$
+$\quad\leq\frac{1}{8}\,\mathbb{E}\Big[\sup_{0\leq r\leq t}|Y_{r}^{\mu}-Y_{r}^{\nu}|^{2}\Big]+2{C_{1}}^{2}\int_{0}^{t}\Big(\tilde{L}_{1}\,\mathbb{E}[\sup_{0\leq r\leq s}|Y_{r}^{\mu}-Y_{r}^{\nu}|^{2}]+\tilde{L}_{2}\,\mathbb{E}[\sup_{0\leq r\leq s}\mathcal{W}_{2}^{2}(\mu_{r},\nu_{r})]\Big)\,\mathrm{d}s.$
+Finally, using
+$\tilde{L}_{1}\,\mathbb{E}[\sup_{0\leq r\leq s}|Y_{r}^{\mu}-Y_{r}^{\nu}|^{2}]+\tilde{L}_{2}\,\mathbb{E}[\sup_{0\leq r\leq s}\mathcal{W}_{2}^{2}(\mu_{r},\nu_{r})]$
+$\quad\leq\max\{\tilde{L}_{1},\tilde{L}_{2}\}\Big(\mathbb{E}[\sup_{0\leq r\leq s}|Y_{r}^{\mu}-Y_{r}^{\nu}|^{2}]+\mathbb{E}[\sup_{0\leq r\leq s}\mathcal{W}_{2}^{2}(\mu_{r},\nu_{r})]\Big),$
+we obtain
+$C_{1}\,\mathbb{E}\Bigg[\Bigg(\int_{0}^{t}|Y_{s}^{\mu}-Y_{s}^{\nu}|^{2}\Big(\tilde{L}_{1}|Y_{s}^{\mu}-Y_{s}^{\nu}|^{2}+\tilde{L}_{2}\mathcal{W}_{2}^{2}(\mu_{s},\nu_{s})\Big)\,\mathrm{d}s\Bigg)^{1/2}\Bigg]$ (2.3)
+$\leq\frac{1}{8}\,\mathbb{E}\Big[\sup_{0\leq r\leq t}|Y_{r}^{\mu}-Y_{r}^{\nu}|^{2}\Big]+2{C_{1}}^{2}\,\max\{\tilde{L}_{1},\tilde{L}_{2}\}$
+$\quad\times\int_{0}^{t}\Big(\mathbb{E}[\sup_{0\leq r\leq s}|Y_{r}^{\mu}-Y_{r}^{\nu}|^{2}]+\mathbb{E}[\sup_{0\leq r\leq s}\mathcal{W}_{2}^{2}(\mu_{r},\nu_{r})]\Big)\,\mathrm{d}s.$
+By using BDG’s inequality and (2.3), we have
+$\mathbb{E}\left[\sup_{0\leq r\leq t}\left|\int_{0}^{r}\langle Y_{s}^{\mu}-Y_{s}^{\nu},(\sigma(Y_{s}^{\mu},\mu_{s})-\sigma(Y_{s}^{\nu},\nu_{s}))\,\mathrm{d}W_{s}\rangle\right|\right]$ (2.4)
 $\quad\leq C_{1}\,\mathbb{E}\left[\left(\int_{0}^{t}|Y_{s}^{\mu}-Y_{s}^{\nu}|^{2}\cdot|\sigma(Y_{s}^{\mu},\mu_{s})-\sigma(Y_{s}^{\nu},\nu_{s})|^{2}\,\mathrm{d}s\right)^{1/2}\right]$
 $\quad\leq C_{1}\,\mathbb{E}\left[\left(\int_{0}^{t}|Y_{s}^{\mu}-Y_{s}^{\nu}|^{2}\cdot\left(\tilde{L}_{1}|Y_{s}^{\mu}-Y_{s}^{\nu}|^{2}+\tilde{L}_{2}\mathcal{W}_{2}^{2}(\mu_{s},\nu_{s})\right)\,\mathrm{d}s\right)^{1/2}\right]$
-$\quad\leq K_{1}\int_{0}^{t}\left(\mathbb{E}[\sup_{0\leq r\leq s}e^{-\lambda r}|Y_{r}^{\mu}-Y_{r}^{\nu}|^{2}+\sup_{0\leq r\leq s}e^{-\lambda r}\mathcal{W}_{2}^{2}(\mu_{r},\nu_{r})]\right){\rm{d}}s,$
-where $K_{1}:=K_{1}(C_{1},\tilde{L}_{1},\tilde{L}_{2})$ and similarly for the $W^{0}$ term and the jump term with $K_{2},K_{3}$ respectively. Therefore,
+$\quad\leq\frac{1}{8}\,\mathbb{E}\Big[\sup_{0\leq r\leq t}|Y_{r}^{\mu}-Y_{r}^{\nu}|^{2}\Big]+K_{1}\int_{0}^{t}\left(\mathbb{E}\Big[\sup_{0\leq r\leq s}|Y_{r}^{\mu}-Y_{r}^{\nu}|^{2}\Big]+\mathbb{E}\Big[\sup_{0\leq r\leq s}\mathcal{W}_{2}^{2}(\mu_{r},\nu_{r})\Big]\right)\,\mathrm{d}s,$
+where $K_{1}\;=\;2{C_{1}}^{2}\cdot\,\max\{\tilde{L}_{1},\tilde{L}_{2}\}$. In the same way for the term driven by $W$, we have
+$\mathbb{E}\left[\sup_{0\leq r\leq t}\int_{0}^{r}\left\langle Y_{s}^{\mu}-Y_{s}^{\nu},(\sigma^{0}(Y_{s}^{\mu},\mu_{s})-\sigma^{0}(Y_{s}^{\nu},\nu_{s}))\mathrm{d}W_{s}^{0}\right\rangle\right]$ (2.5)
+$\quad\leq C_{2}\,\mathbb{E}\left[\left(\int_{0}^{t}|Y_{s}^{\mu}-Y_{s}^{\nu}|^{2}\cdot|\sigma^{0}(Y_{s}^{\mu},\mu_{s})-\sigma^{0}(Y_{s}^{\nu},\nu_{s})|^{2}\,\mathrm{d}s\right)^{1/2}\right]$
+$\quad\leq C_{2}\,\mathbb{E}\left[\left(\int_{0}^{t}|Y_{s}^{\mu}-Y_{s}^{\nu}|^{2}\cdot\left(\tilde{L}_{1}|Y_{s}^{\mu}-Y_{s}^{\nu}|^{2}+\tilde{L}_{2}\mathcal{W}_{2}^{2}(\mu_{s},\nu_{s})\right)\,\mathrm{d}s\right)^{1/2}\right]$
+$\quad\leq\frac{1}{8}\,\mathbb{E}\Big[\sup_{0\leq r\leq t}|Y_{r}^{\mu}-Y_{r}^{\nu}|^{2}\Big]+K_{2}\int_{0}^{t}\Big(\mathbb{E}\big[\sup_{0\leq r\leq s}|Y_{r}^{\mu}-Y_{r}^{\nu}|^{2}\big]+\mathbb{E}\big[\sup_{0\leq r\leq s}\mathcal{W}_{2}^{2}(\mu_{r},\nu_{r})\big]\Big)\,\mathrm{d}s,$
+where $K_{2}=2C_{2}^{2}\cdot\max\{\tilde{L}_{1},\tilde{L}_{2}\}$.
+As for the jump term, let
+$M_{r}:=\int_{0}^{r}\int_{\mathbb{R}_{0}^{d}}\Big(|Y_{s^{-}}^{\mu}-Y_{s^{-}}^{\nu}+\big(c(Y_{s^{-}}^{\mu},\mu_{s})-c(Y_{s^{-}}^{\nu},\nu_{s})\big)\,z|^{2}-|Y_{s^{-}}^{\mu}-Y_{s^{-}}^{\nu}|^{2}\Big)\widetilde{N}(\mathrm{d}s,\mathrm{d}z),\qquad r\in[0,t].$
+Note that
+$\Big|\big|Y_{s^{-}}^{\mu}-Y_{s^{-}}^{\nu}+\big(c(Y_{s^{-}}^{\mu},\mu_{s})-c(Y_{s^{-}}^{\nu},\nu_{s})\big)\,z\big|^{2}-|Y_{s^{-}}^{\mu}-Y_{s^{-}}^{\nu}|^{2}\Big|$
+$\quad\leq 2\,|Y_{s^{-}}^{\mu}-Y_{s^{-}}^{\nu}|\,\cdot\big|c(Y_{s^{-}}^{\mu},\mu_{s})-c(Y_{s^{-}}^{\nu},\nu_{s})\big|\,|z|+\big|c(Y_{s^{-}}^{\mu},\mu_{s})-c(Y_{s^{-}}^{\nu},\nu_{s})\big|^{2}\,|z|^{2}.$
+By the Kunita’s inequality (see [14, Theorem 4.4.23]) for compensated Poisson integrals there exists a constant $C_{3}>0$ such that
+$\mathbb{E}\Big[\sup_{0\leq r\leq t}|M_{r}|\Big]\leq C_{3}\Bigg\{\mathbb{E}\Big[\Big(\int_{0}^{t}\int_{\mathbb{R}^{d}_{0}}\Big|2\big\langle Y_{s^{-}}^{\mu}-Y_{s^{-}}^{\nu},\;c(Y_{s^{-}}^{\mu},\mu_{s})-c(Y_{s^{-}}^{\nu},\nu_{s})\big\rangle\,z$
+$+\big|c(Y_{s^{-}}^{\mu},\mu_{s})-c(Y_{s^{-}}^{\nu},\nu_{s})\big|^{2}\,|z|^{2}\Big|^{2}\,\nu(\mathrm{d}z)\,\mathrm{d}s\Big)^{1/2}\Big]+\;\mathbb{E}\Big[\int_{0}^{t}\int_{\mathbb{R}^{d}_{0}}\Big|2\big\langle Y_{s^{-}}^{\mu}-Y_{s^{-}}^{\nu},\;c(Y_{s^{-}}^{\mu},\mu_{s})-c(Y_{s^{-}}^{\nu},\nu_{s})\big\rangle\,z$
+$+\big|c(Y_{s^{-}}^{\mu},\mu_{s})-c(Y_{s^{-}}^{\nu},\nu_{s})\big|^{2}\,|z|^{2}\Big|\,\nu(\mathrm{d}z)\,\mathrm{d}s\Big]\Bigg\}.$
+Using $|a+b|^{2}\leq 2|a|^{2}+2|b|^{2}$, $|a+b|\leq|a|+|b|$ and the Assumption (A7), we get
+$\int_{0}^{t}\int_{\mathbb{R}^{d}_{0}}\Big|2\langle Y_{s^{-}}^{\mu}-Y_{s^{-}}^{\nu},c(Y_{s^{-}}^{\mu},\mu_{s})-c(Y_{s^{-}}^{\nu},\nu_{s})\rangle z+|c(Y_{s^{-}}^{\mu},\mu_{s})-c(Y_{s^{-}}^{\nu},\nu_{s}|^{2}|z|^{2}\Big|^{2}\nu({\rm{d}}z){\rm{d}}s$
+$\quad\leq\int_{0}^{t}\int_{\mathbb{R}^{d}_{0}}8|Y_{s^{-}}^{\mu}-Y_{s^{-}}^{\nu}|^{2}\cdot|c(Y_{s^{-}}^{\mu},\mu_{s})-c(Y_{s^{-}}^{\nu},\nu_{s})|^{2}|z|^{2}+2|c(Y_{s^{-}}^{\mu},\mu_{s})-c(Y_{s^{-}}^{\nu},\nu_{s})|^{4}|z|^{4}\,\nu({\rm{d}}z){\rm{d}}s$
+$\quad\leq\int_{0}^{t}8|Y_{s^{-}}^{\mu}-Y_{s^{-}}^{\nu}|^{2}\big(\tilde{L}_{1}|Y_{s^{-}}^{\mu}-Y_{s^{-}}^{\nu}|^{2}+\tilde{L}_{2}\mathcal{W}_{2}^{2}(\mu_{s},\nu_{s})\big){\rm{d}}s$
+$\qquad+2\big(\tilde{L}_{1}|Y_{s^{-}}^{\mu}-Y_{s^{-}}^{\nu}|^{2}+\tilde{L}_{2}\mathcal{W}_{2}^{2}(\mu_{s},\nu_{s})\big)^{2}\bigg(\int_{\mathbb{R}^{d}_{0}}|z|^{2}\nu(\mathrm{d}z)\bigg)^{-2}\int_{\mathbb{R}^{d}_{0}}|z|^{4}\nu({\rm{d}}z){\rm{d}}s,$
+and
+$\int_{0}^{t}\int_{\mathbb{R}^{d}_{0}}\Big|2\langle Y_{s^{-}}^{\mu}-Y_{s^{-}}^{\nu},c(Y_{s^{-}}^{\mu},\mu_{s})-c(Y_{s^{-}}^{\nu},\nu_{s})\rangle z+|c(Y_{s^{-}}^{\mu},\mu_{s})-c(Y_{s^{-}}^{\nu},\nu_{s}|^{2}|z|^{2}\Big|\nu({\rm{d}}z){\rm{d}}s$
+$\quad\leq\int_{0}^{t}\int_{\mathbb{R}^{d}_{0}}2|Y_{s^{-}}^{\mu}-Y_{s^{-}}^{\nu}|\cdot|c(Y_{s^{-}}^{\mu},\mu_{s})-c(Y_{s^{-}}^{\nu},\nu_{s})||z|+|c(Y_{s^{-}}^{\mu},\mu_{s})-c(Y_{s^{-}}^{\nu},\nu_{s}|^{2}|z|^{2}\,\nu({\rm{d}}z){\rm{d}}s$
+$\quad\leq\int_{0}^{t}2|Y_{s^{-}}^{\mu}-Y_{s^{-}}^{\nu}|\sqrt{\tilde{L}_{1}|Y_{s^{-}}^{\mu}-Y_{s^{-}}^{\nu}|^{2}+\tilde{L}_{2}\mathcal{W}_{2}^{2}(\mu_{s},\nu_{s})}\bigg(\int_{\mathbb{R}^{d}_{0}}|z|^{2}\nu(\mathrm{d}z)\bigg)^{-1/2}\int_{\mathbb{R}^{d}_{0}}|z|\nu({\rm{d}}z){\rm{d}}s$
+$\qquad+\int_{0}^{t}(\tilde{L}_{1}|Y_{s^{-}}^{\mu}-Y_{s^{-}}^{\nu}|^{2}+\tilde{L}_{2}\mathcal{W}_{2}^{2}(\mu_{s},\nu_{s})){\rm{d}}s.$
+Hence, similar to the term driven by $W$,
+$\mathbb{E}\Big[\sup_{0\leq r\leq t}\Big|\int_{0}^{r}\int_{\mathbb{R}^{d}_{0}}\langle Y_{s}^{\mu}-Y_{s}^{\nu},\;c(Y_{s}^{\mu},\mu_{s},z)-c(Y_{s}^{\nu},\nu_{s},z)\rangle\,\tilde{N}(\mathrm{d}s,\mathrm{d}z)\Big|\Big]$ (2.6)
+$\quad\leq\frac{1}{8}\,\mathbb{E}\Big[\sup_{0\leq r\leq t}|Y_{r}^{\mu}-Y_{r}^{\nu}|^{2}\Big]+K_{3}\int_{0}^{t}\Big(\mathbb{E}[\sup_{0\leq r\leq s}|Y_{r}^{\mu}-Y_{r}^{\nu}|^{2}]+\mathbb{E}[\sup_{0\leq r\leq s}\mathcal{W}_{2}^{2}(\mu_{r},\nu_{r})]\Big)\,\mathrm{d}s,$
+where $K_{3}$ depends on $\tilde{L}_{1},\tilde{L}_{2}$ and the moments $\int_{\mathbb{R}^{d}_{0}}|z|\nu({\rm{d}}z),\int_{\mathbb{R}^{d}_{0}}|z|^{2}\nu({\rm{d}}z),\int_{\mathbb{R}^{d}_{0}}|z|^{4}\nu({\rm{d}}z)$.
+Substituting (2.4), (2.5), (2.6) into (2.2), we obtain
 $\mathbb{E}\Big[\sup_{0\leq s\leq t}\big|Y_{s}^{\mu}-Y_{s}^{\nu}\big|^{2}\Big]$ (2.7)
-$\leq\mathbb{E}\Big{[}\!\sup_{0\leq r\leq t}\int_{0}^{r}\big{(}L_{1}|Y_{s}^{\mu}-Y_{s}^{\nu}|^{2}\big{)}\,\mathrm{d}s\Big{]}+\int_{0}^{t}\mathbb{E}[\sup_{0\leq r\leq s}(K_{1}+K_{2}+K_{3})|Y_{r}^{\mu}-Y_{r}^{\nu}|^{2}$
-$\quad+\sup_{0\leq r\leq s}(L_{2}+K_{1}+K_{2}+K_{3})\mathcal{W}_{2}^{2}(\mu_{r},\nu_{r})]{\rm{d}}s$
-$\leq\int_{0}^{t}\mathbb{E}[\sup_{0\leq r\leq s}(L_{1}+K_{1}+K_{2}+K_{3})|Y_{r}^{\mu}-Y_{r}^{\nu}|^{2}+\sup_{0\leq r\leq s}(L_{2}+K_{1}+K_{2}+K_{3})\mathcal{W}_{2}^{2}(\mu_{r},\nu_{r})]{\rm{d}}s$
-$\leq\hat{C}\int_{0}^{t}\left(\mathbb{E}[\sup_{0\leq r\leq s}|Y_{r}^{\mu}-Y_{r}^{\nu}|^{2}+\sup_{0\leq r\leq s}\mathcal{W}_{2}^{2}(\mu_{r},\nu_{r})]\right){\rm{d}}s,$
-where $\hat{C}:=\hat{C}(L_{1},L_{2},K_{1},K_{2},K_{3})$.
-Note that for any $s\geq 0$,
-$\sup_{0\leq r\leq s}|Y_{r}^{\mu}-Y_{r}^{\nu}|^{2}=\sup_{0\leq r\leq s}\left(e^{\lambda r}\cdot e^{-\lambda r}|Y_{r}^{\mu}-Y_{r}^{\nu}|^{2}\right)\leq e^{\lambda s}\sup_{0\leq r\leq s}e^{-\lambda r}|Y_{r}^{\mu}-Y_{r}^{\nu}|^{2}.$ (2.4)
-For convenience, we denote
-$\phi(t):=\mathbb{E}\left[\sup_{0\leq s\leq t}e^{-\lambda s}\left|Y_{s}^{\mu}-Y_{s}^{\nu}\right|^{2}\right].$
-Taking expectations to (2.4), we obtain
-$\mathbb{E}\left[\sup_{0\leq r\leq s}|Y_{r}^{\mu}-Y_{r}^{\nu}|^{2}\right]\leq e^{\lambda s}\,\phi(s).$
-Similarly,
-$\sup_{0\leq r\leq s}\mathcal{W}_{2}^{2}(\mu_{r},\nu_{r})\leq e^{\lambda s}\sup_{0\leq r\leq s}e^{-\lambda r}\mathcal{W}_{2}^{2}(\mu_{r},\nu_{r}).$
-Hence, according to (2.3) we have
-$\phi(t)$ $\leq\mathbb{E}\Big{[}\sup_{0\leq s\leq t}\big{|}Y_{s}^{\mu}-Y_{s}^{\nu}\big{|}^{2}\Big{]}$
-$\leq\hat{C}\int_{0}^{t}\left(\mathbb{E}[\sup_{0\leq r\leq s}|Y_{r}^{\mu}-Y_{r}^{\nu}|^{2}+\sup_{0\leq r\leq s}\mathcal{W}_{2}^{2}(\mu_{r},\nu_{r})]\right){\rm{d}}s$
-$\leq\hat{C}\int_{0}^{t}\left(e^{\lambda s}\,\phi(s)+e^{\lambda s}\mathbb{E}[\sup_{0\leq r\leq s}e^{-\lambda r}\mathcal{W}_{2}^{2}(\mu_{r},\nu_{r})]\right){\rm{d}}s$
-$\leq\hat{C}e^{\lambda T}\int_{0}^{t}\left(\,\phi(s)+\mathbb{E}[\sup_{0\leq r\leq s}e^{-\lambda r}\mathcal{W}_{2}^{2}(\mu_{r},\nu_{r})]\right){\rm{d}}s.$
-By using Gronwall’s inequality, we have
-$\phi(t)$ $\leq\hat{C}e^{\lambda T}\int_{0}^{t}\mathbb{E}[\sup_{0\leq r\leq s}e^{-\lambda r}\mathcal{W}_{2}^{2}(\mu_{r},\nu_{r})]e^{\hat{C}e^{\lambda T}(t-s)}{\rm{d}}s$
-$\leq\hat{C}e^{\lambda T}\cdot\mathbb{E}[\sup_{0\leq r\leq t}e^{-\lambda r}\mathcal{W}_{2}^{2}(\mu_{r},\nu_{r})]\int_{0}^{t}e^{\hat{C}e^{\lambda T}(t-s)}{\rm{d}}s$
-$=(e^{\hat{C}e^{\lambda T}t}-1)\mathbb{E}[\sup_{0\leq r\leq t}e^{-\lambda r}\mathcal{W}_{2}^{2}(\mu_{r},\nu_{r})].$
+$\leq\mathbb{E}\Big[\!\sup_{0\leq r\leq t}\int_{0}^{r}\big(L_{1}|Y_{s}^{\mu}-Y_{s}^{\nu}|^{2}+L_{2}\mathcal{W}_{2}^{2}(\mu_{s},\nu_{s})\big)\,\mathrm{d}s\Big]$
+$\quad+2\,\mathbb{E}\left[\sup_{0\leq r\leq t}\int_{0}^{r}\left\langle Y_{s}^{\mu}-Y_{s}^{\nu},(\sigma(Y_{s}^{\mu},\mu_{s})-\sigma(Y_{s}^{\nu},\nu_{s}))\mathrm{d}W_{s}\right\rangle\right]$
+$\quad+2\,\mathbb{E}\left[\sup_{0\leq r\leq t}\int_{0}^{r}\left\langle Y_{s}^{\mu}-Y_{s}^{\nu},(\sigma^{0}(Y_{s}^{\mu},\mu_{s})-\sigma^{0}(Y_{s}^{\nu},\nu_{s}))\mathrm{d}W_{s}^{0}\right\rangle\right]$
+$\quad+2\,\mathbb{E}\left[\sup_{0\leq r\leq t}\int_{0}^{r}\int_{\mathbb{R}_{0}^{d}}\Big(|Y_{s^{-}}^{\mu}-Y_{s^{-}}^{\nu}+\big(c(Y_{s^{-}}^{\mu},\mu_{s})-c(Y_{s^{-}}^{\nu},\nu_{s})\big)\,z|^{2}-|Y_{s^{-}}^{\mu}-Y_{s^{-}}^{\nu}|^{2}\Big)\widetilde{N}(\mathrm{d}s,\mathrm{d}z)\right]$
+$\leq\mathbb{E}\Big[\!\sup_{0\leq r\leq t}\int_{0}^{r}\big(L_{1}|Y_{s}^{\mu}-Y_{s}^{\nu}|^{2}+L_{2}\mathcal{W}_{2}^{2}(\mu_{s},\nu_{s})\big)\,\mathrm{d}s\Big]+\frac{3}{4}\,\mathbb{E}\Big[\sup_{0\leq r\leq t}|Y_{r}^{\mu}-Y_{r}^{\nu}|^{2}\Big]$
+$\quad+2(K_{1}+K_{2}+K_{3})\int_{0}^{t}\left(\mathbb{E}\Big[\sup_{0\leq r\leq s}|Y_{r}^{\mu}-Y_{r}^{\nu}|^{2}\Big]+\mathbb{E}\Big[\sup_{0\leq r\leq s}\mathcal{W}_{2}^{2}(\mu_{r},\nu_{r})\Big]\right)\,\mathrm{d}s$
+$\leq\frac{3}{4}\,\mathbb{E}\Big[\sup_{0\leq r\leq t}|Y_{r}^{\mu}-Y_{r}^{\nu}|^{2}\Big]+\frac{\hat{C}}{4}\int_{0}^{t}\left(\mathbb{E}[\sup_{0\leq r\leq s}|Y_{r}^{\mu}-Y_{r}^{\nu}|^{2}+\sup_{0\leq r\leq s}\mathcal{W}_{2}^{2}(\mu_{r},\nu_{r})]\right){\rm{d}}s,$
+where $\hat{C}:=\hat{C}(L_{1},L_{2},K_{1},K_{2},K_{3})$. Hence,
+$\mathbb{E}\Big[\sup_{0\leq s\leq t}\big|Y_{s}^{\mu}-Y_{s}^{\nu}\big|^{2}\Big]$ $\leq\hat{C}\int_{0}^{t}\left(\mathbb{E}[\sup_{0\leq r\leq s}|Y_{r}^{\mu}-Y_{r}^{\nu}|^{2}+\sup_{0\leq r\leq s}\mathcal{W}_{2}^{2}(\mu_{r},\nu_{r})]\right){\rm{d}}s$
+$\leq\hat{C}t\cdot\mathbb{E}[\sup_{0\leq r\leq t}|Y_{r}^{\mu}-Y_{r}^{\nu}|^{2}]+\hat{C}t\cdot\mathbb{E}[\sup_{0\leq r\leq t}\mathcal{W}_{2}^{2}(\mu_{r},\nu_{r})].$
 Therefore, we have
-$\phi(t)$ $=\mathbb{E}[\sup_{0\leq s\leq t}e^{-\lambda s}\left|Y_{s}^{\mu}-Y_{s}^{\nu}\right|^{2}]$
-$\leq(e^{\hat{C}e^{\lambda T}t}-1)\cdot\mathbb{E}[\sup_{0\leq r\leq t}e^{-\lambda r}\mathcal{W}_{2}^{2}(\mu_{r},\nu_{r})]$
-$\leq(e^{\hat{C}e^{\lambda T}t}-1)\cdot\mathbb{E}[\sup_{t\in[0,T]}e^{-\lambda t}|X_{t}-\tilde{X}_{t}|^{2}].$
-Combining the definition of the metric on space $(\mathscr{Y}_{T}(2),\bm{\rho}_{2,\infty})$, it holds that
+$\mathbb{E}\Big[\sup_{0\leq s\leq t}\big|Y_{s}^{\mu}-Y_{s}^{\nu}\big|^{2}\Big]$ $\leq\frac{\hat{C}t}{1-\hat{C}t}\cdot\mathbb{E}[\sup_{0\leq r\leq t}\mathcal{W}_{2}^{2}(\mu_{r},\nu_{r})]$
+$\leq\frac{\hat{C}t}{1-\hat{C}t}\cdot\mathbb{E}[\sup_{0\leq r\leq t}|\eta_{r}-\tilde{\eta}_{r}|^{2}].$
 […]
-$\leq(e^{\hat{C}e^{\lambda T}t}-1)^{1/2}\bigg{[}\mathbb{E}\sup_{t\in[0,T]}e^{-\lambda t}|X_{t}-\tilde{X}_{t}|^{2}\bigg{]}^{1/2}$
+$)\displaystyle\bm{\rho}_{2,\infty}(\Phi(X),\Phi(\tilde{X}$)) $=\bm{\rho}_{2,\infty}(Y,\tilde{Y}):=\bigg[\mathbb{E}\sup_{0\leq r\leq t}|Y_{r}-\tilde{Y}_{r}|^{2}\bigg]^{1/2}$
+$\leq\bigg(\frac{\hat{C}t}{1-\hat{C}t}\bigg)^{1/2}\bigg[\mathbb{E}\sup_{0\leq r\leq t}|X_{r}-\tilde{X}_{r}|^{2}\bigg]^{1/2}$
 […]
-Taking $\lambda=ln{(\frac{ln2}{2T\hat{C}})}^{\frac{1}{T}},$ then we have $0<(e^{\hat{C}e^{\lambda T}t}-1)^{1/2}<1$.
-By the contraction mapping principle, the mapping
+[…]
+By the contraction mapping principle on $0,t_{0}]$, the mapping
 $\Phi:\mathscr{Y}_{T}(2)\to\mathscr{Y}_{T}(2)$
@@
 Therefore, SDE (1.1) with initial distribution $X^{\mu}_{0}$ has a unique solution.
+We now iterate this argument to cover the whole interval $[0,T]$.
+Let $N:=\lceil T/t_{0}\rceil$ and for $k=1,\dots,N-1$ define the shifted time intervals $I_{k}:=[kt_{0},(k+1)t_{0}\wedge T]$.
+Suppose that a unique solution $X^{(k-1)}$ has been constructed on $0,kt_{0}]$.
+Use the terminal value $X^{(k-1)}_{kt_{0}}$ as the initial condition for the problem on the next interval $I_{k}$.
+Define the mapping $\Phi^{(k)}$ on the space $\mathscr{Y}_{I_{k}}(2)$ exactly as before but with initial time $kt_{0}$ and initial law induced by $X^{(k-1)}_{kt_{0}}$.
+The same local estimate (with the same constant $\hat{C}$) is valid on any interval of length $\leq t_{0}$, hence $\Phi^{(k)}$ is again a contraction on $\mathscr{Y}_{I_{k}}(2)$ with contraction constant $q<1$.
+Therefore there exists a unique fixed point $X^{(k)}$ on $I_{k}$.
+By constructing the solutions, we have $X^{(k)}_{kt_{0}}=X^{(k-1)}_{kt_{0}}$, so the pieces $\{X^{(k)}:k=0,\dots,N-1\}$ connect together to give a unique adapted solution $X$ on $[0,T]$ satisfying (1.1).
+Thus the existence on $[0,T]$ is proved.
   1. $\bf(4)$}
-    Proving the second-order moment boundedness
-Apply Itô’s formula to $|X_{t}|^{2}$. Using Assumption (A1)
-$2\langle x,b(x,\mu)\rangle+|\sigma(x,\mu)|^{2}+|\sigma^{0}(x,\mu)|^{2}+\int_{Z}|c(x,\mu)z|^{2}\nu(dz)\leq L(1+|x|^{2}+\mathcal{W}_{2}^{2}(\mu,\delta_{0})).$
-Taking expectation and noting $\mathbb{E}[\mathcal{W}_{2}^{2}(\mathcal{L}^{1}(X_{s}),\delta_{0})]\leq\mathbb{E}|X_{s}|^{2}$, we derive
+    Proving the boundedness of the second-order moment
+Apply Itô’s formula to $|X_{t}|^{2}$. Using Assumption (A1), taking expectation and noting
+$\mathbb{E}[\mathcal{W}_{2}^{2}(\mathcal{L}^{1}(X_{s}),\delta_{0})]\leq\mathbb{E}|X_{s}|^{2},$
+we derive
 $\mathbb{E}|X_{t}|^{2}\leq|x_{0}|^{2}+L\int_{0}^{t}(1+2\mathbb{E}|X_{s}|^{2}){\rm{d}}s.$
-By Gronwall’s inequality, we obtain that
+By Grönwall’s inequality, we obtain
 $\sup_{t\in[0,T]}\mathbb{E}|X_{t}|^{2}\leq\left(|x_{0}|^{2}+LT\right)e^{2LT}:=K.$
-This establishes existence, uniqueness, and the second-order moment boundedness. ∎
-Remark 2.2.
-Similar with Tran et al.[41]. It follows from Assumption (A6) that,
-$\langle x,b(x,\mu)\rangle+\frac{p_{0}-1}{2}(|\sigma(x,\mu)|^{2}+|\sigma^{0}(x,\mu)|^{2})+\frac{1}{2L_{3}}|c(x,\mu)|^{2}\int_{\mathbb{R}^{d}_{0}}|z|\left(\left(1+L_{3}|z|\right)^{p_{0}-1}-1\right)\nu({{\rm{d}}}z)$
-$\leq\gamma_{1}|x|^{2}+\gamma_{2}\mathcal{W}_{2}^{2}(\mu,\delta_{0})+\eta,$
-for any $p\in[2,p_{0}]$, $x\in\mathbb{R}^{d}$ and $\mu\in\mathcal{P}_{2}(\mathbb{R}^{d})$.
-Proposition 2.3.
-[…]
-$\mathbb{E}[|X_{t}|^{p}]\leq\begin{cases}C_{p}(1+e^{\gamma pt})&\text{if }\gamma\neq 0,\\ C_{p}(1+t)^{p/2}&\text{if }\gamma=0,p=2\text{ or }\gamma=0,\gamma_{2}>0,p\in(2,p_{0}],\\ C_{p}(1+t)^{p}&\text{if }\gamma=0,\gamma_{2}=0,p\in(2,p_{0}],\end{cases}$ (2.5)
-where $\gamma=\gamma_{1}+\gamma_{2}$.
-$\text{If }\gamma<0,\text{ then }\sup_{t\geq 0}\mathbb{E}\bigl{[}|X_{t}|^{p}\bigr{]}\leq 2C_{p}.$
-Proof.
-We hereby follow the methodology introduced in [41, Proposition 2.3 ].
-Step 1: It follows from Theorem 2.1 that $\sup_{t\in[0,T]}\mathbb{E}[|X_{t}|^{2}]\leq K,K:=K(\mathbb{E}|X^{\mu}_{0}|^{2},d,L,L_{1},\gamma_{1},\gamma_{2},T).$
-Step 2: We first demonstrate that for any even natural number $p\in[2,p_{0}]$ and for any $T>0$, the following holds,
-$\sup_{t\in[0,T]}\mathbb{E}[|X_{t}|^{p}]\leq C(T,p).$ (2.6)
-Observe that (2.6) is valid for the case $p=2$ as established in Step 1. Assume now, as an induction hypothesis, that (2.6) holds for every even integer $q\in[2,\,p-2]$, namely
-$\sup_{t\in[0,T]}\mathbb{E}\big{[}\,|X_{t}|^{q}\,\big{]}\leq C(T,q).$ (2.7)
-Let $\lambda\in\mathbb{R}$ and consider any even integer $p\in[2,\,p_{0}]$. By applying Itô’s formula to the process $e^{-\lambda t}|X_{t}|^{p}$, we obtain
-$e^{-\lambda t}|X_{t}|^{p}=|x_{0}|^{p}+\int_{0}^{t}e^{-\lambda s}\Big{[}-\lambda|X_{s}|^{p}+p|X_{s}|^{p-2}\langle X_{s},b(X_{s},\mathcal{L}^{1}_{X_{s}})\rangle+\frac{p}{2}|X_{s}|^{p-2}|\sigma(X_{s},\mathcal{L}^{1}_{X_{s}})|^{2}$ (2.8)
-$+\frac{p}{2}|X_{s}|^{p-2}|\sigma^{0}(X_{s},\mathcal{L}^{1}_{X_{s}})|^{2}+\frac{p(p-2)}{2}|X_{s}|^{p-4}\big{|}X_{s}^{\top}\sigma(X_{s},\mathcal{L}^{1}_{X_{s}})\big{|}^{2}$
-$+\frac{p(p-2)}{2}|X_{s}|^{p-4}\big{|}X_{s}^{\top}\sigma^{0}(X_{s},\mathcal{L}^{1}_{X_{s}})\big{|}^{2}\Big{]}{{\rm{d}}}s+p\int_{0}^{t}e^{-\lambda s}|X_{s}|^{p-2}\Big{[}\langle X_{s},\sigma(X_{s},\mathcal{L}^{1}_{X_{s}}){{\rm{d}}}W_{s}\rangle$
-$+\langle X_{s},\sigma^{0}(X_{s},\mathcal{L}^{1}_{X_{s}}){{\rm{d}}}W^{0}_{s}\rangle\Big{]}+\int_{0}^{t}\int_{\mathbb{R}_{0}^{d}}e^{-\lambda s}\Big{(}|X_{s^{-}}+c(X_{s^{-}},\mathcal{L}^{1}_{X_{s^{-}}})z|^{p}-|X_{s^{-}}|^{p}\Big{)}\widetilde{N}({{\rm{d}}}s,{{\rm{d}}}z)$
-$+\int_{0}^{t}\int_{\mathbb{R}_{0}^{d}}e^{-\lambda s}\Big{(}|X_{s}+c(X_{s},\mathcal{L}^{1}_{X_{s}})z|^{p}-|X_{s}|^{p}-p|X_{s}|^{p-2}\langle X_{s},c(X_{s},\mathcal{L}^{1}_{X_{s}})z\rangle\Big{)}\nu({{\rm{d}}}z){{\rm{d}}}s.$
-To handle the last integral in (2.8), it suffices to apply the binomial theorem, which yields, for any $t\geq 0$,
-$|X_{s}+c(X_{s},\mathcal{L}^{1}_{X_{s}})z|^{p}=\left(|X_{s}|^{2}+|c(X_{s},\mathcal{L}^{1}_{X_{s}})z|^{2}+2\langle X_{s},c(X_{s},\mathcal{L}^{1}_{X_{s}})z\rangle\right)^{p/2}$ (2.9)
-$=|X_{s}|^{p}+\frac{p}{2}|X_{s}|^{p-2}\left(|c(X_{s},\mathcal{L}^{1}_{X_{s}})z|^{2}+2\langle X_{s},c(X_{s},\mathcal{L}^{1}_{X_{s}})z\rangle\right)$
-$\quad+\sum_{i=2}^{p/2}\binom{p/2}{i}|X_{s}|^{p-2i}\left(|c(X_{s},\mathcal{L}^{1}_{X_{s}})z|^{2}+2\langle X_{s},c(X_{s},\mathcal{L}^{1}_{X_{s}})z\rangle\right)^{i}$
-$=|X_{s}|^{p}+p|X_{s}|^{p-2}\langle X_{s},c(X_{s},\mathcal{L}^{1}_{X_{s}})z\rangle+\frac{p}{2}|X_{s}|^{p-2}|c(X_{s},\mathcal{L}^{1}_{X_{s}})z|^{2}$
-$\quad+\sum_{i=2}^{p/2}\binom{p/2}{i}|X_{s}|^{p-2i}\left(|c(X_{s},\mathcal{L}^{1}_{X_{s}})z|^{2}+2\langle X_{s},c(X_{s},\mathcal{L}^{1}_{X_{s}})z\rangle\right)^{i}.$
+This establishes existence, uniqueness, and the boundedness of the second-order moment. ∎
+We will need the following Lemma later.
 Lemma 2.4.
 For any $s\geq 0$, we have
-$\mathcal{W}^{2}_{2}(\mathcal{L}^{1}_{X_{s}},\delta_{0})=\mathbb{E}[|X_{s}|^{2}\mid\mathcal{F}^{0}],$
+$\mathcal{W}^{2}_{2}(\mathcal{L}^{1}_{X_{s}},\delta_{0})=\mathbb{E}[|X_{s}|^{2}\mid\mathcal{F}^{0}],$ (2.8)
 where $\mathcal{L}^{1}_{X_{s}}$ denotes the conditional law of $X_{s}$ given $\mathcal{F}^{0}$.
@@
 This completes the proof. ∎
-Next, by repeatedly applying the binomial theorem, using Assumption (A5), the identity
-$\mathcal{W}^{2}_{2}(\mathcal{L}^{1}_{X_{s}},\delta_{0})=\mathbb{E}[|X_{s}|^{2}\mid\mathcal{F}^{0}],$
-and Young’s inequality
+Proposition 2.5.
+Let assumptions (A1)-(A7) hold. Based on Theorem 2.3, we can further obtain for each $p\in[2,p_{0}]$, one can find a constant $C_{p}>0$ such that for all $t\geq 0$,
+$\mathbb{E}[|X_{t}|^{p}]\leq\begin{cases}C_{p}(1+e^{\gamma pt})&\text{if }\gamma\neq 0,\\ C_{p}(1+t)^{p/2}&\text{if }\gamma=0,p=2\text{ or }\gamma=0,\gamma_{2}>0,p\in(2,p_{0}],\\ C_{p}(1+t)^{p}&\text{if }\gamma=0,\gamma_{2}=0,p\in(2,p_{0}],\end{cases}$ (2.9)
+where $\gamma=\gamma_{1}+\gamma_{2}$.
+$\text{If }\gamma<0,\text{ then }\sup_{t\geq 0}\mathbb{E}\bigl[|X_{t}|^{p}\bigr]\leq 2C_{p}.$
+Proof.
+We hereby follow the methodology introduced in [32, Proposition 2.3 ].
+Step 1: It follows from Theorem 2.3 that $\sup_{t\in[0,T]}\mathbb{E}[|X_{t}|^{2}]\leq K,K:=K(\mathbb{E}|X^{\mu}_{0}|^{2},d,L,L_{1},\gamma_{1},\gamma_{2},T).$
+Step 2: We first demonstrate that for any even natural number $p\in[2,p_{0}]$ and for any $T>0$, the following holds,
+$\sup_{t\in[0,T]}\mathbb{E}[|X_{t}|^{p}]\leq C(T,p).$ (2.10)
+Observe that (2.10) is valid for the case $p=2$ as established in Step 1. Assume now, as an induction hypothesis, that (2.10) holds for every even integer $q\in[2,\,p-2]$, namely
+$\sup_{t\in[0,T]}\mathbb{E}\big[\,|X_{t}|^{q}\,\big]\leq C(T,q).$ (2.11)
+Let $\lambda\in\mathbb{R}$ and consider any even integer $p$. By applying Itô’s formula to the process $e^{-\lambda t}|X_{t}|^{p}$, we obtain
+$e^{-\lambda t}|X_{t}|^{p}=|x_{0}|^{p}+\int_{0}^{t}e^{-\lambda s}\Big[-\lambda|X_{s}|^{p}+p|X_{s}|^{p-2}\langle X_{s},b(X_{s},\mathcal{L}^{1}_{X_{s}})\rangle+\frac{p}{2}|X_{s}|^{p-2}|\sigma(X_{s},\mathcal{L}^{1}_{X_{s}})|^{2}$ (2.12)
+$+\frac{p}{2}|X_{s}|^{p-2}|\sigma^{0}(X_{s},\mathcal{L}^{1}_{X_{s}})|^{2}+\frac{p(p-2)}{2}|X_{s}|^{p-4}\big|X_{s}^{\top}\sigma(X_{s},\mathcal{L}^{1}_{X_{s}})\big|^{2}$
+$+\frac{p(p-2)}{2}|X_{s}|^{p-4}\big|X_{s}^{\top}\sigma^{0}(X_{s},\mathcal{L}^{1}_{X_{s}})\big|^{2}\Big]{{\rm{d}}}s+p\int_{0}^{t}e^{-\lambda s}|X_{s}|^{p-2}\Big[\langle X_{s},\sigma(X_{s},\mathcal{L}^{1}_{X_{s}}){{\rm{d}}}W_{s}\rangle$
+$+\langle X_{s},\sigma^{0}(X_{s},\mathcal{L}^{1}_{X_{s}}){{\rm{d}}}W^{0}_{s}\rangle\Big]+\int_{0}^{t}\int_{\mathbb{R}_{0}^{d}}e^{-\lambda s}\Big(|X_{s^{-}}+c(X_{s^{-}},\mathcal{L}^{1}_{X_{s^{-}}})z|^{p}-|X_{s^{-}}|^{p}\Big)\widetilde{N}({{\rm{d}}}s,{{\rm{d}}}z)$
+$+\int_{0}^{t}\int_{\mathbb{R}_{0}^{d}}e^{-\lambda s}\Big(|X_{s}+c(X_{s},\mathcal{L}^{1}_{X_{s}})z|^{p}-|X_{s}|^{p}-p|X_{s}|^{p-2}\langle X_{s},c(X_{s},\mathcal{L}^{1}_{X_{s}})z\rangle\Big)\nu({{\rm{d}}}z){{\rm{d}}}s.$
+To handle the last integral in (2.12), it suffices to apply the binomial theorem, which yields, for any $t\geq 0$,
+$|X_{s}+c(X_{s},\mathcal{L}^{1}_{X_{s}})z|^{p}=\left(|X_{s}|^{2}+|c(X_{s},\mathcal{L}^{1}_{X_{s}})z|^{2}+2\langle X_{s},c(X_{s},\mathcal{L}^{1}_{X_{s}})z\rangle\right)^{p/2}$ (2.13)
+$=|X_{s}|^{p}+\frac{p}{2}|X_{s}|^{p-2}\left(|c(X_{s},\mathcal{L}^{1}_{X_{s}})z|^{2}+2\langle X_{s},c(X_{s},\mathcal{L}^{1}_{X_{s}})z\rangle\right)$
+$\quad+\sum_{i=2}^{p/2}\binom{p/2}{i}|X_{s}|^{p-2i}\left(|c(X_{s},\mathcal{L}^{1}_{X_{s}})z|^{2}+2\langle X_{s},c(X_{s},\mathcal{L}^{1}_{X_{s}})z\rangle\right)^{i}$
+$=|X_{s}|^{p}+p|X_{s}|^{p-2}\langle X_{s},c(X_{s},\mathcal{L}^{1}_{X_{s}})z\rangle+\frac{p}{2}|X_{s}|^{p-2}|c(X_{s},\mathcal{L}^{1}_{X_{s}})z|^{2}$
+$\quad+\sum_{i=2}^{p/2}\binom{p/2}{i}|X_{s}|^{p-2i}\left(|c(X_{s},\mathcal{L}^{1}_{X_{s}})z|^{2}+2\langle X_{s},c(X_{s},\mathcal{L}^{1}_{X_{s}})z\rangle\right)^{i}.$
+Next, by repeatedly applying the binomial theorem, using Assumption (A5), the identity (2.4) and Young’s inequality
 $y|x|^{p-3}\leq\frac{1}{2}(|x|^{p-2}+y^{2}|x|^{p-4}),$
 which holds for any $x\in\mathbb{R}^{d}$, $y>0$, as well as the Binomial identities
-$\sum_{j=0}^{i}\binom{i}{j}a^{j}=(1+a)^{i},\quad\sum_{j=0}^{i}\binom{i}{j}ja^{j}=ia(1+a)^{i-1}$
-valid for any $a\in\mathbb{R}$, we obtain,
+$\sum_{j=0}^{i}\binom{i}{j}a^{j}=(1+a)^{i}\quad\textup{and}\quad\sum_{j=0}^{i}\binom{i}{j}ja^{j}=ia(1+a)^{i-1},$
+which valid for any $a\in\mathbb{R}$, we obtain,
 $|X_{s}|^{p-2i}\Big(|c(X_{s},\mathcal{L}^{1}_{X_{s}})z|^{2}+2\langle X_{s},c(X_{s},\mathcal{L}^{1}_{X_{s}})z\rangle\Big)^{i}$
@@
 $\leq|c(X_{s},\mathcal{L}^{1}_{X_{s}})|^{2}|X_{s}|^{p-2}\left(\frac{i}{L_{3}^{2}}(L_{3}^{2}|z|^{2}+2L_{3}|z|)^{i}-\frac{|z|i}{L_{3}}(L_{3}^{2}|z|^{2}+2L_{3}|z|)^{i-1}\right)$
-$\quad+{L_{3}^{2}}\Big{(}1+|X_{s}|+\sqrt{\mathbb{E}[|X_{s}|^{2}|\mathcal{F}^{0}]}\Big{)}^{2}\sum_{j=0}^{i}\binom{i}{j}2^{j}|z|^{2i-j}L_{3}^{2i-j-2}\Bigg{(}(i-\frac{j}{2}-1)(1+\sqrt{\mathbb{E}[|X_{s}|^{2}|\mathcal{F}^{0}]})^{2}|X_{s}|^{p-4}$
-$\quad+\sum_{k=2}^{2i-j-2}\binom{2i-j-2}{k}(1+\sqrt{\mathbb{E}[|X_{s}|^{2}|\mathcal{F}^{0}]})^{k}|X_{s}|^{p-2-k}\Bigg{)}$
+$\quad+{L_{3}^{2}}\Big(1+|X_{s}|+\sqrt{\mathbb{E}[|X_{s}|^{2}|\mathcal{F}^{0}]}\Big)^{2}\sum_{j=0}^{i}\binom{i}{j}2^{j}|z|^{2i-j}L_{3}^{2i-j-2}\Bigg((i-\frac{j}{2}-1)(1+\sqrt{\mathbb{E}[|X_{s}|^{2}|\mathcal{F}^{0}]})^{2}$
+$\quad\times|X_{s}|^{p-4}+\sum_{k=2}^{2i-j-2}\binom{2i-j-2}{k}(1+\sqrt{\mathbb{E}[|X_{s}|^{2}|\mathcal{F}^{0}]})^{k}|X_{s}|^{p-2-k}\Bigg)$
 $=|c(X_{s},\mathcal{L}^{1}_{X_{s}})|^{2}|X_{s}|^{p-2}\frac{i|z|}{L_{3}}(1+L_{3}|z|)(L_{3}^{2}|z|^{2}+2L_{3}|z|)^{i-1}$
-$\quad+\sum_{j=0}^{i}|z|^{2i-j}Q_{p}\big{(}p-2,2i-j,|X_{s}|,1+\sqrt{\mathbb{E}[|X_{s}|^{2}|\mathcal{F}^{0}]}\big{)}$
-$where\quad 2i-j\geq 2\quad\text{and}\quad$
+$\quad+\sum_{j=0}^{i}|z|^{2i-j}Q_{p}\big(p-2,2i-j,|X_{s}|,1+\sqrt{\mathbb{E}[|X_{s}|^{2}|\mathcal{F}^{0}]}\big),$
+$\text{where}\quad 2i-j\geq 2\quad\text{and}\quad$
 $Q_{q}(m,n,x,y)=\sum_{\ell_{1}\leq m,\,\ell_{2}\leq n,\,\ell_{1}+\ell_{2}=q}c_{\ell_{1}\ell_{2}}x^{\ell_{1}}y^{\ell_{2}}.$
-Combined with the identity
+This, together with the identity
 $\sum_{i=2}^{p/2}\binom{p/2}{i}ia^{i-1}=\frac{p}{2}\left((1+a)^{p/2-1}-1\right),$
-which holds for any $a\in\mathbb{R}$, this yields,
+which holds for any $a\in\mathbb{R}$, yields,
 $\sum_{i=2}^{p/2}\binom{p/2}{i}|X_{s}|^{p-2i}\left(|c(X_{s},\mathcal{L}^{1}_{X_{s}})z|^{2}+2\langle X_{s},c(X_{s},\mathcal{L}^{1}_{X_{s}})z\rangle\right)^{i}$ (2.14)
@@
 $\tau_{N}:=\inf\{t\geq 0:|X_{t}|\geq N\}.$
-By choosing $\lambda=\gamma_{1}p$ and applying Assumption (A6), Remark 2.2, together with the identity
+By choosing $\lambda=\gamma_{1}p$ and applying Assumption (A6), Remark 2.1, together with Lemma 2.4
 $\mathcal{W}_{2}^{2}(\mathcal{L}^{1}_{X_{s}},\delta_{0})=\mathbb{E}[|X_{s}|^{2}\mid\mathcal{F}^{0}],$
@@
 $\sup_{t\in[0,T]}\mathbb{E}\left[|X_{t}|^{p}\right]\leq C(T,p).$
-Thus, by the principle of induction, we have established (2.6).
+Thus, by the principle of induction, we establish (2.10).
 Step 3: We next aim to prove (2.9) for every even natural number $p\in[2,p_{0}]$.
-First, by applying (2.8) with $\lambda=2\gamma$ and $p=2$, and using Assumption (A6), we obtain,
-$e^{-2\gamma t}|X_{t}|^{2}=|x_{0}|^{2}+\int_{0}^{t}e^{-2\gamma s}\bigg{[}-2\gamma|X_{s}|^{2}+2\langle X_{s},b(X_{s},\mathcal{L}^{1}_{X_{s}})\rangle+|\sigma(X_{s},\mathcal{L}^{1}_{X_{s}})|^{2}+|\sigma^{0}(X_{s},\mathcal{L}^{1}_{X_{s}})|^{2}\bigg{]}{{\rm{d}}}s$
-$\quad+2\int_{0}^{t}e^{-2\gamma s}\bigg{[}\langle X_{s},\sigma(X_{s},\mathcal{L}^{1}_{X_{s}}){{\rm{d}}}W_{s}\rangle+\langle X_{s},\sigma^{0}(X_{s},\mathcal{L}^{1}_{X_{s}}){{\rm{d}}}W^{0}_{s}\rangle\Bigg{]}$
-$\quad+\int_{0}^{t}\int_{\mathbb{R}^{d}_{0}}e^{-2\lambda s}\Big{(}|X_{s-}+c(X_{s-},\mathcal{L}^{1}_{X_{s}-})z|^{2}-|X_{s-}|^{2}\Big{)}\tilde{N}({{\rm{d}}}s,{{\rm{d}}}z)$
-$\quad+\int_{0}^{t}\int_{\mathbb{R}^{d}_{0}}e^{-2\gamma s}|c(X_{s},\mathcal{L}^{1}_{X_{s}})z|^{2}\nu({{\rm{d}}}z){{\rm{d}}}s$
-$\leq|x_{0}|^{2}+2\int_{0}^{t}e^{-2\gamma s}\bigg{[}-\gamma|X_{s}|^{2}+\langle X_{s},b(X_{s},\mathcal{L}^{1}_{X_{s}})\rangle+\frac{1}{2}|\sigma(X_{s},\mathcal{L}^{1}_{X_{s}})|^{2}+\frac{1}{2}|\sigma^{0}(X_{s},\mathcal{L}^{1}_{X_{s}})|^{2}$
-$\quad+\frac{1}{2}|c(X_{s},\mathcal{L}^{1}_{X_{s}})|^{2}\int_{\mathbb{R}^{d}_{0}}|z|^{2}\nu({{\rm{d}}}z)\bigg{]}{{\rm{d}}}s$
-$\quad+2\int_{0}^{t}e^{-2\gamma s}\bigg{[}\langle X_{s},\sigma(X_{s},\mathcal{L}^{1}_{X_{s}}){{\rm{d}}}W_{s}\rangle+\langle X_{s},\sigma^{0}(X_{s},\mathcal{L}^{1}_{X_{s}}){{\rm{d}}}W_{s}^{0}\rangle\bigg{]}$
-$\quad+\int_{0}^{t}\int_{\mathbb{R}^{d}_{0}}e^{-2\gamma s}\Big{(}|X_{s-}+c(X_{s-},\mathcal{L}^{1}_{X_{s}-})z|^{2}-|X_{s-}|^{2}\Big{)}\tilde{N}({{\rm{d}}}s,{{\rm{d}}}z)$
-$\leq|x_{0}|^{2}+2\int_{0}^{t}e^{-2\gamma s}\left(-\gamma_{2}|X_{s}|^{2}+\gamma_{2}\mathcal{W}_{2}^{2}(\mathcal{L}^{1}_{X_{s}},\delta_{0})+\eta\right){{\rm{d}}}s$
+First, by applying (2.12) with $\lambda=2\gamma$ and $p=2$, and using Assumption (A6), we obtain
+$e^{-2\gamma t}|X_{t}|^{2}\leq|x_{0}|^{2}+2\int_{0}^{t}e^{-2\gamma s}\left(-\gamma_{2}|X_{s}|^{2}+\gamma_{2}\mathcal{W}_{2}^{2}(\mathcal{L}^{1}_{X_{s}},\delta_{0})+\eta\right){{\rm{d}}}s$
 $\quad+2\int_{0}^{t}e^{-2\gamma s}\bigg[\langle X_{s},\sigma(X_{s},\mathcal{L}^{1}_{X_{s}}){{\rm{d}}}W_{s}+\sigma^{0}(X_{s},\mathcal{L}^{1}_{X_{s}}){{\rm{d}}}W_{s}^{0}\rangle\bigg]$
 $\quad+\int_{0}^{t}\int_{\mathbb{R}^{d}_{0}}e^{-2\gamma s}\Big(|X_{s-}+c(X_{s-},\mathcal{L}^{1}_{X_{s}-})z|^{2}-|X_{s-}|^{2}\Big)\tilde{N}({{\rm{d}}}s,{{\rm{d}}}z).$
-From the identity
-$\mathcal{W}_{2}^{2}\big{(}\mathcal{L}^{1}_{X_{s}},\delta_{0}\big{)}=\mathbb{E}\big{[}\,|X_{s}|^{2}\,\big{|}\,\mathcal{F}^{0}\big{]},$
-and by applying the estimate (2.6), we deduce
+Due to Lemma 2.4, we have $\mathcal{W}_{2}^{2}\big(\mathcal{L}^{1}_{X_{s}},\delta_{0}\big)=\mathbb{E}\big[\,|X_{s}|^{2}\,\big|\,\mathcal{F}^{0}\big],$ and by applying the estimate (2.10), we deduce
 $\mathbb{E}\!\left[e^{-2\gamma t}|X_{t}|^{2}\right]\leq|x_{0}|^{2}+2\int_{0}^{t}e^{-2\gamma s}\mathbb{E}\!\left[-\gamma_{2}|X_{s}|^{2}+\gamma_{2}\mathbb{E}[|X_{s}|^{2}\mid\mathcal{F}^{0}]+\eta\right]\mathrm{d}s\leq|x_{0}|^{2}+2\eta\int_{0}^{t}e^{-2\gamma s}\,\mathrm{d}s.$
 This yields
-$\mathbb{E}[|X_{t}|^{2}]\leq\begin{cases}\big{(}|x_{0}|^{2}+\tfrac{\eta}{\gamma}\big{)}e^{2\gamma t}-\tfrac{\eta}{\gamma},&\gamma\neq 0,\\[3.99994pt] |x_{0}|^{2}+2\eta t,&\gamma=0,\end{cases}$
-so that (2.5) holds in the case $p=2$. Assume now, as an induction hypothesis, that (2.5) is valid for every even integer $q\in[2,\,p-2]$, that is,
-$\mathbb{E}[|X_{t}|^{q}]\leq\begin{cases}C_{q}\big{(}1+e^{\gamma qt}\big{)},&\gamma\neq 0,\\[3.99994pt] C_{q}(1+t)^{q/2},&\gamma=0,\ \big{(}q=2\ \text{or}\ (\gamma_{2}>0,\ q\in(2,p-2])\big{)},\\[3.99994pt] C_{q}(1+t)^{q},&\gamma=0,\ \gamma_{2}=0,\ q\in(2,p-2].\end{cases}$ (2.14)
+$\mathbb{E}[|X_{t}|^{2}]\leq\begin{cases}\big(|x_{0}|^{2}+\tfrac{\eta}{\gamma}\big)e^{2\gamma t}-\tfrac{\eta}{\gamma}&\gamma\neq 0,\\[3.99994pt] |x_{0}|^{2}+2\eta t&\gamma=0,\end{cases}$
+so that (2.9) holds in the case $p=2$. Assume now, as an induction hypothesis, (2.9) is valid for every even integer $q\in[2,\,p-2]$, that is,
+$\mathbb{E}[|X_{t}|^{q}]\leq\begin{cases}C_{q}\big(1+e^{\gamma qt}\big)&\gamma\neq 0,\\[3.99994pt] C_{q}(1+t)^{q/2}&\gamma=0,\ \big(q=2\ \text{or}\ (\gamma_{2}>0,\ q\in(2,p-2])\big),\\[3.99994pt] C_{q}(1+t)^{q}&\gamma=0,\ \gamma_{2}=0,\ q\in(2,p-2].\end{cases}$ (2.18)
 We prove that (2.9) also holds for the even integer $p$. The argument relies on (2.16), the inductive assumption (2.18), and Assumption (A4).
@@
 Letting $N\to\infty$ yields $\mathbb{E}[|X_{t}|^{p}]\leq C(1+t)^{p}$.
-Therefore, (2.5) holds for the given even $p$.
+Therefore, (2.9) holds for the given even pp.
 By induction, it follows that (2.9) is valid for all even integers $p\in[2,\,p_{0}]$, and by Hölder’s inequality, for all real $p\in[2,\,p_{0}]$.
 This completes the proof.
+∎
 3 Propagation of chaos
 For each $N\in\mathbb{N}$, let $(W^{i},Z^{i})$, $i=1,\ldots,N$, be independent copies of the pair $(W,Z)$.
-Denote by $N^{i}(dt,dz)$ the Poisson random measure corresponding to the jumps of the Lévy process $Z^{i}$ with intensity measure $\nu(dz)dt$, and define the compensated Poisson random measure by
+Denote by $N^{i}({\rm{d}}t,{\rm{d}}z)$ the Poisson random measure corresponding to the jumps of the Lévy process $Z^{i}$ with intensity measure $\nu({\rm{d}}z){\rm{d}}t$, and define the compensated Poisson random measure by
 $\widetilde{N}^{i}({\rm{d}}t,{\rm{d}}z):=N^{i}({\rm{d}}t,{\rm{d}}z)-\nu({\rm{d}}z){\rm{d}}t.$
@@
 $Z_{t}^{i}=\int_{0}^{t}\int_{\mathbb{R}_{0}^{d}}z\,\widetilde{N}^{i}({\rm{d}}s,{\rm{d}}z),\quad t\geq 0.$
-We now consider the system of non-interacting particles associated with the Lévy-driven McKean-Vlasov SDE with common noise (1.1), where the state $X^{i}=(X_{t}^{i})_{t\geq 0}$ of the $i$-th particle is defined by
+We now consider the system of non-interacting particles associated with the McKean-Vlasov SDE driven by Lévy processes with common noise (1.1), where the state $X^{i}=(X_{t}^{i})_{t\geq 0}$ of the $i$-th particle is defined by
 $\begin{split}X_{t}^{i}&=x_{0}+\int_{0}^{t}b(X_{s}^{i},\mathcal{L}^{1}_{X_{s}^{i}}){{\rm{d}}}s+\int_{0}^{t}\sigma(X_{s}^{i},\mathcal{L}^{1}_{X_{s}^{i}}){{\rm{d}}}W_{s}^{i}+\int_{0}^{t}\sigma^{0}(X_{s}^{i},\mathcal{L}^{1}_{X_{s}^{i}}){{\rm{d}}}W^{0}_{s}\\ &+\int_{0}^{t}\int_{\mathbb{R}_{0}^{d}}c\left(X_{s-}^{i},\mathcal{L}^{1}_{X_{s-}^{i}}\right)z\widetilde{N}^{i}({{\rm{d}}}s,{{\rm{d}}}z),\end{split}$ (3.1)
 for any $t\geq 0$ and $i\in\{1,\ldots,N\}$.
-For $\bm{x}^{N}:=(x_{1},x_{2},\ldots,x_{N}),\bm{y}^{N}:=(y_{1},y_{2},\ldots,y_{N})\in\mathbb{R}^{dN}$, we have
-$\mathcal{W}_{2}^{2}(\mu^{\bm{x}^{N}},\delta_{0})=\frac{1}{N}\sum_{i=1}^{N}|x_{i}|^{2}.$
+For $\bm{X}^{N}:=(X^{1},X^{2},\ldots,X^{N}),\bm{Y}^{N}:=(Y^{1},Y^{2},\ldots,Y^{N})\in\mathbb{R}^{dN}$, we have
+$\mathcal{W}_{2}^{2}(\mu^{\bm{X}^{N}},\delta_{0})=\frac{1}{N}\sum_{i=1}^{N}|X^{i}|^{2}.$
 Here, the empirical measure is defined by
-$\mu^{\bm{x}^{N}}({\rm{d}}x):=\frac{1}{N}\sum_{i=1}^{N}\delta_{x_{i}}({\rm{d}}x),$
+$\mu^{\bm{X}^{N}}({\rm{d}}x):=\frac{1}{N}\sum_{i=1}^{N}\delta_{X^{i}}({\rm{d}}x),$ (3.2)
 where $\delta_{X}$ denotes the Dirac measure concentrated at $X$.
 Moreover, a standard estimate for the Wasserstein distance between two empirical measures $\mu^{\bm{X}^{N}}$ and $\mu^{\bm{Y}^{N}}$ is given by
-$\mathcal{W}_{2}^{2}(\mu^{\bm{x}^{N}},\mu^{\bm{y}^{N}})\leq\frac{1}{N}\sum_{i=1}^{N}|x_{i}-y_{i}|^{2}=\frac{1}{N}\left|\bm{x}^{N}-\bm{y}^{N}\right|^{2},$
-(see inequality (1.24) in [18]). The true measure $\mathcal{L}_{X_{t}}$ at each time $t$ is approximated by the empirical measure
-$\mu_{t}^{\bm{X}^{N}}({\rm{d}}x):=\frac{1}{N}\sum_{i=1}^{N}\delta_{X_{t}^{i,N}}({\rm{d}}x),$
+$\mathcal{W}_{2}^{2}(\mu^{\bm{X}^{N}},\mu^{\bm{Y}^{N}})\leq\frac{1}{N}\sum_{i=1}^{N}|X^{i}-Y^{i}|^{2}=\frac{1}{N}\left|\bm{X}^{N}-\bm{Y}^{N}\right|^{2},$ (3.3)
+(see inequality (1.24) in [12]). The true measure $\mathcal{L}^{1}_{X_{t}}$ at each time $t$ is approximated by the empirical measure
+$\mu_{t}^{\bm{X}^{N}}({\rm{d}}x):=\frac{1}{N}\sum_{i=1}^{N}\delta_{X_{t}^{i,N}}({\rm{d}}x),$ (3.4)
 where $\bm{X}^{N}=(\bm{X}_{t}^{N})_{t\geq 0}=(X_{t}^{1,N},\ldots,X_{t}^{N,N})_{t\geq 0}^{\top}$ denotes the system of interacting particles, which is the solution to an $\mathbb{R}^{dN}$-valued SDE driven by Lévy processes with common noise.
@@
 Proposition 3.1.
-(see [41]) For each $p\in[2,p_{0}]$, there exists a constant $C_{p}>0$ such that, for all $t\geq 0$,
-$\max_{1\leq i\leq N}\mathbb{E}\!\left[|X_{t}^{i,N}|^{p}\right]\leq\begin{cases}C_{p}\big{(}1+e^{\gamma pt}\big{)},&\gamma\neq 0,\\[3.00003pt] C_{p}(1+t)^{p/2},&\gamma=0,\ p=2\ \text{or}\ (\gamma=0,\ \gamma_{2}>0,\ p\in(2,p_{0}]),\\[3.00003pt] C_{p}(1+t)^{p},&\gamma=0,\ \gamma_{2}=0,\ p\in(2,p_{0}],\end{cases}$
+([32, Proposition 3.1]) Let assumptions (A1)-(A7) hold, for each $p\in[2,p_{0}]$, there exists a constant $C_{p}>0$ such that, for all $t\geq 0$,
+$\max_{1\leq i\leq N}\mathbb{E}\!\left[|X_{t}^{i,N}|^{p}\right]\leq\begin{cases}C_{p}\big(1+e^{\gamma pt}\big)&\gamma\neq 0,\\[3.00003pt] C_{p}(1+t)^{p/2}&\gamma=0,\ p=2\ \text{or}\ (\gamma=0,\ \gamma_{2}>0,\ p\in(2,p_{0}]),\\[3.00003pt] C_{p}(1+t)^{p}&\gamma=0,\ \gamma_{2}=0,\ p\in(2,p_{0}],\end{cases}$
 where $\gamma:=\gamma_{1}+\gamma_{2}$. In the special case $\gamma<0$, one further has
@@
 Proof.
-The argument follows the same reasoning as in the proof of Proposition 2.3 and is therefore not repeated here. ∎
-Next, we state a result concerning the propagation of chaos, which plays a central role in the convergence analysis as $N\to\infty$. For clarity, we first introduce the following definition,
-$\varphi(N)=\begin{cases}N^{-1/2}&\text{if }d<4,\\ N^{-1/2}\ln N&\text{if }d=4,\\ N^{-2/d}&\text{if }d>4,\end{cases}$ (3.3)
-(see Theorem 1 of [21]).
+The argument follows the same reasoning as in the proof of Proposition 2.5 and is therefore omitted here. ∎
 Proposition 3.2.
-Assume that all conditions in Proposition 3.1 hold and that Assumption (A2) holds for $L_{1}\in\mathbb{R}$, $L_{2}\geq 0$. Then, we have
-$\max_{i\in\{1,\ldots,N\}}\sup_{t\in[0,T]}\mathbb{E}\left[|X_{t}^{i}-X_{t}^{i,N}|^{2}\right]\leq C_{T}\varphi(N),$ (3.4)
-for any $N\in\mathbb{N}$, where the positive constant $C_{T}$ does not depend on $N$.
-Assume further that $L_{1}+L_{2}<0$ and $\gamma=\gamma_{1}+\gamma_{2}<0$. Then, we have
-$\max_{i\in\{1,\ldots,N\}}\sup_{t\geq 0}\mathbb{E}\left[|X_{t}^{i}-X_{t}^{i,N}|^{2}\right]\leq C\varphi(N),$ (3.5)
-for any $N\in\mathbb{N}$, where the positive constant $C$ does not depend on $N$ and $T$.
+Define the rate function
+$\varphi(N)=\begin{cases}N^{-1/2}&\text{for }d<4,\\ N^{-1/2}\ln N&\text{for }d=4,\\ N^{-2/d}&\text{for }d>4.\end{cases}$ (3.6)
+Under the conditions of Proposition 3.1 and Assumption (A2) with $L_{1}\in\mathbb{R}$, $L_{2}\geq 0$, there exists a constant $C_{T}>0$, independent of $N$, such that for all $N\in\mathbb{N}$,
+$\max_{i\in\{1,\ldots,N\}}\sup_{t\in[0,T]}\mathbb{E}\left[|X_{t}^{i}-X_{t}^{i,N}|^{2}\right]\leq C_{T}\varphi(N).$ (3.7)
+Furthermore, if $L_{1}+L_{2}<0$ and $\gamma<0$, then the constant can be made independent of time, i.e., there exists $C>0$, independent of $N$ and $T$, such that
+$\max_{i\in\{1,\ldots,N\}}\sup_{t\geq 0}\mathbb{E}\left[|X_{t}^{i}-X_{t}^{i,N}|^{2}\right]\leq C\varphi(N).$ (3.8)
 Proof.
@@
 $X_{t}^{i}-X_{t}^{i,N}=\int_{0}^{t}\left(b(X_{s}^{i},\mathcal{L}^{1}_{X_{s}^{i}})-b(X_{s}^{i,N},\mu_{s}^{\bm{X}^{N}})\right){{\rm{d}}}s+\int_{0}^{t}\left(\sigma(X_{s}^{i},\mathcal{L}^{1}_{X_{s}^{i}})-\sigma(X_{s}^{i,N},\mu_{s}^{\bm{X}^{N}})\right){\rm{d}}W_{s}^{i}$
-$\quad+\int_{0}^{t}\left(\sigma^{0}(X_{s}^{i},\mathcal{L}^{1}_{X_{s}^{i}})-\sigma^{0}(X_{s}^{i,N},\mu_{s}^{\bm{X}^{N}})\right){\rm{d}}W_{s}^{0}$
-$\quad+\int_{0}^{t}\int_{\mathbb{R}_{0}^{d}}\left(c(X_{s-}^{i},\mathcal{L}^{1}_{X_{s-}^{i}})-c(X_{s-}^{i,N},\mu_{s-}^{\bm{X}^{N}})\right)z\widetilde{N}^{i}({\rm{d}}s,{\rm{d}}z).$
+$+\int_{0}^{t}\left(\sigma^{0}(X_{s}^{i},\mathcal{L}^{1}_{X_{s}^{i}})-\sigma^{0}(X_{s}^{i,N},\mu_{s}^{\bm{X}^{N}})\right){\rm{d}}W_{s}^{0}+\int_{0}^{t}\int_{\mathbb{R}_{0}^{d}}\left(c(X_{s-}^{i},\mathcal{L}^{1}_{X_{s-}^{i}})-c(X_{s-}^{i,N},\mu_{s-}^{\bm{X}^{N}})\right)z\widetilde{N}^{i}({\rm{d}}s,{\rm{d}}z).$
 Next, for $\lambda\in\mathbb{R}$, by applying Itô’s formula and invoking Assumption (A2), $L_{1}\in\mathbb{R}$, and $L_{2}\geq 0$, we deduce that for all $t\geq 0$,
-$e^{-\lambda t}|X_{t}^{i}-X_{t}^{i,N}|^{2}=\int_{0}^{t}e^{-\lambda s}\Big{[}-\lambda|X_{s}^{i}-X_{s}^{i,N}|^{2}+2\langle X_{s}^{i}-X_{s}^{i,N},b(X_{s}^{i},\mathcal{L}^{1}_{X_{s}^{i}})-b(X_{s}^{i,N},\mu_{s}^{\bm{X}^{N}})\rangle$
+$e^{-\lambda t}|X_{t}^{i}-X_{t}^{i,N}|^{2}=\int_{0}^{t}e^{-\lambda s}\Big[-\lambda|X_{s}^{i}-X_{s}^{i,N}|^{2}+2\langle X_{s}^{i}-X_{s}^{i,N},b(X_{s}^{i},\mathcal{L}^{1}_{X_{s}^{i}})-b(X_{s}^{i,N},\mu_{s}^{\bm{X}^{N}})\rangle$ (3.9)
 $\quad+|\sigma(X_{s}^{i},\mathcal{L}^{1}_{X_{s}^{i}})-\sigma(X_{s}^{i,N},\mu_{s}^{\bm{X}^{N}})|^{2}+|\sigma^{0}(X_{s}^{i},\mathcal{L}^{1}_{X_{s}^{i}})-\sigma^{0}(X_{s}^{i,N},\mu_{s}^{\bm{X}^{N}})|^{2}\Big]{\rm{d}}s$
@@
 $\quad+\langle X_{s}^{i}-X_{s}^{i,N},(\sigma^{0}(X_{s}^{i},\mathcal{L}^{1}_{X_{s}^{i}})-\sigma^{0}(X_{s}^{i,N},\mu_{s}^{\bm{X}^{N}})){\rm{d}}W_{s}^{0}\rangle\bigg]$
-$\quad+\int_{0}^{t}\int_{\mathbb{R}_{0}^{d}}e^{-\lambda s}\left(|X_{s-}^{i}-X_{s-}^{i,N}+(c(X_{s-}^{i},\mathcal{L}^{1}_{X_{s-}^{i}})-c(X_{s-}^{i,N},\mu_{s-}^{\bm{X}^{N}}))z|^{2}-|X_{s-}^{i}-X_{s-}^{i,N}|^{2}\right)\widetilde{N}^{i}({\rm{d}}s,{\rm{d}}z)$
+$\quad+\int_{0}^{t}\int_{\mathbb{R}_{0}^{d}}e^{-\lambda s}\left(|X_{s-}^{i}-X_{s-}^{i,N}+\big(c(X_{s-}^{i},\mathcal{L}^{1}_{X_{s-}^{i}})-c(X_{s-}^{i,N},\mu_{s-}^{\bm{X}^{N}})\big)z|^{2}-|X_{s-}^{i}-X_{s-}^{i,N}|^{2}\right)\widetilde{N}^{i}({\rm{d}}s,{\rm{d}}z)$
 $\quad+\int_{0}^{t}\int_{\mathbb{R}_{0}^{d}}e^{-\lambda s}|(c(X_{s-}^{i},\mathcal{L}^{1}_{X_{s-}^{i}})-c(X_{s-}^{i,N},\mu_{s-}^{\bm{X}^{N}}))z|^{2}\nu({\rm{d}}z){\rm{d}}s$
-$\leq\int_{0}^{t}e^{-\lambda s}\left(-\lambda|X_{s}^{i}-X_{s}^{i,N}|^{2}+2\left\langle X_{s}^{i}-X_{s}^{i,N},b(X_{s}^{i},\mathcal{L}^{1}_{X_{s}^{i}})-b(X_{s}^{i,N},\mu_{s}^{X^{N}})\right\rangle\right){\rm{d}}s$
-$\quad+\int_{0}^{t}e^{-\lambda s}\left(|\sigma(X_{s}^{i},\mathcal{L}^{1}_{X_{s}^{i}})-\sigma(X_{s}^{i,N},\mu_{s}^{X^{N}})|^{2}+|\sigma^{0}(X_{s}^{i},\mathcal{L}^{1}_{X_{s}^{i}})-\sigma^{0}(X_{s}^{i,N},\mu_{s}^{X^{N}})|^{2}\right){\rm{d}}s$
-$\quad+\int_{0}^{t}e^{-\lambda s}|c(X_{s-}^{i},\mathcal{L}^{1}_{X_{s-}^{i}})-c(X_{s-}^{i,N},\mu_{s-}^{X^{N}})|^{2}\int_{\mathbb{R}_{0}^{d}}|z|^{2}\nu({\rm{d}}z){\rm{d}}s$
-$\quad+2\int_{0}^{t}e^{-\lambda s}\bigg{[}\left\langle X_{s}^{i}-X_{s}^{i,N},\left(\sigma(X_{s}^{i},\mathcal{L}^{1}_{X_{s}^{i}})-\sigma(X_{s}^{i,N},\mu_{s}^{X^{N}})\right){\rm{d}}W_{s}^{i}\right\rangle$
-$\quad+\left\langle X_{s}^{i}-X_{s}^{i,N},\left(\sigma^{0}(X_{s}^{i},\mathcal{L}^{1}_{X_{s}^{i}})-\sigma^{0}(X_{s}^{i,N},\mu_{s}^{X^{N}})\right){\rm{d}}W_{s}^{0}\right\rangle\bigg{]}$
-$\quad+\int_{0}^{t}\int_{\mathbb{R}_{0}^{d}}e^{-\lambda s}\left(|X_{s-}^{i}-X_{s-}^{i,N}+\left(c(X_{s-}^{i},\mathcal{L}^{1}_{X_{s-}^{i}})-c(X_{s-}^{i,N},\mu_{s-}^{X^{N}})\right)z|^{2}-|X_{s-}^{i}-X_{s-}^{i,N}|^{2}\right)\widetilde{N}^{i}({\rm{d}}s,{\rm{d}}z)$
-$\leq\int_{0}^{t}e^{-\lambda s}\left(-\lambda|X_{s}^{i}-X_{s}^{i,N}|^{2}+L_{1}|X_{s}^{i}-X_{s}^{i,N}|^{2}+L_{2}\mathcal{W}_{2}^{2}(\mathcal{L}^{1}_{X_{s}^{i}},\mu_{s}^{X^{N}})\right){\rm{d}}s$
-$\quad+2\int_{0}^{t}e^{-\lambda s}\bigg{[}\left\langle X_{s}^{i}-X_{s}^{i,N},\left(\sigma(X_{s}^{i},\mathcal{L}^{1}_{X_{s}^{i}})-\sigma(X_{s}^{i,N},\mu_{s}^{X^{N}})\right){\rm{d}}W_{s}^{i}\right\rangle$
-$\quad+\left\langle X_{s}^{i}-X_{s}^{i,N},\left(\sigma^{0}(X_{s}^{i},\mathcal{L}^{1}_{X_{s}^{i}})-\sigma^{0}(X_{s}^{i,N},\mu_{s}^{X^{N}})\right){\rm{d}}W_{s}^{0}\right\rangle\bigg{]}$
-$\quad+\int_{0}^{t}\int_{\mathbb{R}_{0}^{d}}e^{-\lambda s}\left(|X_{s-}^{i}-X_{s-}^{i,N}+\left(c(X_{s-}^{i},\mathcal{L}^{1}_{X_{s-}^{i}})-c(X_{s-}^{i,N},\mu_{s-}^{X^{N}})\right)z|^{2}-|X_{s-}^{i}-X_{s-}^{i,N}|^{2}\right)\widetilde{N}^{i}({\rm{d}}s,{\rm{d}}z).$
-By taking the expectation and utilizing the bound
-$\mathbb{E}\left[\mathcal{W}_{2}^{2}\big{(}\mathcal{L}^{1}_{X_{s}^{i}},\mathcal{L}^{1}_{X_{s}^{i,N}}\big{)}\right]\leq\mathbb{E}\left[|X_{s}^{i}-X_{s}^{i,N}|^{2}\right],$
-it follows that, for any $\varepsilon>0$,
+$\leq\int_{0}^{t}e^{-\lambda s}\left(-\lambda|X_{s}^{i}-X_{s}^{i,N}|^{2}+2\left\langle X_{s}^{i}-X_{s}^{i,N},b(X_{s}^{i},\mathcal{L}^{1}_{X_{s}^{i}})-b(X_{s}^{i,N},\mu_{s}^{\bm{X}^{N}})\right\rangle\right){\rm{d}}s$
+$\quad+\int_{0}^{t}e^{-\lambda s}\left(|\sigma(X_{s}^{i},\mathcal{L}^{1}_{X_{s}^{i}})-\sigma(X_{s}^{i,N},\mu_{s}^{\bm{X}^{N}})|^{2}+|\sigma^{0}(X_{s}^{i},\mathcal{L}^{1}_{X_{s}^{i}})-\sigma^{0}(X_{s}^{i,N},\mu_{s}^{\bm{X}^{N}})|^{2}\right){\rm{d}}s$
+$\quad+\int_{0}^{t}e^{-\lambda s}|c(X_{s-}^{i},\mathcal{L}^{1}_{X_{s-}^{i}})-c(X_{s-}^{i,N},\mu_{s-}^{\bm{X}^{N}})|^{2}\int_{\mathbb{R}_{0}^{d}}|z|^{2}\nu({\rm{d}}z){\rm{d}}s$
+$\quad+2\int_{0}^{t}e^{-\lambda s}\bigg[\left\langle X_{s}^{i}-X_{s}^{i,N},\left(\sigma(X_{s}^{i},\mathcal{L}^{1}_{X_{s}^{i}})-\sigma(X_{s}^{i,N},\mu_{s}^{\bm{X}^{N}})\right){\rm{d}}W_{s}^{i}\right\rangle$
+$\quad+\left\langle X_{s}^{i}-X_{s}^{i,N},\left(\sigma^{0}(X_{s}^{i},\mathcal{L}^{1}_{X_{s}^{i}})-\sigma^{0}(X_{s}^{i,N},\mu_{s}^{\bm{X}^{N}})\right){\rm{d}}W_{s}^{0}\right\rangle\bigg]$
+$\quad+\int_{0}^{t}\int_{\mathbb{R}_{0}^{d}}e^{-\lambda s}\left(|X_{s-}^{i}-X_{s-}^{i,N}+\left(c(X_{s-}^{i},\mathcal{L}^{1}_{X_{s-}^{i}})-c(X_{s-}^{i,N},\mu_{s-}^{\bm{X}^{N}})\right)z|^{2}-|X_{s-}^{i}-X_{s-}^{i,N}|^{2}\right)\widetilde{N}^{i}({\rm{d}}s,{\rm{d}}z)$
+$\leq\int_{0}^{t}e^{-\lambda s}\left(-\lambda|X_{s}^{i}-X_{s}^{i,N}|^{2}+L_{1}|X_{s}^{i}-X_{s}^{i,N}|^{2}+L_{2}\mathcal{W}_{2}^{2}(\mathcal{L}^{1}_{X_{s}^{i}},\mu_{s}^{\bm{X}^{N}})\right){\rm{d}}s$
+$\quad+2\int_{0}^{t}e^{-\lambda s}\bigg[\left\langle X_{s}^{i}-X_{s}^{i,N},\left(\sigma(X_{s}^{i},\mathcal{L}^{1}_{X_{s}^{i}})-\sigma(X_{s}^{i,N},\mu_{s}^{\bm{X}^{N}})\right){\rm{d}}W_{s}^{i}\right\rangle$
+$\quad+\left\langle X_{s}^{i}-X_{s}^{i,N},\left(\sigma^{0}(X_{s}^{i},\mathcal{L}^{1}_{X_{s}^{i}})-\sigma^{0}(X_{s}^{i,N},\mu_{s}^{\bm{X}^{N}})\right){\rm{d}}W_{s}^{0}\right\rangle\bigg]$
+$\quad+\int_{0}^{t}\int_{\mathbb{R}_{0}^{d}}e^{-\lambda s}\left(|X_{s-}^{i}-X_{s-}^{i,N}+\left(c(X_{s-}^{i},\mathcal{L}^{1}_{X_{s-}^{i}})-c(X_{s-}^{i,N},\mu_{s-}^{\bm{X}^{N}})\right)z|^{2}-|X_{s-}^{i}-X_{s-}^{i,N}|^{2}\right)\widetilde{N}^{i}({\rm{d}}s,{\rm{d}}z).$
+By taking the expectation to (3.9), we have for any $\varepsilon>0$,
 $\begin{split}&e^{-\lambda t}\mathbb{E}\left[|X_{t}^{i}-X_{t}^{i,N}|^{2}\right]\\ &\leq\int_{0}^{t}e^{-\lambda s}\bigg((-\lambda+L_{1})\mathbb{E}\left[|X_{s}^{i}-X_{s}^{i,N}|^{2}\right]+L_{2}\mathbb{E}\left[\mathcal{W}_{2}^{2}(\mathcal{L}^{1}_{X_{s}^{i}},\mu_{s}^{\bm{X}^{N}})\right]\bigg){\rm{d}}s\\ &\leq\int_{0}^{t}e^{-\lambda s}\left((-\lambda+L_{1}+L_{2}+\varepsilon)\mathbb{E}\left[|X_{s}^{i}-X_{s}^{i,N}|^{2}\right]+L_{2}\left(1+\frac{L_{2}}{\varepsilon}\right)\mathbb{E}\left[\mathcal{W}_{2}^{2}(\mathcal{L}^{1}_{X_{s}^{i,N}},\mu_{s}^{\bm{X}^{N}})\right]\right){\rm{d}}s.\end{split}$ (3.10)
-Moreover, by Proposition 3.1, for any $p\in(4,p_{0}]$, we have
+Next, according to (3.2), (3.3), (3.4) and $\mathbb{E}\left[\mathcal{W}_{2}^{2}\big(\mathcal{L}^{1}_{X_{s}^{i}},\mathcal{L}^{1}_{X_{s}^{i,N}}\big)\right]\leq\mathbb{E}\left[|X_{s}^{i}-X_{s}^{i,N}|^{2}\right]$, let $|X_{s}^{j}-X_{s}^{j,N}|^{2}:=\max_{i\in\{1,\ldots,N\}}|X_{s}^{i}-X_{s}^{i,N}|^{2}$ we have
+$\begin{split}\mathbb{E}\left[\mathcal{W}_{2}^{2}(\mathcal{L}^{1}_{X_{s}^{i,N}},\mu_{s}^{\bm{X}^{N}})\right]&\leq\mathbb{E}[\mathcal{W}_{2}(\mathcal{L}^{1}_{X_{s}^{i,N}},\mathcal{L}^{1}_{X_{s}^{i}})+\mathcal{W}_{2}(\mathcal{L}^{1}_{X_{s}^{i}},\mu^{\bm{X}^{N}})+\mathcal{W}_{2}(\mu^{\bm{X}^{N}},\mu_{s}^{\bm{X}^{N}})]^{2}\\ &\leq 3\mathbb{E}[\mathcal{W}^{2}_{2}(\mathcal{L}^{1}_{X_{s}^{i,N}},\mathcal{L}^{1}_{X_{s}^{i}})+\mathcal{W}^{2}_{2}(\mathcal{L}^{1}_{X_{s}^{i}},\mu^{\bm{X}^{N}})+\mathcal{W}^{2}_{2}(\mu^{\bm{X}^{N}},\mu_{s}^{\bm{X}^{N}})]\\ &\leq 3\mathbb{E}[\mathbb{E}\left[|X_{s}^{i}-X_{s}^{i,N}|^{2}\right]+\mathcal{W}^{2}_{2}(\mathcal{L}^{1}_{X_{s}^{i}},\mu^{\bm{X}^{N}})+\frac{1}{N}\sum_{i=1}^{N}|X_{s}^{i}-X_{s}^{i,N}|^{2}]\\ &\leq 3\mathbb{E}[\mathbb{E}\left[|X_{s}^{i}-X_{s}^{i,N}|^{2}\right]+\mathcal{W}^{2}_{2}(\mathcal{L}^{1}_{X_{s}^{i}},\mu^{\bm{X}^{N}})+|X_{s}^{j}-X_{s}^{j,N}|^{2}]\\ &\leq 6\mathbb{E}[|X_{s}^{j}-X_{s}^{j,N}|^{2}]+3\mathbb{E}[\mathcal{W}^{2}_{2}(\mathcal{L}^{1}_{X_{s}^{i}},\mu^{\bm{X}^{N}})].\end{split}$ (3.11)
+Moreover, by Proposition 3.1, for any $p\in[2,p_{0}]$, we have
 $\max_{i\in\{1,\dots,N\}}\sup_{t\in[0,T]}\mathbb{E}\left[\left|X^{i,N}_{t}\right|^{p}\right]\leq C_{T},$
-for some constant $C_{T}>0$.
-Following Fournier and Guillin [21] and Sznitman [38], for any $\mu\in\mathcal{P}_{q}(\mathbb{R}^{d})$ with $q>4$, there exists a constant $C=C(d,q,M_{q}(\mu))$ such that the empirical measure $\mu^{N}$ satisfies
-$\mathbb{E}\left[\mathcal{W}_{2}^{2}(\mu^{N},\mu)\right]\leq C\varphi(N),$
-where $\varphi(N)$ is defined in (3.3).
-Although the original result in Fournier and Guillin [21] assumes independent particles, this bound still holds for our system with common noise $W^{0}$ due to the exchangeability of the particles conditioned on $\mathcal{F}^{0}$.
-The rest of the proof proceeds similarly by controlling the term $\mathbb{E}\left[\mathcal{W}_{2}^{2}\big{(}\mathcal{L}^{1}_{X^{i,N}_{s}},\mu^{X^{N}}_{s}\big{)}\right]$ using the triangle inequality and the exchangeability of the particles.
-If $\mu\in\mathcal{P}_{q}(\mathbb{R}^{d})$ for some $q>4$, that is, if
-$M_{q}(\mu)^{q}:=\int_{\mathbb{R}^{d}}|x|^{q}\,\mu({\rm{d}}x)<\infty,$
-then for each dimension $d\geq 1$, there exists a constant $C=C(d,q,M_{q}(\mu))$ such that, for all $N\geq 2$,
-$\begin{split}\mathbb{E}\left[\mathcal{W}_{2}(\bar{\mu}^{N},\mu)^{2}\right]&\leq C\begin{cases}N^{-1/2}&\text{if }d<4,\\ N^{-1/2}\log N&\text{if }d=4,\\ N^{-2/d}&\text{if }d>4,\end{cases}\\ &=C\varphi(N).\end{split}$
-Next, let $(X_{t}^{i,N})_{i=1}^{N}$ be a collection of random variables that are identically distributed but not independent, and define the empirical measure by
-$\mu_{t}^{\bm{x}^{N}}:=\frac{1}{N}\sum_{i=1}^{N}\delta_{x_{t}^{i,N}}.$
-Meanwhile, let $(X_{t}^{i})_{i=1}^{N}$ be a sequence of independent and identically distributed random variables, and define the corresponding empirical measure by
-$\bar{\mu}_{t}^{N}:=\frac{1}{N}\sum_{i=1}^{N}\delta_{x_{t}^{i}}.$
-In this case, we have
-[…]
+for some constant $C_{T}>0$, together with Carmona and Delarue [10, Theorem 5.8] deduce that
+$\mathbb{E}[\mathcal{W}^{2}_{2}(\mathcal{L}^{1}_{X_{s}^{i}},\mu^{\bm{X}^{N}})]\leq C\varphi(N).$ (3.12)
+Combining (3.11) and (3.12), we have
+$\mathbb{E}\left[\mathcal{W}_{2}^{2}(\mathcal{L}^{1}_{X_{s}^{i,N}},\mu_{s}^{\bm{X}^{N}})\right]$ $\leq 6\mathbb{E}[|X_{s}^{j}-X_{s}^{j,N}|^{2}]+3\mathbb{E}[\mathcal{W}^{2}_{2}(\mathcal{L}^{1}_{X_{s}^{i}},\mu^{\bm{X}^{N}})]$ (3.13)
+$\leq 6\mathbb{E}[|X_{s}^{j}-X_{s}^{j,N}|^{2}]+3C\varphi(N).$
 Consequently, it suffices to choose $\lambda=L_{1}+L_{2}+\epsilon$ in (LABEL:3.6) to conclude that
-$\begin{split}&e^{-\lambda t}\mathbb{E}\left[|X_{t}^{i}-X_{t}^{i,N}|^{2}\right]\\ &\leq\int_{0}^{t}e^{-\lambda s}\left(L_{2}\left(1+\frac{L_{2}}{\varepsilon}\right)\mathbb{E}\left[\mathcal{W}_{2}^{2}(\mathcal{L}^{1}_{X_{s}^{i,N}},\mu_{s}^{\bm{X}^{N}})\right]\right){{\rm{d}}}s\\ &\leq L_{2}\left(1+\frac{L_{2}}{\varepsilon}\right)\int_{0}^{t}e^{-\lambda s}\left(6\mathbb{E}[|X_{s}^{i}-X_{s}^{i,N}|^{2}]+3C\varphi(N)\right){{{\rm{d}}}}s\\ &=6L_{2}\left(1+\frac{L_{2}}{\varepsilon}\right)\int_{0}^{t}e^{-\lambda s}\mathbb{E}[|X_{s}^{i}-X_{s}^{i,N}|^{2}]{{{\rm{d}}}}s+3C\varphi(N)L_{2}\left(1+\frac{L_{2}}{\varepsilon}\right)\int_{0}^{t}e^{-\lambda s}{{\rm{d}}}s.\end{split}$
-Then using Gronwall’s inequality, we have
+$\begin{split}&e^{-\lambda t}\mathbb{E}\left[|X_{t}^{i}-X_{t}^{i,N}|^{2}\right]\\ &\leq\int_{0}^{t}e^{-\lambda s}\left(L_{2}\left(1+\frac{L_{2}}{\varepsilon}\right)\mathbb{E}\left[\mathcal{W}_{2}^{2}(\mathcal{L}^{1}_{X_{s}^{i,N}},\mu_{s}^{\bm{X}^{N}})\right]\right){{\rm{d}}}s\\ &\leq L_{2}\left(1+\frac{L_{2}}{\varepsilon}\right)\int_{0}^{t}e^{-\lambda s}\left(6\mathbb{E}[|X_{s}^{j}-X_{s}^{j,N}|^{2}]+3C\varphi(N)\right){{{\rm{d}}}}s\\ &=6L_{2}\left(1+\frac{L_{2}}{\varepsilon}\right)\int_{0}^{t}e^{-\lambda s}\mathbb{E}[|X_{s}^{j}-X_{s}^{j,N}|^{2}]{{{\rm{d}}}}s+3C\varphi(N)L_{2}\left(1+\frac{L_{2}}{\varepsilon}\right)\int_{0}^{t}e^{-\lambda s}{{\rm{d}}}s.\end{split}$ (3.14)
+Thus, we have
+$e^{-\lambda t}\mathbb{E}\left[|X_{t}^{j}-X_{t}^{j,N}|^{2}\right]\leq 6L_{2}\left(1+\frac{L_{2}}{\varepsilon}\right)\int_{0}^{t}e^{-\lambda s}\mathbb{E}[|X_{s}^{j}-X_{s}^{j,N}|^{2}]{{{\rm{d}}}}s+3C\varphi(N)L_{2}\left(1+\frac{L_{2}}{\varepsilon}\right)\int_{0}^{t}e^{-\lambda s}{{\rm{d}}}s.$
+Then using Grönwall’s inequality, we have
 […]
 When $L_{1}+L_{2}<0$ and $\gamma=\gamma_{1}+\gamma_{2}<0$, one can choose $\epsilon>0$ sufficiently small so that $\lambda<0$. Hence, we have
-$\begin{split}\mathbb{E}\left[|X_{t}^{i}-X_{t}^{i,N}|^{2}\right]&\leq\frac{3C\varphi(N)L_{2}\left(1+\frac{L_{2}}{\varepsilon}\right)}{\lambda}e^{-6L_{2}\left(1+\frac{L_{2}}{\varepsilon}\right)(1-e^{\lambda t})/\lambda}\\ &\leq\frac{3C\varphi(N)L_{2}\left(1+\frac{L_{2}}{\varepsilon}\right)}{\lambda}e^{-6L_{2}\left(1+\frac{L_{2}}{\varepsilon}\right)/\lambda}\\ &=C\varphi(N),\end{split}$
-for any $t\in[0,T]$, where the positive constant $C$ does not depend on time. Finally, when $\gamma<0$, it follows from Proposition 3.1 that for any $p\in(4,p_{0}]$,
-$\max_{i\in\{1,\dots,N\}}\sup_{t\geq 0}\mathbb{E}\left[\left|X^{i,N}_{t}\right|^{q}\right]\leq C,$
-where the positive constant $C$ does not depend on $T$. The desired proof follows. ∎
-4 Tamed-adaptive Euler-Maruyama method
-4.1 Tamed coefficients and adaptive time control
-For each $\Delta\in(0,1)$ define the tamed coefficients
-$\sigma_{\Delta}(x,\mu)$ $:=\frac{\sigma(x,\mu)}{1+\Delta^{1/2}|\sigma(x,\mu)|(1+|x|)},$ (4.1)
-$c_{\Delta}(x,\mu)$ $:=\frac{c(x,\mu)}{1+\Delta^{1/2}|c(x,\mu)|(1+|x|+|b(x,\mu)|)},$ (4.2)
-$\sigma^{0}_{\Delta}(x,\mu)$ $:=\frac{\sigma^{0}(x,\mu)}{1+\Delta^{1/2}|\sigma^{0}(x,\mu)|(1+|x|)}.$ (4.3)
-The adaptive time control function is taken as
-$h(x,\mu)=\frac{h_{0}}{(1+|b(x,\mu)|+|\sigma(x,\mu)|+|x|^{l})^{2}+|c(x,\mu)|^{p_{0}}},$ (4.4)
-with $h_{0}>0$ (we set $h_{0}=1$ in proofs), and $l,p_{0}$ as in (A3) and (A4). We will use the standard taming bounds
-$|\sigma_{\Delta}(x,\mu)|(1+|x|)\leq\Delta^{-1/2},$ (4.5)
-$|c_{\Delta}(x,\mu)|(1+|x|)\leq\Delta^{-1/2},$ (4.6)
-$|\sigma(x,\mu)-\sigma_{\Delta}(x,\mu)|\leq C\sqrt{\Delta}\,|\sigma(x,\mu)|^{2}(1+|x|),$ (4.7)
-$|\sigma^{0}(x,\mu)-\sigma^{0}_{\Delta}(x,\mu)|\leq C\sqrt{\Delta}\,|\sigma^{0}(x,\mu)|^{2}(1+|x|).$ (4.8)
-4.2 Continuous-time tamed adaptive scheme with common noise
-Fix $N\in\mathbb{N}$.
-For $i=1,\dots,N$, let $W^{i}$ be independent Brownian motions, $W^{0}$ a common Brownian motion, and $Z^{i}$ independent Lévy processes with common Lévy measure $\nu$.
-Define the empirical measure $\mu_{\widehat{X}^{N}_{t}}:=\frac{1}{N}\sum_{j=1}^{N}\delta_{\widehat{X}^{j,N}_{t}}$.
-The continuous-time tamed adaptive Euler-Maruyama approximation $\widehat{X}^{i,N}_{t}$ is the Itô process
-$\widehat{X}^{i,N}_{t}$ $=x_{0}+\int_{0}^{t}b\bigl{(}\widehat{X}^{i,N}_{s},\mu_{\widehat{X}^{N}_{s}}\bigr{)}\,{\rm{d}}s+\int_{0}^{t}\sigma_{\Delta}\!\bigl{(}\widehat{X}^{i,N}_{s},\mu_{\widehat{X}^{N}_{s}}\bigr{)}\,{\rm{d}}W^{i}_{s}$
-$\quad+\int_{0}^{t}\sigma^{0}_{\Delta}\!\bigl{(}\widehat{X}^{i,N}_{s},\mu_{\widehat{X}^{N}_{s}}\bigr{)}\,{\rm{d}}W^{0}_{s}+\int_{0}^{t}\int_{\mathbb{R}^{d}_{0}}c_{\Delta}\!\bigl{(}\widehat{X}^{i,N}_{s-},\mu_{\widehat{X}^{N}_{s-}}\bigr{)}z\,\widetilde{N}^{i}({\rm{d}}s,{\rm{d}}z),$ (4.9)
-where $\widetilde{N}^{i}$ is the compensated Poisson random measure of $Z^{i}$.
-Remark 4.1.
-[…]
-$\max_{1\leq i\leq N}\mathbb{E}\Big{[}\sup_{0\leq t\leq T}|\widehat{X}^{i,N}_{t}|^{p}\Big{]}\leq C.$
+$\begin{split}\mathbb{E}\left[|X_{t}^{j}-X_{t}^{j,N}|^{2}\right]&\leq\frac{3C\varphi(N)L_{2}\left(1+\frac{L_{2}}{\varepsilon}\right)}{\lambda}e^{-6L_{2}\left(1+\frac{L_{2}}{\varepsilon}\right)(1-e^{\lambda t})/\lambda}\\ &\leq\frac{3C\varphi(N)L_{2}\left(1+\frac{L_{2}}{\varepsilon}\right)}{\lambda}e^{-6L_{2}\left(1+\frac{L_{2}}{\varepsilon}\right)/\lambda}\\ &=C\varphi(N),\end{split}$
+for any $t\in[0,T]$, where the positive constant $C$ does not depend on time. This completes the proof. ∎
+4 Tamed-Adaptive Euler-Maruyama (TAEM) Method
+4.1 Discrete-time and Continuous-time TAEM Method
+Set
+$t_{0}:=0,\qquad X_{t_{0}}:=X_{0},$
+where $X_{0}$ is the given initial random variable in (1.1). We now consider the tamed-adaptive EM scheme with step
+$\Delta_{t_{k}}:=\frac{h}{1+\beta\,|b(X_{t_{k}},\mathcal{L}^{1}(X_{t_{k}}))|},\qquad\beta>0,\ h\in(0,1],\qquad i.e.\qquad t_{k+1}=t_{k}+\Delta_{t_{k}},$ (4.1)
+and tamed drift
+$b_{\Delta_{t_{k}}}(x,\mu):=\frac{b(x,\mu)}{1+\Delta_{t_{k}}^{\alpha}|b(x,\mu)|},\qquad 0<\alpha<\frac{1}{2}.$ (4.2)
+The TAEM discrete approximation is then defined by
+$X_{t_{k+1}}$ $=X_{t_{k}}+b_{\Delta_{t_{k}}}\bigl(X_{t_{k}},\mathcal{L}^{1}(X_{t_{k}})\bigr)\,\Delta_{t_{k}}+\sigma\bigl(X_{t_{k}},\mathcal{L}^{1}(X_{t_{k}})\bigr)\,\Delta W_{t_{k}}$ (4.3)
+$\quad+c\bigl(X_{t_{k}},\mathcal{L}^{1}(X_{t_{k}})\bigr)\,\Delta Z_{t_{k}}+\sigma^{0}\bigl(X_{t_{k}},\mathcal{L}^{1}(X_{t_{k}})\bigr)\,\Delta W^{0}_{t_{k}},$
+where $\Delta W_{t_{k}}=W(t_{k+1})-W(t_{k})$, $\Delta Z_{t_{k}}=Z(t_{k+1})-Z(t_{k})$, and $\Delta W^{0}_{t_{k}}=W^{0}(t_{k+1})-W^{0}(t_{k})$.
+The iteration stops when $t_{k}\geq T$.
+Note that at each time $t_{k}$, the random variables $\Delta_{t_{k}}$ and all coefficients are $\mathcal{F}_{t_{k}}$-measurable.
+The piecewise constant interpolant is defined by
+$\overline{X}^{\Delta}(t):=X_{t_{k}},\qquad t\in[t_{k},t_{k+1}),$ (4.4)
+and the continuous-time TAEM approximation is
+$X^{\Delta}(t)$ $=X_{0}+\int_{0}^{t}b_{\Delta_{t_{s}}}\bigl(\overline{X}^{\Delta}(s),\mathcal{L}^{1}(\overline{X}^{\Delta}(s))\bigr)\,\mathrm{d}s+\int_{0}^{t}\sigma\bigl(\overline{X}^{\Delta}(s),\mathcal{L}^{1}(\overline{X}^{\Delta}(s))\bigr)\,\mathrm{d}W(s)$ (4.5)
+$\quad+\int_{0}^{t}c\bigl(\overline{X}^{\Delta}(s),\mathcal{L}^{1}(\overline{X}^{\Delta}(s))\bigr)\,\mathrm{d}Z(s)+\int_{0}^{t}\sigma^{0}\bigl(\overline{X}^{\Delta}(s),\mathcal{L}^{1}(\overline{X}^{\Delta}(s))\bigr)\,\mathrm{d}W^{0}(s),$
+where
+$\Delta_{t_{s}}:=\Delta_{t_{k}}=\frac{h}{1+\beta|b(X_{t_{k}},\mathcal{L}^{1}(X_{t_{k}}))|}\leq h,\quad{s}\in[t_{k},t_{k+1}).$
+4.2 Reachability of $\Delta_{t_{k}}$ over time $T$
+Theorem 4.1.
+Suppose that assumptions (A1)-(A7) hold.
+Let $\Delta_{t_{k}}$ be given by (4.1) and define the stopping time $\tau_{T}:=\inf\left\{n\geq 0:S_{n}:=\sum_{k=0}^{n-1}\Delta_{t_{k}}\geq T\right\}.$ Then, $\mathbb{E}[\tau_{T}]<\infty$.
+In particular, $\tau_{T}$ is almost surely finite.
+In order to prove the aforementioned theorem, the following lemmas are presented.
+Lemma 4.2.
+Suppose that assumptions (A1)-(A7) hold.
+Let $X^{\Delta}(t)$ be the continuous-time TAEM approximation given in (4.5).
+Then, for any real number $l\geq 2$, there exists a constant $C_{T,l}>0$, depending only on $T$, $l$ such that
+$\sup_{t\leq T}\mathbb{E}[|X^{\Delta}(t)|^{l}]\leq C_{T,l}.$
+In particular, since $X^{\Delta}(t_{k})=X_{t_{k}}$ for all grid points $t_{k}$, we also have
+$\sup_{t_{k}\leq T}\mathbb{E}\big[|X_{t_{k}}|^{l}\big]\leq C_{T,l}.$
 Proof.
-The proof of this remark can be established by adapting the methodology presented in [41]. ∎
+We divide the proof into several steps.
+Step 1: Itô’s formula for $|X^{\Delta}(t)|^{l}$.
+Let $\mu_{s}=\mathcal{L}^{1}(\overline{X}^{\Delta}(s))$ and define
+$f(s)$ $:=b_{\Delta_{t_{s}}}(\overline{X}^{\Delta}(s),\mu_{s}),\qquad g(s):=\sigma(\overline{X}^{\Delta}(s),\mu_{s}),$
+$g_{0}(s)$ $:=\sigma^{0}(\overline{X}^{\Delta}(s),\mu_{s}),\qquad j(s):=c(\overline{X}^{\Delta}(s),\mu_{s}).$
+Applying Itô’s formula to $|X^{\Delta}(t)|^{l}$ for $l\geq 2$
+$|X^{\Delta}(t)|^{l}=|X_{0}|^{l}+l\int_{0}^{t}|X^{\Delta}(s)|^{l-2}\,\langle X^{\Delta}(s),f(s)\rangle\,\mathrm{d}s$ (4.6)
+$\quad+l\int_{0}^{t}|X^{\Delta}(s)|^{l-2}\,\langle X^{\Delta}(s),g(s)\,\mathrm{d}W(s)\rangle+l\int_{0}^{t}|X^{\Delta}(s)|^{l-2}\,\langle X^{\Delta}(s),g_{0}(s)\,\mathrm{d}W^{0}(s)\rangle$
+$\quad+\frac{l(l-2)}{2}\int_{0}^{t}|X^{\Delta}(s)|^{l-4}\Big(|g(s)^{\top}X^{\Delta}(s)|^{2}+|g_{0}(s)^{\top}X^{\Delta}(s)|^{2}\Big)\,\mathrm{d}s$
+$\quad+\frac{l}{2}\int_{0}^{t}|X^{\Delta}(s)|^{l-2}\Big(\|g(s)\|^{2}+\|g_{0}(s)\|^{2}\Big)\,\mathrm{d}s$
+$\quad+2\int_{0}^{t}\int_{\mathbb{R}_{0}^{d}}\Big[|X^{\Delta}(s-)+j(s)z|^{l}-|X^{\Delta}(s-)|^{l}\Big]\,\tilde{N}(\mathrm{d}s,\mathrm{d}z)$
+$\quad+\int_{0}^{t}\int_{\mathbb{R}_{0}^{d}}\Big[2|X^{\Delta}(s-)+j(s)z|^{l}-2|X^{\Delta}(s-)|^{l}-l|X^{\Delta}(s-)|^{l-2}\,\langle X^{\Delta}(s-),j(s)z\rangle\Big]\,\nu(\mathrm{d}z)\,\mathrm{d}s.$
+Step 2: Drift term estimation.
+For the drift term $f(s):=b_{\Delta_{t_{s}}}(\overline{X}^{\Delta}(s),\mu_{s})$, by the taming property, we obtain
+$|f(s)|\leq|b(\overline{X}^{\Delta}(s),\mu_{s})|.$
+Using (A3) with $\overline{x}=0$ and $\overline{\mu}=\delta_{0}$, there exist constants $\tilde{L}>0$ and $\ell>0$ such that
+$|b(\overline{X}^{\Delta}(s),\mu_{s})|\leq|b(0,\delta_{0})|+\tilde{L}(1+|\overline{X}^{\Delta}(s)|^{\ell})\left(|\overline{X}^{\Delta}(s)|+\mathcal{W}_{2}(\mu_{s},\delta_{0})\right).$
+Since $\mathcal{W}_{2}(\mu_{s},\delta_{0})=\left(\mathbb{E}[|\overline{X}^{\Delta}(s)|^{2}\mid\mathcal{F}^{0}]\right)^{1/2}$, we have
+$|b(\overline{X}^{\Delta}(s),\mu_{s})|$ $\leq C+\tilde{L}(1+|\overline{X}^{\Delta}(s)|^{\ell})|\overline{X}^{\Delta}(s)|+\tilde{L}(1+|\overline{X}^{\Delta}(s)|^{\ell})\left(\mathbb{E}[|\overline{X}^{\Delta}(s)|^{2}\mid\mathcal{F}^{0}]\right)^{1/2}$
+$\leq C+\tilde{L}(1+|\overline{X}^{\Delta}(s)|^{\ell})|\overline{X}^{\Delta}(s)|+\frac{\tilde{L}}{2}(1+|\overline{X}^{\Delta}(s)|^{\ell})^{2}+\frac{\tilde{L}}{2}\mathbb{E}[|\overline{X}^{\Delta}(s)|^{2}\mid\mathcal{F}^{0}]$
+$\leq C\left(1+|\overline{X}^{\Delta}(s)|^{\ell+1}+|\overline{X}^{\Delta}(s)|^{2\ell}+\mathbb{E}[|\overline{X}^{\Delta}(s)|^{2}\mid\mathcal{F}^{0}]\right).$
+Now estimate the drift term, by choosing $l\geq\max\{\,2\ell,\ell+1,2\},$ we have
+$\left|l\int_{0}^{t}|X^{\Delta}(s)|^{l-2}\langle X^{\Delta}(s),f(s)\rangle\,\mathrm{d}s\right|\leq l\int_{0}^{t}|X^{\Delta}(s)|^{l-1}|f(s)|\,\mathrm{d}s$
+$\leq C_{l}\int_{0}^{t}|X^{\Delta}(s)|^{l-1}\left(1+|\overline{X}^{\Delta}(s)|^{\ell+1}+|\overline{X}^{\Delta}(s)|^{2\ell}+\mathbb{E}[|\overline{X}^{\Delta}(s)|^{2}\mid\mathcal{F}^{0}]\right)\,\mathrm{d}s$
+$\leq C_{l}\int_{0}^{t}\left(1+|X^{\Delta}(s)|^{l}+|\overline{X}^{\Delta}(s)|^{l}+\mathbb{E}[|\overline{X}^{\Delta}(s)|^{l}\mid\mathcal{F}^{0}]\right)\,\mathrm{d}s.$
+Step 3: Diffusion terms estimation.
+From Assumption (A1), we have
+$|g(s)|^{2}+|g_{0}(s)|^{2}\leq L\left(1+|\overline{X}^{\Delta}(s)|^{2}+\mathcal{W}_{2}^{2}(\mu_{s},\delta_{0})\right)\leq L\left(1+|\overline{X}^{\Delta}(s)|^{2}+\mathbb{E}[|\overline{X}^{\Delta}(s)|^{2}\mid\mathcal{F}^{0}]\right).$
+Then we obtain
+$\frac{l(l-2)}{2}\int_{0}^{t}|X^{\Delta}(s)|^{l-4}\Big(|g(s)^{\top}X^{\Delta}(s)|^{2}+|g_{0}(s)^{\top}X^{\Delta}(s)|^{2}\Big)\,\mathrm{d}s+\frac{l}{2}\int_{0}^{t}|X^{\Delta}(s)|^{l-2}\Big(\|g(s)\|^{2}+\|g_{0}(s)\|^{2}\Big)\,\mathrm{d}s$
+$\leq C_{l}\int_{0}^{t}|X^{\Delta}(s)|^{l-2}\left(|g(s)|^{2}+|g_{0}(s)|^{2}\right)\,\mathrm{d}s$
+$\leq C_{l,L}\int_{0}^{t}\left[1+|X^{\Delta}(s)|^{l}+|\overline{X}^{\Delta}(s)|^{l}+\mathbb{E}[|\overline{X}^{\Delta}(s)|^{l}\mid\mathcal{F}^{0}]\right]\,\mathrm{d}s.$
+Step 4: Jump term estimation.
+Write the jump contribution as $M_{J}(t)+D_{J}(t)$, where the martingale part is
+$M_{J}(t)=2\int_{0}^{t}\!\!\int_{\mathbb{R}_{0}^{d}}\Big[|X^{\Delta}(s{-})+j(s)z|^{l}-|X^{\Delta}(s{-})|^{l}\Big]\,\tilde{N}(\mathrm{d}z,\mathrm{d}s),$
+and the compensator is
+$D_{J}(t)=\int_{0}^{t}\int_{\mathbb{R}_{0}^{d}}\Big[2|X^{\Delta}(s-)+j(s)z|^{l}-2|X^{\Delta}(s-)|^{l}-l|X^{\Delta}(s-)|^{l-2}\,\langle X^{\Delta}(s-),j(s)z\rangle\Big]\,\nu(\mathrm{d}z)\,\mathrm{d}s.$
+For $l\geq 2$, by Taylor’s formula there exists $C_{l}>0$ such that for all $x,z$,
+$\big||x+jz|^{l}-|x|^{l}-l|x|^{l-2}\langle x,jz\rangle\big|\;\leq\;C_{l}\big(|x|^{l-2}|jz|^{2}+|jz|^{l}\big),$
+and also
+$\big||x+jz|^{l}-|x|^{l}\big|\;\leq\;C_{l}\big(|x|^{l-1}|jz|^{2}+|jz|^{l}\big).$
+Hence,
+$|D_{J}(t)|\leq C_{l}\int_{0}^{t}\!\!\int_{\mathbb{R}_{0}^{d}}\Big(|X^{\Delta}(s)|^{l-2}|j(s)|^{2}|z|^{2}+2|j(s)|^{l}|z|^{l}+|X^{\Delta}(s)|^{l-1}|j(s)|^{2}|z|^{2}\Big)\,\nu(\mathrm{d}z)\,\mathrm{d}s.$
+By Assumption (A5), we have
+$|j(s)|\;\leq\;L_{3}\!\left(1+|\overline{X}^{\Delta}(s)|+\big[\mathbb{E}\!\left(|\overline{X}^{\Delta}(s)|^{2}\mid\mathcal{F}^{0}\right)\big]^{1/2}\right),$
+so that $|j(s)|^{2}$ and $|j(s)|^{l}$ are bounded by $1+|\overline{X}^{\Delta}(s)|^{l}+\mathbb{E}\!\left(|\overline{X}^{\Delta}(s)|^{l}\mid\mathcal{F}^{0}\right).$ Moreover, Assumption (A4) ensures $\int_{|z|<1}|z|^{2}\nu(\mathrm{d}z)<\infty$ and $\int_{|z|\geq 1}|z|^{l}\nu(\mathrm{d}z)<\infty.$ Collecting the above, we obtain
+$|D_{J}(t)|\;\leq\;C_{l,L_{3},\nu}\!\int_{0}^{t}\Big[1+|X^{\Delta}(s)|^{l}+|\overline{X}^{\Delta}(s)|^{l}+\mathbb{E}\!\left(|\overline{X}^{\Delta}(s)|^{l}\mid\mathcal{F}^{0}\right)\Big]\,\mathrm{d}s.$
+Step 5: Taking expectations and Grönwall’s inequality.
+Taking expectations to (4.6), all martingale terms vanish. Let
+$m(t)=\sup_{0\leq u\leq t}\mathbb{E}[|X^{\Delta}(u)|^{l}],$
+we obtain
+$\mathbb{E}[|X^{\Delta}(t)|^{l}]\leq\mathbb{E}[|X_{0}|^{l}]+C\int_{0}^{t}\left[1+\mathbb{E}[|X^{\Delta}(s)|^{l}]+\mathbb{E}[|\overline{X}^{\Delta}(s)|^{l}]\right]\,\mathrm{d}s.$
+Since $\overline{X}^{\Delta}(s)=X_{t_{k}}$ for $s\in[t_{k},t_{k+1})$, we have $\mathbb{E}[|\overline{X}^{\Delta}(s)|^{l}]\leq m(s)$ for $s\leq t$. Therefore,
+$m(t)\leq\mathbb{E}[|X_{0}|^{l}]+C\int_{0}^{t}\left[1+2m(s)\right]\,\mathrm{d}s.$
+By Grönwall’s inequality, we get
+$m(T)\leq\left(\mathbb{E}[|X_{0}|^{l}]+CT\right)e^{2CT}=:C_{T,l}.$
+This completes the proof of the lemma. ∎
+Proof of Theorem 4.1.
+According to Assumption (A3), we have
+$|b(x,\mu)-b(0,\delta_{0})|\leq\tilde{L}(1+|x|^{\ell}+|0|^{\ell})\,(|x-0|+\mathcal{W}_{2}(\mu,\delta_{0}))=\tilde{L}(1+|x|^{\ell})(|x|+\mathcal{W}_{2}(\mu,\delta_{0})).$
+Taking $x=X_{t_{k}},\ \mu=\mathcal{L}^{1}(X_{t_{k}})$, we have
+$|b(X_{t_{k}},\mathcal{L}^{1}(X_{t_{k}})|$ $\leq C_{0}+L(1+|X_{t_{k}}|^{\ell})(|X_{t_{k}}|+\mathcal{W}_{2}(\mathcal{L}^{1}(X_{t_{k}},\delta_{0}))$
+$\leq C_{0}+L(1+|X_{t_{k}}|^{\ell})\Big(|X_{t_{k}}|+(\mathbb{E}[|X_{t_{k}}|^{2}\mid\mathcal{F}^{0}])^{1/2}\Big),$
+where $C_{0}=|b(0,\delta_{0})|.$ By Hölder’s inequality and Lemma 4.2, both $\mathbb{E}[(1+|X_{t_{k}}|^{\ell})|X_{t_{k}}|]$ and $\mathbb{E}[1+|X_{t_{k}}|^{\ell}\cdot(\mathbb{E}|X_{t_{k}}|^{2})^{1/2}]$ are bounded by a constant depending only on $L$ and $C_{T,\ell}$.
+Hence there exists $M=M(L,C_{T,\ell})$ such that,
+$\mathbb{E}[|b(X_{t_{k}},\mathcal{L}^{1}(X_{t_{k}}))|]\leq C_{0}+L\mathbb{E}\big[(1+|X_{t_{k}}|^{\ell})|X_{t_{k}}|\big]+L\mathbb{E}[1+|X_{t_{k}}|^{\ell}]\cdot(\mathbb{E}[|X_{t_{k}}|^{2}])^{1/2}\leq M.$
+Define $f(y)=\dfrac{h}{1+\beta y}$ for $y\geq 0$. Note $f$ is convex and nonnegative on $0,\infty)$. By Jensen’s inequality,
+$\mathbb{E}[\Delta_{t_{k}}]=\mathbb{E}\big[f(|b(X_{t_{k}},\mathcal{L}^{1}(X_{t_{k}}))|)\big]\geq f\big(\mathbb{E}[|b(X_{t_{k}},\mathcal{L}^{1}(X_{t_{k}}))|]\big)\geq\frac{h}{1+\beta M}=:\,c>0.$
+Due to $S_{\tau_{T}}-T\leq\Delta_{t_{\tau_{T}-1}}\leq h$. Fix a constant truncation integer $N$, we have
+$S_{\tau_{T}\wedge N}\leq S_{\tau_{T}}\leq T+h.$ (4.7)
+Taking expectations to (4.7), we have
+$\mathbb{E}[S_{\tau_{T}\wedge N}]$ $=\mathbb{E}[\sum_{k=0}^{\tau_{T}\wedge N-1}\Delta_{t_{k}}]=\mathbb{E}[\sum_{k=0}^{N-1}\Delta_{t_{k}}\mathbf{1}_{\{\tau_{T}\wedge N>k\}}]=\sum_{k=0}^{N-1}\mathbb{E}[\Delta_{t_{k}}\mathbf{1}_{\{\tau_{T}>k\}}]$
+$\geq\sum_{k=0}^{N-1}\mathbb{E}[\Delta_{t_{k}}]\,\mathbb{E}[\mathbf{1}_{\{\tau_{T}>k\}}]\geq c\,\mathbb{E}[\sum_{k=0}^{N-1}\mathbf{1}_{\{\tau_{T}>k\}}]=c\,\mathbb{E}[\tau_{T}\wedge N].$
+Combining with $\mathbb{E}[S_{\tau_{T}\wedge N}]\leq T+h$ yields
+$c\,\mathbb{E}[\tau_{T}\wedge N]\leq T+h.$
+Let $N\to\infty$ and apply monotone convergence to get $c\,\mathbb{E}[\tau_{T}]\leq T+h$, so we have $\mathbb{E}[\tau_{T}]<\infty$.
+Therefore $\mathbb{P}(\tau_{T}=\infty)=0$ and $\tau_{T}<\infty$ almost surely.
+∎
+4.3 Convergence of TAEM Method under assumptions (A1)-(A7)
+We now establish strong convergence of the TAEM scheme (4.3)-(4.5) under assumptions (A1)-(A7). Fix $R>0$ and define the stopping time
+$\tau_{R}:=\inf\bigl\{t\geq 0:|X(t)|\geq R\ \text{or}\ |X^{\Delta}(t)|\geq R\bigr\}.$ (4.8)
+Consider the stopped processes
+$X^{R}(t):=X(t\wedge\tau_{R}),\qquad X^{\Delta,R}(t):=X^{\Delta}(t\wedge\tau_{R}),$
+and set
+$e^{R}(t):=X^{\Delta,R}(t)-X^{R}(t),\qquad e(t):=X^{\Delta}(t)-X(t).$
 Theorem 4.3.
-Assume (A1)-(A7) and use the definitions (4.1)-(4.4).
-Let $X^{i,N}$ be the interacting particle system (4.2) and $\widehat{X}^{i,N}$ the continuous tamed-adaptive approximation (4.2).
-Then for any $T>0$ there exists $C_{T}>0$ such that
-$\max_{1\leq i\leq N}\sup_{0\leq t\leq T}\mathbb{E}\bigl{[}|X^{i,N}_{t}-\widehat{X}^{i,N}_{t}|^{2}\bigr{]}\leq C_{T}\Delta.$
-If moreover $L_{1}+L_{2}<0$ and $\gamma<0$, the constant can be chosen independent of $T$, i.e.
-$\max_{1\leq i\leq N}\sup_{t\geq 0}\mathbb{E}\bigl{[}|X^{i,N}_{t}-\widehat{X}^{i,N}_{t}|^{2}\bigr{]}\leq C\Delta.$
+Suppose that assumptions (A1)-(A7) hold. Let $X(t)$ be the unique strong solution of (1.1), and let $X^{\Delta}(t)$ be the continuous-time TAEM approximation given in (4.5), with adaptive time step
+$\Delta_{t_{k}}=\frac{h}{1+\beta\,|b(X_{t_{k}},\mathcal{L}^{1}(X_{t_{k}}))|}.$
+Then for any $\varepsilon\in(0,1)$ and any $p>2$, there exist constants $C_{T,\varepsilon,p}>0$ and $C_{T,p}>0$, independent of $h$, and a constant $C>0$ depending only on $\ell>0$, such that
+$\mathbb{E}\!\left[\sup_{0\leq t\leq T}|X^{\Delta}(t)-X(t)|^{2}\right]\;\leq\;C_{T,\varepsilon,p}\,h^{1-\varepsilon}\,(\log(1/h))^{C}\;+\;C_{T,p}\,(\log(1/h))^{-\frac{p-2}{2\ell}}.$
+In particular, the TAEM method is strongly convergent in the sense that
+$\lim_{h\to 0}\mathbb{E}\!\left[\sup_{0\leq t\leq T}|X^{\Delta}(t)-X(t)|^{2}\right]=0.$
+In order to prove the aforementioned theorem, the following lemmas are presented.
+Lemma 4.4.
+Let $\overline{X}^{\Delta}(t)$ be the piecewise constant interpolant defined in (4.4), $X^{\Delta}(t)$ be the continuous-time TAEM approximation defined in (4.5), and $\tau_{R}$ be the stopping time defined in (4.8), there exists a constant $C_{R}>0$ such that
+$\int_{0}^{T}\mathbb{E}|X^{\Delta}(s)-\overline{X}^{\Delta}(s)|^{2}\,{\rm{d}}s\ \leq\ T\,C_{R}\,h.$
 Proof.
-We denote by $X^{i,N}$ the exact interacting particle system with common noise
-$X^{i,N}_{t}$ $=x_{0}+\int_{0}^{t}b\bigl{(}X^{i,N}_{s},\mu_{X^{N}_{s}}\bigr{)}\,{\rm{d}}s+\int_{0}^{t}\sigma\bigl{(}X^{i,N}_{s},\mu_{X^{N}_{s}}\bigr{)}\,{\rm{d}}W^{i}_{s}$
-$\quad+\int_{0}^{t}\sigma^{0}\bigl{(}X^{i,N}_{s},\mu_{X^{N}_{s}}\bigr{)}\,{\rm{d}}W^{0}_{s}+\int_{0}^{t}\int_{\mathbb{R}^{d}_{0}}c\bigl{(}X^{i,N}_{s-},\mu_{X^{N}_{s-}}\bigr{)}z\,\widetilde{N}^{i}({\rm{d}}s,{\rm{d}}z).$ (4.10)
-Set the error $e^{i,N}_{t}:=X^{i,N}_{t}-\widehat{X}^{i,N}_{t}$. Applying Itô’s formula to $|e^{i,N}_{t}|^{2}$ yields
-$|e^{i,N}_{t}|^{2}$ $=2\int_{0}^{t}\big{\langle}e^{i,N}_{s},\,b(X^{i,N}_{s},\mu_{X^{N}_{s}})-b(\widehat{X}^{i,N}_{s},\mu_{\widehat{X}^{N}_{s}})\big{\rangle}{\rm{d}}s$
-$\quad+\int_{0}^{t}\big{\|}\sigma(X^{i,N}_{s},\mu_{X^{N}_{s}})-\sigma_{\Delta}(\widehat{X}^{i,N}_{s},\mu_{\widehat{X}^{N}_{s}})\big{\|}^{2}{\rm{d}}s$
-$\quad+\int_{0}^{t}\big{\|}\sigma^{0}(X^{i,N}_{s},\mu_{X^{N}_{s}})-\sigma^{0}_{\Delta}(\widehat{X}^{i,N}_{s},\mu_{\widehat{X}^{N}_{s}})\big{\|}^{2}{\rm{d}}s$
-$\quad+\int_{0}^{t}\int_{\mathbb{R}^{d}_{0}}\big{|}c(X^{i,N}_{s-},\mu_{X^{N}_{s-}})-c_{\Delta}(\widehat{X}^{i,N}_{s-},\mu_{\widehat{X}^{N}_{s-}})\big{|}^{2}|z|^{2}\nu(dz)\,{\rm{d}}s$
-$\quad+\mathcal{M}^{i,N}_{t},$ (4.11)
-where $\mathcal{M}^{i,N}_{t}$ is a local martingale collecting the stochastic integral remainders.
-Decompose each coefficient difference into a consistency part and a taming part, e.g.
-$).\sigma(X,\mu)-\sigma_{\Delta}(\widehat{X},\widehat{\mu})=\big{(}\sigma(X,\mu)-\sigma(\widehat{X},\widehat{\mu})\big{)}+\big{(}\sigma(\widehat{X},\widehat{\mu})-\sigma_{\Delta}(\widehat{X},\widehat{\mu}$$\big{)}$.
-Using Assumptions (A2)-(A3) and the taming control (4.8), for any small $\varepsilon>0$ we can bound
-$\big{\|}\sigma(X,\mu)-\sigma_{\Delta}(\widehat{X},\widehat{\mu})\big{\|}^{2}$ $\leq(1+\varepsilon)\big{\|}\sigma(X,\mu)-\sigma(\widehat{X},\widehat{\mu})\big{\|}^{2}+C(\varepsilon)\Delta(1+|\widehat{X}|^{q_{1}}),$ (4.12)
-$\big{\|}\sigma^{0}(X,\mu)-\sigma^{0}_{\Delta}(\widehat{X},\widehat{\mu})\big{\|}^{2}$ $\leq(1+\varepsilon)\big{\|}\sigma^{0}(X,\mu)-\sigma^{0}(\widehat{X},\widehat{\mu})\big{\|}^{2}+C(\varepsilon)\Delta(1+|\widehat{X}|^{q_{2}}),$ (4.13)
-with analogous estimates for $c$; here $q_{1},q_{2}$ reflect coefficient growth and are handled by moment bounds.
-Plugging these decompositions into (4.2), using (A2), Young’s and BDG’s inequalities, yields for suitable constants $C_{1},C_{2}$ and a polynomial moment term $R_{s}$, which is bounded in expectation on $[0,T]$
-$|e^{i,N}_{t}|^{2}\leq\int_{0}^{t}\big{(}C_{1}|e^{i,N}_{s}|^{2}+C_{2}\mathcal{W}_{2}^{2}(\mu_{X^{N}_{s}},\mu_{\widehat{X}^{N}_{s}})\big{)}\,{\rm{d}}s+C\Delta\int_{0}^{t}(1+R_{s})\,{\rm{d}}s+\mathcal{M}^{i,N}_{t}.$ (4.14)
-Take expectations in (4.14) and use the standard bound
-$\mathcal{W}_{2}^{2}(\mu_{X^{N}_{s}},\mu_{\widehat{X}^{N}_{s}})\leq\frac{1}{N}\sum_{j=1}^{N}|e^{j,N}_{s}|^{2}.$
-Let $\mathcal{E}_{t}:=\max_{1\leq i\leq N}\mathbb{E}[|e^{i,N}_{t}|^{2}]$. Then from (4.14) we obtain
-$\mathcal{E}_{t}\leq C\int_{0}^{t}\mathcal{E}_{s}\,{\rm{d}}s+C\Delta\int_{0}^{t}(1+\mathbb{E}[R_{s}])\,{\rm{d}}s.$
-By uniform moment bounds, $\mathbb{E}[R_{s}]\leq C_{T}$ on $[0,T]$. Hence
-$\mathcal{E}_{t}\leq C\int_{0}^{t}\mathcal{E}_{s}\,{\rm{d}}s+C_{T}\Delta t,$
-and Gronwall’s inequality gives
-$\mathcal{E}_{t}\leq C_{T}\Delta,\qquad t\in[0,T].$
-Therefore, we have
-$\max_{1\leq i\leq N}\sup_{0\leq t\leq T}\mathbb{E}\big{[}|X^{i,N}_{t}-\widehat{X}^{i,N}_{t}|^{2}\big{]}\leq C_{T}\Delta.$
-If in addition the contractivity conditions $L_{1}+L_{2}<0$ and $\gamma=\gamma_{1}+\gamma_{2}<0$ hold, then the same reasoning yields a time-uniform estimate
-$\max_{1\leq i\leq N}\sup_{t\geq 0}\mathbb{E}\big{[}|X^{i,N}_{t}-\widehat{X}^{i,N}_{t}|^{2}\big{]}\leq C\Delta,$
-with $C$ independent of $T$. ∎
-5 Numerical Example
-In this section, I will provide corresponding numerical examples based on the aforementioned theoretical background.
-Each of these conditional McKean-Vlasov SDEs involves both idiosyncratic and common sources of randomness, as well as jump components, and the coefficients satisfy the conditions required in (A1)-(A7).
-Model 1 (LQ conditional mean-field control I).
-${\rm{d}}X_{t}=u(t)\,{\rm{d}}t+\theta\,{\rm{d}}B_{t}^{0}+\sigma\,\mathbb{E}[X_{t}\mid\mathcal{F}_{t}^{B^{0}}]\,{\rm{d}}W_{t}+\int_{\mathbb{R}}\gamma_{0}(z)\,\mathbb{E}[X_{t}\mid\mathcal{F}_{t}^{B^{0}}]\,\tilde{N}({\rm{d}}t,{\rm{d}}z),\quad X_{0}=x_{0}\in\mathbb{R},$
-where $B^{0}$ is a common Brownian motion, $W$ is an idiosyncratic Brownian motion, and $\tilde{N}$ is a compensated Poisson random measure.
-The control process $u$ is adapted to the common filtration, and the jump amplitude $\gamma_{0}(z)$ is assumed to be bounded.
-Under these conditions, the coefficients satisfy (A1)-(A5) and (A7).
-In the numerical experiments, the jump distribution is taken to be Gaussian, so that (A4) also holds.
+$\exists~t_{k}<\tau_{R}$, for $s\in[t_{k},t_{k+1})\cap[0,\tau_{R}]$, by (4.5) we have
+$X^{\Delta}(s)$ $=X_{t_{k}}+\int_{t_{k}}^{s}b_{\Delta_{t_{k}}}\bigl(X_{t_{k}},\mathcal{L}^{1}(X_{t_{k}})\bigr)\,\mathrm{d}r+\int_{t_{k}}^{s}\sigma\bigl(X_{t_{k}},\mathcal{L}^{1}(X_{t_{k}})\bigr)\,\mathrm{d}W(r)$
+$\quad+\int_{t_{k}}^{s}c\bigl(X_{t_{k}},\mathcal{L}^{1}(X_{t_{k}})\bigr)\,\mathrm{d}Z(r)+\int_{t_{k}}^{s}\sigma^{0}\bigl(X_{t_{k}},\mathcal{L}^{1}(X_{t_{k}})\bigr)\,\mathrm{d}W^{0}(r).$
+Hence,
+$\mathbb{E}\big|X^{\Delta}(s)-\overline{X}^{\Delta}(s)\big|^{2}$ $\leq 4\mathbb{E}\bigg[\Big|\int_{t_{k}}^{s}b_{\Delta_{t_{k}}}(X_{t_{k}},\mathcal{L}^{1}(X_{t_{k}}))\,{\rm{d}}r\Big|^{2}+\Big|\int_{t_{k}}^{s}\sigma(X_{t_{k}},\mathcal{L}^{1}(X_{t_{k}}))\,{\rm{d}}W(r)\Big|^{2}$
+$\quad+\Big|\int_{t_{k}}^{s}c(X_{t_{k}},\mathcal{L}^{1}(X_{t_{k}}))\,{\rm{d}}Z(r)\Big|^{2}+\Big|\int_{t_{k}}^{s}\sigma^{0}(X_{t_{k}},\mathcal{L}^{1}(X_{t_{k}}))\,{\rm{d}}W^{0}(r)\Big|^{2}\bigg].$
+Step 1. Bound of the drift term. By Assumption (A3) and the definition of $\tau_{R}$, we have $|X_{t_{k}}|\leq R$ and $\mathcal{W}_{2}(\mathcal{L}^{1}(X_{t_{k}}),\delta_{0})\leq R$. Hence,
+$|b_{\Delta_{t_{k}}}(X_{t_{k}},\mathcal{L}^{1}(X_{t_{k}}))|\leq|b(X_{t_{k}},\mathcal{L}^{1}(X_{t_{k}}))|\leq|b(0,\delta_{0})|+2\tilde{L}(1+R^{\ell})R=:B_{R}.$
+Then
+$\mathbb{E}\Big|\int_{t_{k}}^{s}b_{\Delta_{t_{k}}}(X_{t_{k}},\mathcal{L}^{1}(X_{t_{k}}))\,\mathrm{d}r\Big|^{2}\leq(s-t_{k})^{2}B_{R}^{2}.$
+Step 2. Bound of the diffusion and jump terms. By Assumption (A7), for any $|x|$, $\mathcal{W}_{2}(\mathcal{L}^{1}(X_{t_{k}}),\delta_{0})\leq R$, we have
+$|\sigma(x,\mu)|^{2}+|\sigma^{0}(x,\mu)|^{2}+|c(x,\mu)|^{2}\!\int_{\mathbb{R}_{0}^{d}}|z|^{2}\nu(\mathrm{d}z)\leq 2\bigl(|\sigma(0,\delta_{0})|^{2}+|\sigma^{0}(0,\delta_{0})|^{2}+|c(0,\delta_{0})|^{2}I_{2}\bigr)+2(\tilde{L}_{1}+\tilde{L}_{2})R^{2},$
+where $I_{2}:=\int_{\mathbb{R}_{0}^{d}}|z|^{2}\nu(\mathrm{d}z)$. Denote this constant by
+$K_{R}:=2\bigl(|\sigma(0,\delta_{0})|^{2}+|\sigma^{0}(0,\delta_{0})|^{2}+|c(0,\delta_{0})|^{2}I_{2}\bigr)+2(\tilde{L}_{1}+\tilde{L}_{2})R^{2}.$
+Then, by the Itô isometry and the corresponding property for Lévy jumps,
+$\mathbb{E}\Big|\int_{t_{k}}^{s}\sigma(X_{t_{k}},\mathcal{L}^{1}(X_{t_{k}}))\,\mathrm{d}W(r)\Big|^{2}+\mathbb{E}\Big|\int_{t_{k}}^{s}\sigma^{0}(X_{t_{k}},\mathcal{L}^{1}(X_{t_{k}}))\,\mathrm{d}W^{0}(r)\Big|^{2}$
+$\qquad+\mathbb{E}\Big|\int_{t_{k}}^{s}c(X_{t_{k}},\mathcal{L}^{1}(X_{t_{k}}))\,\mathrm{d}Z(r)\Big|^{2}\leq 3K_{R}(s-t_{k}).$
+Combining the above bounds yields
+$\mathbb{E}|X^{\Delta}(s)-\overline{X}^{\Delta}(s)|^{2}\leq 4\Big[(s-t_{k})^{2}B_{R}^{2}+3K_{R}(s-t_{k})\Big].$
+Therefore, since $s\in[0,T]$ is arbitrary, we obtain
+$\mathbb{E}|X^{\Delta}(s)-\overline{X}^{\Delta}(s)|^{2}\leq 4\big(B_{R}^{2}+3K_{R}\big)h=:C_{R}\,h,\qquad\forall s\in[0,T].$
+Hence,
+$\int_{0}^{T}\mathbb{E}|X^{\Delta}(s)-\overline{X}^{\Delta}(s)|^{2}\,{\rm{d}}s\ \leq\ T\,C_{R}\,h,$
+where
+$C_{R}=4\big(B_{R}^{2}+3K_{R}\big),\quad B_{R}=|b(0,\delta_{0})|+2\tilde{L}(1+R^{\ell})R,$ (4.9)
+$K_{R}=2\bigl(|\sigma(0,\delta_{0})|^{2}+|\sigma^{0}(0,\delta_{0})|^{2}+|c(0,\delta_{0})|^{2}I_{2}\bigr)+2(\tilde{L}_{1}+\tilde{L}_{2})R^{2}.$
+Lemma 4.5.
+Suppose that there exist constants $C_{R,1},C_{R}^{(\ast)}>0$ such that
+$\mathbb{E}\bigg[\sup_{0\leq t\leq T}|e^{R}(t)|^{2}\bigg]\leq 5C_{R,1}h^{2\alpha}+C_{R}^{(\ast)}\int_{0}^{T}\mathbb{E}\big[|\overline{X}^{\Delta}(s)-X(s)|^{2}\cdot\mathbf{1}_{\{s\leq\tau_{R}\}}\big]\,{\rm d}s.$
+Then,
+$\mathbb{E}\bigg[\sup_{0\leq t\leq T}|e^{R}(t)|^{2}\bigg]\leq\Big(5C_{R,1}h^{2\alpha}+8C_{R}^{(\ast)}T(B_{R}^{2}+3K_{R})h\Big)e^{2C_{R}^{(\ast)}T}.$
+Proof.
+Due to Lemma 4.4, we have
+$\mathbb{E}\Big[\sup_{0\leq t\leq T}|e^{R}(t)|^{2}\Big]\leq 5C_{R,1}h^{2\alpha}+C_{R}^{(\ast)}\int_{0}^{T}\mathbb{E}\big|\overline{X}^{\Delta}(s)-X(s)\big|^{2}\mathbf{1}_{\{s\leq\tau_{R}\}}\,{\rm d}s$ (4.10)
+$\leq 5C_{R,1}h^{2\alpha}+C_{R}^{(\ast)}\int_{0}^{T}\!\Big(2\,\mathbb{E}\Big[\sup_{0\leq r\leq s}|e^{R}(r)|^{2}\Big]+2\,\mathbb{E}\big|X^{\Delta}(s)-\overline{X}^{\Delta}(s)\big|^{2}\Big){\rm d}s$
+$\leq 5C_{R,1}h^{2\alpha}+2C_{R}^{(\ast)}\int_{0}^{T}\mathbb{E}\Big[\sup_{0\leq r\leq s}|e^{R}(r)|^{2}\Big]\,{\rm d}s+8C_{R}^{(\ast)}T\big(B_{R}^{2}+3K_{R}\big)\,h.$
+By using Grönwall’s inequality to (4.10), we have
+$\mathbb{E}\Big[\sup_{0\leq t\leq T}|e^{R}(t)|^{2}\Big]\leq\big(5C_{R,1}h^{2\alpha}+8C_{R}^{(\ast)}T(B_{R}^{2}+3K_{R})\,h\big)\,e^{2C_{R}^{(\ast)}T}.$
+Proof of Theorem 4.3.
+For $t\in[0,\tau_{R}]$, the coefficients are globally Lipschitz with constants depending only on $R$. Note that $e^{R}(t)=X^{\Delta,R}(t)-X^{R}(t)$. Subtracting the integral equations gives
+$e^{R}(t)$ $=\int_{0}^{t\wedge\tau_{R}}\big(b_{\Delta_{t_{s}}}(\overline{X}^{\Delta}(s),\bar{\mu}_{s})-b(\overline{X}^{\Delta}(s),\bar{\mu}_{s})\big)\,\mathrm{d}s+\int_{0}^{t\wedge\tau_{R}}\big(b(\overline{X}^{\Delta}(s),\bar{\mu}_{s})-b(X(s),\mu_{s})\big)\,\mathrm{d}s$
+$\quad+\int_{0}^{t\wedge\tau_{R}}\big(\sigma(\overline{X}^{\Delta}(s),\bar{\mu}_{s})-\sigma(X(s),\mu_{s})\big)\,\mathrm{d}W(s)+\int_{0}^{t\wedge\tau_{R}}\big(c(\overline{X}^{\Delta}(s),\bar{\mu}_{s})-c(X(s),\mu_{s})\big)\,\mathrm{d}Z(s)$
+$\quad+\int_{0}^{t\wedge\tau_{R}}\big(\sigma^{0}(\overline{X}^{\Delta}(s),\bar{\mu}_{s})-\sigma^{0}(X(s),\mu_{s})\big)\,\mathrm{d}W^{0}(s)=:e_{1}(t)+e_{2}(t)+e_{3}(t)+e_{4}(t)+e_{5}(t),$
+where $\bar{\mu}_{s}:=\mathcal{L}^{1}(\overline{X}^{\Delta}(s))$, $\mu_{s}:=\mathcal{L}^{1}(X(s))$.
+Applying $(a_{1}+\cdots+a_{5})^{2}\leq 5(a_{1}^{2}+\cdots+a_{5}^{2})$ and the BDG inequality, we obtain
+$\mathbb{E}\Big[\sup_{0\leq t\leq T}|e^{R}(t)|^{2}\Big]$ $\leq 5\,\mathbb{E}\Big[\sup_{0\leq t\leq T}|e_{1}(t)|^{2}\Big]+5\,\mathbb{E}\Big[\sup_{0\leq t\leq T}|e_{2}(t)|^{2}\Big]+5\,\mathbb{E}\Big[\sup_{0\leq t\leq T}|e_{3}(t)|^{2}\Big]$ (4.11)
+$\quad+5\,\mathbb{E}\Big[\sup_{0\leq t\leq T}|e_{4}(t)|^{2}\Big]+5\,\mathbb{E}\Big[\sup_{0\leq t\leq T}|e_{5}(t)|^{2}\Big]$
+$=:5(I_{1}+I_{2}+I_{3}+I_{4}+I_{5}),$
+where $I_{i}=\mathbb{E}\big[\sup_{0\leq t\leq T}|e_{i}(t)|^{2}\big]$.
+Step 1: Estimate of $I_{1}$. On $0,\tau_{R}]$ we have $|\overline{X}^{\Delta}(s)|\leq R$ and $\mathcal{W}_{2}(\bar{\mu}_{s},\delta_{0})\leq R$. By Assumption (A3),
+$|b(\overline{X}^{\Delta}(s),\bar{\mu}_{s})|\leq|b(0,\delta_{0})|+2\tilde{L}(1+R^{\ell})R=B_{R}.$
+Hence for $s\leq\tau_{R}$,
+$\big|b_{\Delta_{t_{s}}}(\overline{X}^{\Delta}(s),\bar{\mu}_{s})-b(\overline{X}^{\Delta}(s),\bar{\mu}_{s})\big|\leq\Delta_{t_{s}}^{\alpha}B_{R}^{2}\leq h^{\alpha}B_{R}^{2}.$
+Therefore,
+$\sup_{0\leq t\leq T}\Big|\int_{0}^{t\wedge\tau_{R}}\big(b_{\Delta_{t_{s}}}(\overline{X}^{\Delta}(s),\bar{\mu}_{s})-b(\overline{X}^{\Delta}(s),\bar{\mu}_{s})\big){\rm d}s\Big|^{2}\leq(Th^{\alpha}B_{R}^{2})^{2}=T^{2}B_{R}^{4}\,h^{2\alpha}.$
+Taking expectation gives
+$I_{1}\leq C_{R,1}h^{2\alpha}\quad\text{with}\quad C_{R,1}:=T^{2}B_{R}^{4}=T^{2}\big[|b(0,\delta_{0})|+2\tilde{L}(1+R^{\ell})R\big]^{4}.$ (4.12)
+Step 2: Estimate of $I_{2}$. By Assumption (A3), for $s\leq\tau_{R}$,
+$|b(\overline{X}^{\Delta}(s),\bar{\mu}_{s})-b(X(s),\mu_{s})|\leq L_{b}(R)\big(|\overline{X}^{\Delta}(s)-X(s)|+\mathcal{W}_{2}(\bar{\mu}_{s},\mu_{s})\big),$
+where
+$L_{b}(R):=\tilde{L}(1+2R^{\ell}).$
+Then, using $\sup_{0\leq t\leq T}|\int_{0}^{t}f(s){\rm d}s|^{2}\leq T\int_{0}^{T}|f(s)|^{2}{\rm d}s$ and $\mathcal{W}_{2}^{2}(\bar{\mu}_{s},\mu_{s})\leq\mathbb{E}|\overline{X}^{\Delta}(s)-X(s)|^{2}$, we have
+$I_{2}$ $=\mathbb{E}\Big[\sup_{0\leq t\leq T}\Big|\int_{0}^{t\wedge\tau_{R}}\big(b(\overline{X}^{\Delta},\bar{\mu})-b(X,\mu)\big){\rm d}s\Big|^{2}\Big]$
+$\leq T\mathbb{E}\int_{0}^{T\wedge\tau_{R}}|b(\overline{X}^{\Delta}(s),\bar{\mu}_{s})-b(X(s),\mu_{s})|^{2}{\rm d}s$
+$\leq T\cdot 2L_{b}(R)^{2}\int_{0}^{T\wedge\tau_{R}}\mathbb{E}\Big[\,|\overline{X}^{\Delta}(s)-X(s)|^{2}+\mathcal{W}_{2}^{2}(\bar{\mu}_{s},\mu_{s})\Big]{\rm d}s$
+$\leq T\cdot 2L_{b}(R)^{2}\int_{0}^{T\wedge\tau_{R}}\mathbb{E}\Big[\,|\overline{X}^{\Delta}(s)-X(s)|^{2}+\mathbb{E}|\overline{X}^{\Delta}(s)-X(s)|^{2}\Big]{\rm d}s$
+$\leq 4TL_{b}(R)^{2}\int_{0}^{T\wedge\tau_{R}}\mathbb{E}|\overline{X}^{\Delta}(s)-X(s)|^{2}{\rm d}s.$
+Hence we can take
+$C_{R,2}:=4T\bigl(\tilde{L}(1+2R^{\ell})\bigr)^{2}\quad\text{and}\quad I_{2}\leq C_{R,2}\int_{0}^{T\wedge\tau_{R}}\mathbb{E}|\overline{X}^{\Delta}(s)-X(s)|^{2}{\rm d}s.$ (4.13)
+Step 3: Estimate of $I_{3}$. Let $g(s):=\sigma(\overline{X}^{\Delta}(s),\bar{\mu}_{s})-\sigma(X(s),\mu_{s})$. By the BDG inequality,
+$\mathbb{E}\Big[\sup_{0\leq t\leq T}\Big|\int_{0}^{t\wedge\tau_{R}}g(s)\,{\rm d}W(s)\Big|^{2}\Big]\leq C_{\mathrm{BDG}}\,\mathbb{E}\int_{0}^{T\wedge\tau_{R}}|g(s)|^{2}{\rm d}s.$
+By Assumption (A7),
+$|g(s)|^{2}\leq\tilde{L}_{1}|\overline{X}^{\Delta}(s)-X(s)|^{2}+\tilde{L}_{2}\mathcal{W}_{2}^{2}(\bar{\mu}_{s},\mu_{s})\leq(\tilde{L}_{1}+\tilde{L}_{2})\,\mathbb{E}|\overline{X}^{\Delta}(s)-X(s)|^{2}.$
+Hence,
+$I_{3}\leq C_{\mathrm{BDG}}(\tilde{L}_{1}+\tilde{L}_{2})\int_{0}^{T\wedge\tau_{R}}\mathbb{E}|\overline{X}^{\Delta}(s)-X(s)|^{2}{\rm d}s.$
+So we set
+$C_{R,3}:=C_{\mathrm{BDG}}(\tilde{L}_{1}+\tilde{L}_{2})\quad\text{and}\quad I_{3}\leq C_{R,3}\int_{0}^{T\wedge\tau_{R}}\mathbb{E}|\overline{X}^{\Delta}(s)-X(s)|^{2}{\rm d}s.$ (4.14)
+Step 4: Estimate of $I_{4}$. For the jump part, let
+$h(s,z):=\big(c(\overline{X}^{\Delta}(s),\bar{\mu}_{s})-c(X(s),\mu_{s})\big)z,$
+and let $M_{t}:=\int_{0}^{t}\int_{\mathbb{R}_{0}^{d}}h(s,z)\,\tilde{N}({\rm d}s,{\rm d}z)$. By Kunita’s inequality,
+$I_{4}=\mathbb{E}\Big[\sup_{0\leq t\leq T}|M_{t\wedge\tau_{R}}|^{2}\Big]\leq C_{\mathrm{Ku}}\,\mathbb{E}\int_{0}^{T\wedge\tau_{R}}\int_{\mathbb{R}_{0}^{d}}|h(s,z)|^{2}\nu({\rm d}z)\,{\rm d}s.$
+Using assumptions (A4) and (A7),
+$\int_{\mathbb{R}_{0}^{d}}|h(s,z)|^{2}\nu({\rm d}z)$ $=|c(\overline{X}^{\Delta}(s),\bar{\mu}_{s})-c(X(s),\mu_{s})|^{2}\int_{\mathbb{R}_{0}^{d}}|z|^{2}\nu({\rm d}z)$
+$\leq(\tilde{L}_{1}+\tilde{L}_{2})\big(|\overline{X}^{\Delta}(s)-X(s)|^{2}+\mathcal{W}_{2}^{2}(\bar{\mu}_{s},\mu_{s})\big).$
+Due to $\mathcal{W}_{2}^{2}(\bar{\mu}_{s},\mu_{s})\leq\mathbb{E}|\overline{X}^{\Delta}(s)-X(s)|^{2}$, we have
+$I_{4}\leq 2C_{\mathrm{Ku}}(\tilde{L}_{1}+\tilde{L}_{2})\int_{0}^{T\wedge\tau_{R}}\mathbb{E}|\overline{X}^{\Delta}(s)-X(s)|^{2}{\rm d}s.$
+Hence,
+$C_{R,4}:=2C_{\mathrm{Ku}}(\tilde{L}_{1}+\tilde{L}_{2})\quad\text{and}\quad I_{4}\leq C_{R,4}\int_{0}^{T\wedge\tau_{R}}\mathbb{E}|\overline{X}^{\Delta}(s)-X(s)|^{2}{\rm d}s.$ (4.15)
+Step 5: Estimate of $I_{5}$. This is completely analogous to the estimate of $I_{3}$. Thus,
+$I_{5}\leq C_{\mathrm{BDG}}(\tilde{L}_{1}+\tilde{L}_{2})\int_{0}^{T\wedge\tau_{R}}\mathbb{E}|\overline{X}^{\Delta}(s)-X(s)|^{2}{\rm d}s=:C_{R,5}\int_{0}^{T\wedge\tau_{R}}\mathbb{E}|\overline{X}^{\Delta}(s)-X(s)|^{2}{\rm d}s.$ (4.16)
+Putting (4.12), (4.13), (4.14), (4.15), (4.16) back into (4.11), we obtain
+$\mathbb{E}\Big[\sup_{0\leq t\leq T}|e^{R}(t)|^{2}\Big]\leq 5C_{R,1}h^{2\alpha}+5\big(C_{R,2}+C_{R,3}+C_{R,4}+C_{R,5}\big)\int_{0}^{T\wedge\tau_{R}}\mathbb{E}|\overline{X}^{\Delta}(s)-X(s)|^{2}{\rm d}s.$ (4.17)
+If we denote
+$C_{R}^{(\ast)}:=5\big(C_{R,2}+C_{R,3}+C_{R,4}+C_{R,5}\big),$ (4.18)
+then the estimate (4.17) becomes
+$\mathbb{E}\Big[\sup_{0\leq t\leq T}|e^{R}(t)|^{2}\Big]\leq 5C_{R,1}h^{2\alpha}+C_{R}^{(\ast)}\int_{0}^{T\wedge\tau_{R}}\mathbb{E}|\overline{X}^{\Delta}(s)-X(s)|^{2}{\rm d}s.$
+To obtain the global error estimate for $e(t)=X^{\Delta}(t)-X(t)$,
+$\mathbb{E}\left[\sup_{0\leq t\leq T}|e(t)|^{2}\right]$ $\leq\mathbb{E}\left[\sup_{0\leq t\leq T}|e^{R}(t)|^{2}\right]+\mathbb{E}\left[\sup_{0\leq t\leq T}|e(t)|^{2}\cdot\mathbf{1}_{\{\tau_{R}\leq T\}}\right].$
+According to Proposition 2.5 and Lemma 4.2, for $p\geq 2$ we have,
+$\mathbb{E}\left[\sup_{0\leq t\leq T}|X(t)|^{p}\right]\leq C_{T,p},\quad\mathbb{E}\left[\sup_{0\leq t\leq T}|X^{\Delta}(t)|^{p}\right]\leq C_{T,p}.$ (4.19)
+Now, by using Hölder’s inequality,
+$\mathbb{E}\left[\sup_{0\leq t\leq T}|e(t)|^{2}\cdot\mathbf{1}_{\{\tau_{R}\leq T\}}\right]$ $\leq\left(\mathbb{E}\left[\sup_{0\leq t\leq T}|e(t)|^{p}\right]\right)^{2/p}\cdot\left(\mathbb{P}(\tau_{R}\leq T)\right)^{1-2/p}.$
+By the triangle inequality and (4.19), we obtain
+$\mathbb{E}\left[\sup_{0\leq t\leq T}|e(t)|^{p}\right]\leq 2^{p-1}\left(\mathbb{E}\left[\sup_{0\leq t\leq T}|X(t)|^{p}\right]+\mathbb{E}\left[\sup_{0\leq t\leq T}|X^{\Delta}(t)|^{p}\right]\right)\leq 2^{p}C_{T,p}.$
+Now, by Markov’s inequality,
+$\mathbb{P}(\tau_{R}\leq T)\leq\mathbb{P}\left(\sup_{0\leq t\leq T}|X(t)|\geq R\right)+\mathbb{P}\left(\sup_{0\leq t\leq T}|X^{\Delta}(t)|\geq R\right)\leq\frac{2C_{T,p}}{R^{p}}.$
+Thus,
+$\mathbb{E}\left[\sup_{0\leq t\leq T}|e(t)|^{2}\cdot\mathbf{1}_{\{\tau_{R}\leq T\}}\right]\leq(2^{p}C_{T,p})^{2/p}\cdot\left(\frac{2C_{T,p}}{R^{p}}\right)^{1-2/p}\leq C_{T,p}R^{2-p},$ (4.20)
+combining with Lemma 4.5 we have
+$\mathbb{E}\left[\sup_{0\leq t\leq T}|e(t)|^{2}\right]$ $\leq 5C_{R,1}h^{2\alpha}+C_{R}^{(\ast)}\int_{0}^{T\wedge\tau_{R}}\mathbb{E}|\overline{X}^{\Delta}(s)-X(s)|^{2}{\rm d}s+\mathbb{E}\left[\sup_{0\leq t\leq T}|e(t)|^{2}\cdot\mathbf{1}_{\{\tau_{R}\leq T\}}\right]$ (4.21)
+$\leq\Big(5C_{R,1}h^{2\alpha}+8C_{R}^{(\ast)}T\big(B_{R}^{2}+3K_{R}\big)\,h\Big)\,e^{2C_{R}^{(\ast)}\,T}\ +\,C_{T,p}R^{2-p}$
+$=\Big(5T^{2}\big[|b(0,\delta_{0})|+2\tilde{L}(1+R^{\ell})R\big]^{4}h^{2\alpha}+8C_{R}^{(\ast)}T\big(B_{R}^{2}+3K_{R}\big)\,h\Big)\,e^{2C_{R}^{(\ast)}\,T}\ +\,C_{T,p}R^{2-p}$
+$=:(g_{1}+g_{2})\,e^{2C_{R}^{(\ast)}T}\;+\;g_{3}.$
+Due to (4.9) and (4.18), we have
+$C_{R}^{(\ast)}\leq\bar{C}_{1}\bigl(1+R^{2\ell}\bigr),\qquad B_{R}^{2}+3K_{R}\leq\bar{C}_{2}\bigl(1+R^{2\ell+2}\bigr),$ (4.22)
+for some constants $\bar{C}_{1},\bar{C}_{2}>0$ independent of $h$. Let
+$R=\left(\frac{\varepsilon}{4\bar{C}_{1}T}\,\log\!\frac{1}{h}\right)^{\!1/(2\ell)},\qquad 0<\varepsilon<1.$ (4.23)
+Hence, we have
+$e^{2C_{R}^{(\ast)}T}\leq e^{\,2\bar{C}_{1}T\,(1+R^{2\ell})}=e^{2\bar{C}_{1}T}\,\exp\!\Big(2\bar{C}_{1}T\cdot\frac{\varepsilon}{4\bar{C}_{1}T}\,\log\!\tfrac{1}{h}\Big)=e^{2\bar{C}_{1}T}\,h^{-\varepsilon/2}.$ (4.24)
+We now bound each term in (4.21) under the choice (4.23):
+Step 1: Estimate of $g_{1}$. Since (4.23), we have
+$5T^{2}\left[|b(0,\delta_{0})|+2\tilde{L}(1+R^{\ell})R\right]^{4}h^{2\alpha}\leq C_{1}\,h^{2\alpha}\,\bigl(\log(1/h)\bigr)^{\frac{2(\ell+1)}{\ell}}.$ (4.25)
+Step 2: Estimate of $g_{2}$. Due to (4.22), we get
+$C_{R}^{(\ast)}(B_{R}^{2}+3K_{R})\leq\bar{C}_{1}(1+R^{2\ell})\cdot\bar{C}_{2}(1+R^{2\ell+2})=\bar{C}_{1}\bar{C}_{2}\bigl(1+R^{2\ell}+R^{2\ell+2}+R^{4\ell+2}\bigr).$
+Hence,
+$C_{R}^{(\ast)}(B_{R}^{2}+3K_{R})\leq C_{2}\,\bigl(\log(1/h)\bigr)^{\frac{4\ell+2}{2\ell}}.$ (4.26)
+Therefore, since in (4.21) we have $g_{2}=8\,C_{R}^{(\ast)}T\,(B_{R}^{2}+3K_{R})\,h$, combining with (4.26) yields $g_{2}\leq 8T\,C_{2}\,h\,(\log(1/h))^{\frac{4\ell+2}{2\ell}}.$
+Step 3: Estimate of $g_{3}$.
+$C_{T,p}R^{2-p}=C_{T,p}\,\bigl(\log(1/h)\bigr)^{\frac{2-p}{2\ell}}.$ (4.27)
+Substituing (4.24), (4.25), (4.26), (4.27) into (4.21), we obtain
+$\mathbb{E}\left[\sup_{0\leq t\leq T}|e(t)|^{2}\right]$ $\leq\Big[C_{1}\,h^{2\alpha}\bigl(\log(1/h)\bigr)^{\frac{2(\ell+1)}{\ell}}\ +\,8T\,C_{2}\,h\,\bigl(\log(1/h)\bigr)^{\frac{4\ell+2}{2\ell}}\Big]\cdot h^{-\varepsilon/2}\ +\,C_{T,p}\,\bigl(\log(1/h)\bigr)^{\frac{2-p}{2\ell}}$
+$=C_{1}\,h^{2\alpha-\frac{\varepsilon}{2}}\,\bigl(\log(1/h)\bigr)^{\frac{2(\ell+1)}{\ell}}\ +\,8T\,C_{2}\,h^{1-\frac{\varepsilon}{2}}\,\bigl(\log(1/h)\bigr)^{\frac{4\ell+2}{2\ell}}\ +\,C_{T,p}\,\bigl(\log(1/h)\bigr)^{-\frac{p-2}{2\ell}}.$
+Choosing $\alpha\in\bigl(\tfrac{1}{2}-\tfrac{\varepsilon}{4},\,\tfrac{1}{2}\bigr)$ gives $2\alpha-\tfrac{\varepsilon}{2}\geq 1-\varepsilon$. Hence,
+$\mathbb{E}\left[\sup_{0\leq t\leq T}|e(t)|^{2}\right]\leq C_{T,\varepsilon}\,h^{1-\varepsilon}\,(\log(1/h))^{C}\ +\,C_{T,p}\,\bigl(\log(1/h)\bigr)^{-\frac{p-2}{2\ell}},$
+for a constant $C>0$ depending only on $\ell$. Consequently, the above bound guarantees strong convergence,
+$\lim_{h\to 0}\mathbb{E}\left[\sup_{0\leq t\leq T}|e(t)|^{2}\right]=0.$
+5 Numerical Examples
+In this section we present four conditional McKean-Vlasov SDE test problems with common noise and jumps, and examine the performance of the TAEM scheme.
+All model coefficients satisfy (A1)-(A7).
+For a final time $T$ and a family of dyadic time grids indexed by $l\in\{1,\ldots,6\}$ with time step $h_{l}=T/2^{l}$, we define the level-$l$ mean-squared error (MSE) against the finest reference level $L_{\max}$ under a fully coupled driver construction:
+$\mathrm{MSE}(l,T):=\frac{1}{M}\sum_{m=1}^{M}\big|X^{(l)}_{m}(T)-X^{(L_{\max})}_{m}(T)\big|^{2},\qquad\beta:=-\mathrm{Slope}\big(\,l\mapsto\log_{2}\mathrm{MSE}(l,T)\,\big).$
+Hence $\beta\approx 1$ corresponds to an almost first-order decay of the MSE with respect to $h$ (since $\mathrm{MSE}\approx O(h)$ implies a root mean-square error of order $O(h^{1/2})$).
+We report the fitted $\beta$ for $T\in\{1,5,10\}$ with each model and show the aggregate plot in Figure 1.
+Model 1: Ginzburg-Landau model with common noise.
+${\rm d}X_{t}=0.1\bigl(X_{t}-X_{t}^{3}+m(\mu_{t})\bigr)\,{\rm d}t+0.1X_{t}\,{\rm d}W_{t}+0.05X_{t}\,{\rm d}W_{t}^{0}+\sin(X_{t-})\,{\rm d}Z_{t}^{0},\qquad X_{0}=1.0,$
+where $\mu_{t}=\mathcal{L}^{1}(X_{t})$ and $m(\mu_{t})=\int_{\mathbb{R}}y\,\mu_{t}({\rm d}y)$.
+Here $W$ denotes an idiosyncratic Brownian motion, $W^{0}$ a common Brownian motion, and $Z^{0}$ a pure-jump Lévy process (e.g., tempered-stable).
+Due to Example 1, assumptions (A1)-(A7) are satisfied.
+For Model 1, Figure 1 shows that the TAEM scheme exhibits an almost first-order decay of the MSE at short horizon, with $\beta\approx 1.15$ for $T=1$.
+At $T=5$, we observe a pre-asymptotic superlinear behaviour with $\beta\approx 1.78$.
+At $T=10$, the fitted rate turns negative ($\beta\approx-0.48$) and the MSE increases upon refinement, reflecting long-horizon instability driven by strong nonlinearity under common noise.
+Model 2: LQ conditional mean-field control I model.
+${\rm d}X_{t}=u(t)\,{\rm d}t+0.15\,{\rm d}B_{t}^{0}+0.12\,\mathbb{E}[X_{t}\mid\mathcal{F}_{t}^{B^{0}}]\,{\rm d}W_{t}+\int_{\mathbb{R}}0.08\tanh(z)\,\tilde{N}({\rm d}t,{\rm d}z),\quad X_{0}=1.0,$
+with control
+$u(t)=-0.5\left(X_{t}+\mathbb{E}[X_{t}\mid\mathcal{F}_{t}^{B^{0}}]\right).$
+Here $B^{0}$ denotes the common Brownian motion, $W$ an idiosyncratic Brownian motion, and $\tilde{N}$ a compensated Poisson random measure.
+The jump distribution is Gaussian with variance $\varsigma^{2}=0.25$, and the jump intensity is $\lambda=1.0$.
+Under these specifications, assumptions (A1)-(A7) are satisfied.
+For Model 2, Figure 1 shows that we obtain $\beta\approx 1.01,\,1.59,\,1.91$ for $T=1,5,10$, respectively.
+Accuracy remains robust and even improves within our tested range, reflecting that the control effectively neutralizes both idiosyncratic and common fluctuations.
+Model 3: LQ conditional mean-field control II model.
+${\rm d}X_{t}=u(t)\,{\rm d}t+\int_{\mathbb{R}}0.18\tanh(z)\,\tilde{N}_{1}({\rm d}t,{\rm d}z)+\int_{\mathbb{R}}0.10\tanh(z)\,\mathbb{E}[X_{t}\mid\mathcal{F}_{t}^{\tilde{N}_{1}}]\,\tilde{N}_{2}({\rm d}t,{\rm d}z),\quad X_{0}=1.0,$
+with control
+$u(t)=-0.4\left(X_{t}+\mathbb{E}[X_{t}\mid\mathcal{F}_{t}^{\tilde{N}_{1}}]\right).$
+Here $\tilde{N}_{1}$ and $\tilde{N}_{2}$ are independent compensated Poisson random measures, representing idiosyncratic and common jumps, respectively.
+Both jump distributions are Gaussian with variance $0.16$, with intensities $\lambda_{1}=0.8$ and $\lambda_{2}=0.6$.
+By the Lipschitz and linear growth conditions, (A1)-(A7) are fulfilled.
+For Model 3, Figure 1 shows that the fitted rates are $\beta\approx 1.04,\,1.33,\,1.79$ for $T=1,5,10$, respectively, again with slopes around or above $1$ and stable behaviour as $T$ grows.
+Model 4: Interbank lending model with jumps.
+${\rm d}X_{t}=0.6\bigl(\bar{X}_{t}-X_{t}\bigr)\,{\rm d}t+0.25\,{\rm d}W_{t}+0.18\,{\rm d}B_{t}^{0}+\int_{\mathbb{R}}0.12\tanh(z)\,\tilde{N}({\rm d}t,{\rm d}z),\quad X_{0}=1.0,$
+where $\bar{X}_{t}=\mathbb{E}[X_{t}\mid\mathcal{F}_{t}^{B^{0}}]$ denotes the conditional mean with respect to the common noise.
+This model captures mean-reversion toward the conditional average across banks, together with both idiosyncratic and common volatility.
+The jump distribution is Gaussian with variance $0.2025$, and the jump intensity is $\lambda=0.9$.
+Again, (A1)-(A7) are satisfied.
+For Model 4, Figure 1 shows that we obtain $\beta\approx 1.38,\,1.64,\,1.77$ for $T=1,5,10$, respectively, with consistently large slopes and good large-$T$ behaviour; mean reversion stabilizes the dynamics against the combined effect of common Brownian noise and Gaussian jumps.
+Figure 1: $\log_{2}\mathrm{MSE}(l,T)$ against $l$ for the four models ($T\!\in\!\{1,5,10\}$); dashed line: slope 1-1.
+Conclusion.
+In summary, the numerical experiments display an almost first-order decay of the MSE in most cases (with fitted slopes $\beta\approx 1$), which complements the strong convergence result in Theorem 4.3.
+Its long-horizon robustness hinges on the balance among nonlinearity strength, common-noise coupling, and jump activity.
+Across the four models, LQ-I/II and Interbank remain stable over extended horizons, while the Ginzburg-Landau model shows a transient non-asymptotic behaviour only at very large $T$ due to strong nonlinearity.
+Nevertheless, across practically relevant horizons all models demonstrate satisfactory convergence behaviour and accuracy at the tested resolutions.
+Moreover, the inclusion of common noise in all four models is essential.
+In practical financial and economic systems, individual agents are not only exposed to their own idiosyncratic shocks, but also to systemic sources of randomness such as market-wide fluctuations, macroeconomic policy shifts, or collective risk factors.
+This type of shared uncertainty governs the correlation structure among agents and strongly affects the evolution of the mean field as well as the effectiveness of control strategies.
+Neglecting common noise would lead to an overestimation of convergence and stability, and would fail to capture the mechanism of error accumulation over long horizons.
+Hence, incorporating common noise is crucial for accurately describing systemic risk propagation, assessing the robustness of numerical schemes, and ensuring that the models remain relevant for real-world applications.
+Acknowledgments
+We would like to express our gratitude to all those who helped us with this research. Special thanks to Professor Jianhai Bao and Professor Jian Wang.